A Complete Bibliography of The Journal of the American Statistical Association: 1930–1939

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.05

Title word cross-reference

2 × s [Bra33a]. χ^{2} [Fry38].

-Test [Fry38].

13-Month [Ano34a]. 1929 [Ano30e]. 1929-31 [Fab35]. 1930 [Ano30f, Ano30h, Ano31e]. 1931 [Ano31g, Ano31f, Ano31h, Ano32g]. 1932 [Ano32h, Ano32j, Ano33h]. 1932-33 [Ano32e]. 1933 [Ano33j, Ano33i, Ano33k, Ano34l, Sch35]. 1933-34 [Ano33b]. 1934 [Ano34n, Ano34m, Ano34o]. 1936 [Asc37, Pea36b]. 1937 [Ano38t].

A.A.A. [CB36]. A.M.I [Eva36]. A.M.II [Eva36]. Abroad [Kin34e]. Absolute [Spi34]. Absorption [Wei32c]. Abstract [Mar34b]. Accident [Ano37a, Ano37i, Kos39]. Accidents [Cam31]. According [Ayl37, SN30]. Accounting [Met36, Met38]. Accuracy [Dav33a, Dow31]. Acreage [Bea30b]. Act [Kin33g]. Activities [Ano35c, Ano35d, Ano35e, Ano36a, Ano36b, Ano36c, Ano36d, Ano37c, Ano37d, Ano37e, Ano37f, Ano38b, Ano38c, Ano38d, Ano38e, Ano39b, Ano39c, Ano39d, Ano39e, Bra36a, MF35, Ste35a]. Activity [Bra34]. Actual [Fra31]. Adaptation [Dun32, Sma33]. Added [Ano30e, Ano30g, Ano30f, Ano30h, Ano31e, Ano31g, Ano31f, Ano31h, Ano32g, Ano32i, Ano32h, Ano32j, Ano33h, Ano33j, Ano33i, Ano33k, Ano34l, Ano34n, Ano34m, Ano34o]. Additional [Ano31a, Ano33a]. Adequacy [Han34b, Kin33a]. Adjusting [Hom33, Vee30, Wel30]. Adjustment [Lor35, Pis34, Sar34, vH33c]. Adjustments [Hum37]. Administration [Alt37, Big37, Bow36a, Cla37, Sar34, Ros34]. Advertising [Kin32f]. Advisory [Giv34c]. Affecting [Mit39b]. Age [Lor35, SN30, Tay32]. Agencies [DuB36]. Agents [Per36]. Ages [Smi35c]. Aggregates [Art37]. Agreement [Sar34]. Agreements [DF37, Fet37]. Agricultural [CB36, Dav33a, Sar34, Wai38, Wis36]. Agriculture [H.34e, H.35d]. Agronomic [Ric30a, Ric30b]. Aid [WD32]. Algebraic [Ham33b]. Allocation [Ayl37, Dor32, Lin37a, Sch33a]. Alternative [Ste31]. Amendment [Wil31]. America [Cha33b]. American [Ano30b, Ano30i, Ano30q, Ano33f, Bad31, Bad33, Dod37, Dor39, H.34d, Kin30f, KSF36, Kuz35a, RG30, Ste35a, Ste38b, Wil34b]. Among [Mal36, Ogb35, Sch33a, Tho37b]. Amplitude [Kuz32b]. Analysis [Ano36h, Bai31, Bra33a, Coc39, Fri37a, Fri39b, GS36, Kem34, Lot38, McI37, Men39b, Ric31, SS33, Sti34, Sto36, Tho37a, Wau36, Wei38, Yat34, Wor34]. Analytical [Cru33]. Analyzing [Kin31b, MK37]. Annals [Ano38a]. Annual [Ano34r, Ano36j, Ano36y, Ano37s, Ano38t, Ano39n, Fis30, Kin34c, Ror30]. Apparatus [Sib30]. Application [Bea30b, Ber38, Bow36b, Dod37, HG30, Rom35, Sto34c]. Applications [Ano31d, Ney35, Ric30a, Ric30b, Rie31c]. Applied [Ric37]. Applying [Sto33]. Appointments [Ano31a, Ano33a]. Appraising [Kin30a]. Approach [Kos39]. Area [HCW^{+}34]. Areas [Cha34a]. Arising [Ano38f, Mue38, Sel34]. Arithmetic [vH30c]. Arrangement [Ano36f]. Art [Ogb32]. Artisans [Tuc36b]. Aspects [Bra36a, Dav39, Giv31b, Kin31f, Wil38e]. Assets [Car38]. Assistance [WW38]. Associated [HG30]. Association [Ano30b, Ano30q, Ano33f, Bra36b, Hal36, RG30, Ste35a, Ste38b, Wai38, Ano30i, Kuz35a]. Associations [Bra36a]. Assumption [Fri37a, Fri39b]. Athens [Wil37b]. Atlantic [Ano38t]. Attempt [Nyb37, Wil31]. Attitudes [Kin31b, Kuz35a]. Attorney [San36a]. Authorities [Com34]. Authority [Che34, Kin35b]. Automatic [McI37]. Automobile [Kin39d, RvS39a, RvS39b, Sel34]. Automobiles [Kin39c]. Average [Spi34, vH30c]. Avoid [Fri37a, Fri39b]. Awards [Ano30a, Ano31b, Ano33b].

Back [Ano31c, Ano32a, Ano33c, Ano34b, Ano34c, Ano37b, Ano39a]. Balance [Dic38, Kuc30, Lot30, Tay37]. Balance-of-Payments [Tay37]. Bank [Bar38, Fox36a, Jon38, Kin30b, Tho38a]. Banking [Fox36b, Gol38].

Banks [Gol36]. Bars [Cro32b]. Base [Ano38w]. Based [Sma33, SN30]. Bases [Ray36]. Basis [Gir39, Pat39, Slo37]. be [Kin34d, Kin35b, Kin39c]. Before [Wol32]. Being [Kin34a]. Benefit [Tol34]. Better [Ano35a, Kin30d]. Between [HG30, Imm37, Lot36, Zub39, vH30c, vH33b]. Bi [Wal34]. Bi-Centenary [Wal34]. Bias [Hin38, Vic38]. Biennial [Wil30a]. Big [TBSM38]. Biology [Sne36]. Biostatistical [CSS39]. Birth [CJ32, HH31, Kop32b, Whe34a, Whe34b]. Births [Kuc30, Lot30]. Board [Alt37, Ano33d, Ano34d, Ano35b, Ano36w, Dic35, MF35]. Bond [Tho33, Tur36, Tur39]. Book [Abb38, Abe34, Abr31, Abr35, Ach30, Ach34a, Ada33, Ada39, Ale37a, Ale37b, AE38, Alt35, Alt36, Amo37, Amo38, And35, And39, And30, Ang30, Ang32, Ang33, Ang39, Ano39n, Arn31, Art36, Bac31, Bal36, Bar37, Bar33, Bau36, Bau33, Bea31a, Bea30a, Bea31b, Bea35, Bec37, Bec38b, Bel31, Ben37a, Ber36a, Bez39, Bid34, Bla34, Bla37, Bla38, Bla36, Bla33, Bli32, Blo31, Blo37, Bob38, Bod39, Bog35, Bog36, Bog37, Bog38a, Boi36, Bor38, Bow35b, Boy33, Bra30, Bre39, Bri35, Bri31, Bro31, Bro38, Bro39, Bro36, Bru36, Bru37, Bru35, Buc38a, Buc38b, Bue38a, Bue38b, Bue38c, Bur32a, Bur30, Bur36a, Bur31, Bur36b, Bur39, Cam38a, Can39, Car36, Car30, Car35, Car31, Cav37, Cha30a, Cha33a, Cha39, Cha34b, Cha35a, Cha38, Cio37, Cio38, Cio39]. Book [Cla31, Cla32, Cla33, Cla39, Cli39, Coa31a, Col37, Col38, Com33, Cop30, Cop32b, Cor34, Cov37, Cov38, Cox36, Cox38, Cox39a, Cox39b, Cox39c, Cra35, Cra36, Cra37, Cre38, Cro38a, Cro32a, Cro30, Cro33, Cro36, Cro38b, Cro38c, Cru31, Cru35a, Cru36, Cun33, Cus36, Cus39, D.37, Dam37, Dau35, Dau36, Dau39, Dav30a, Dav32, Dav34a, Dav33b, Dav34b, Dav37a, Dea30, Dea32, Dem36a, Dem37, Dem39, Den37, Der39, Des38, Dew35, Dew36, Dew38, Die36, Dod36, Dod38, Dol39, Dou38a, Dou30, Dou38b, Dow38, Dre39b, Dre39a, Dru36, Dru38, DuB35, Dub31, Dul31, Dul34, Dul36a, Dul36b, Dul37, Dun39, Dun37, Ede37, Edi31, Edw39, Eld35, Eld36a, Eld36b, Ell37, Els36, Emm31, Eng38, Eps33, Eps34b, Eps34a, Eps34c, Eps36, Eva33, Eze32a, Eze37, Fab36, Fal38]. Book [Fal37, Fel30, Fel37, Fel36, Fel39, Fer31a, Fer32a, Fer32b, Fer34, Fer39, Fic35, Fis32, Fis37a, Fis37b, Fis36c, Fle35, Fle38a, Fle38b, Fol35, For39, Fos35b, Fos35a, Fos39a, Fra30, Fre38, Fri30a, Fri39a, Fri39c, Fri36a, Fri36b, Fri37b, Fri38, Fri39d, Fri39e, Fry36, Gay31a, Gay31b, Gay31c, Geh31, Geh37, Geo37, Ger39, Gev35, Gid33, Gid39, Gil33a, Gil34, Gil35, Gil36a, Gil36b, Gil37, Giv31a, Gla30, Glo31, Gok36, Goo34, Gra38, Gre36a, Gre37, Gre36b, Gre38, Gre39, Gri36, Gri35, Had38, Hag37, Hag39, Hai32, Hai33, Hal33, Hal37, Hal38, Ham33a, Ham35, Ham36, Han30, Han33, Han35, Han37, Har37a, Har37b, Har30, Har32a, Har34a, Har39a, Har32b, Har36b, Har39b, Har39d, Har37c, Hau37, Hay36a, Hed36, Hei39, Hen32, Hen39, Her30, Hew30]. Book [Hil38, Hin39, Hin32, His32, Hob39, Hof34a, Hof34b, Hog31a, Hog31b, Hog34a, Hol32, Hol36, Hom34, Hoo33, Hoo37, Hoo38a, Hoo38b, Hoo39, Hor35a, Hot30a, Hot32a, Hot32b, Hot33b, Hot33a, Hot35a, Hot35b, Hot35c, Hot35d, Hot37a, Hot37b, Hot38, Hot39, Hou39, How36a, How36b, How37, Hoy39, Hub37, Hur34a, Ing39, Inm39, Isa34, Isa36, Isa38, Ise36, Jam35, JP36, Jen34, Jer31, Jer32, Jer36, Joh38a, Joh39, Jon37, Joy30a, Joy33, Kah39, Kel38, Kel33, Kel34, Kel36, Kes36, Kil38, Kin39a, Kin31a, Kin32a, Kin33c, Kin39b, Kis34a, Kis36, Kni37, Kni38, Kni39, Koe36, Kop30, Kop31, Kop32a, Kre31, Kre32, Kro34, Kuc31a, Kul34, Kul38, Kuz30a, Kuz30b, Kuz31, Kuz32a, Kuz33, Kuz35b, Kuz38a, Kyr30, Kyr33, Lan33, Lan35, Lat30, Lat34, Lea38, Lea39]. Book [Led36, Leo36b, Leo37, Ler38, Lev30, Lev37, Lew38b, Lew38a, Lew39, Lin32, Lin39, Liv39, Liv38, Loc36, Lor39, Lou39, Lov36, Lub30, Lun34, Lun35, Lun36a, Lun37, Lun38, Lun39, Lun36b, Lut33, Lut37, Lut39, Lyo31, Lyo36, Mac33, Mac38, Mac39a, Mac39b, Mad38, Mal35, Mar30, Mar33, Mar34a, Mar35a, Mar35b, Mar39a, Mar39b, Mar39c, Mas33, Mas35, Mas36, Mas37, Mas38, McC36a, McC36b, McD38, McD39a, McD39b, McI36, McI38a, McK36, McL36, McL38, McL39, Mea31, Mea33, Mea37, Men39a, Men30, Men32, Men35, Mer33, Mey38, Mil38, Mil30, Min30, Min33, Min35, Mit30, Mit32, Mit36, Mon38, Moo37, Mor39, Mor36a, Mor36b, Mud37, Mud39, Mus33, Mye39, Nad39, Nag39, Nat35, Nat37, Nat39, Nel39, New38, Ney39, Nic39, Nys34, Orc30]. Book [Orc31a, Orc31b, Orc33a, Orc33b, Orc34, Orc35, Orc36, Orc37, Ost31, Pal38a, Pal39a, Pal36, Pal38b, Pal39c, Pal39d, Par37a, Par39, Par38, Par37b, Pat30, Pat34a, Pat34b, Pat37, Pea38, Pea36a, Pea36d, Pec31, Pec33, Pen31, Pen32a, Per37, Per31, Pet33, Pet39, Phe31, Phe37a, Pri35, Pro39a, Pro39b, Puf39, Ree33, Ree35, Ree37a, Ree39, Ric38b, Rid37, Rie32, Rie30, Rie31a, Rie31b, Rie37, Rig33a, Rig39, Ril38, Rob33, Rob38, Rob39, Rod32, Roe30, Roe38, Rog34, Rol30a, Rol30b, Rol31, Roo36a, Roo36b, Rub30, Rub33, San36b, San39, San31, Sch39a, Sch39b, Sch38a, Sch38b, Sch39c, Sch31, Sch36a, Sch39e, Sch36b, Sch36c, Sch39f, Sco37, Sco39, Sco38, Sea36, Sel38, Sel32, Sel33, She39, Shi37a, Shi37b, Shi38, Shl39, Sho33, Sho35, Sho36]. Book [Shr39, Sie39, Sim37, Sim38, Slo30, Slo31, Sma34, Smi32, Smi33, Smi35a, Smi35b, Smi36, Smi37b, Smi37a, Smi39, Sne39, Sou30, Sou34, Sou36, Spa35, Spe30, Spe31, Spe39, Spi39, Ste34a, Ste34b, Ste35d, Ste36b, Ste36c, Ste37b, Ste39c, Sto32, Sto31, Sto37, Sto38, Sto39b, Str37, Str39, Stu37a, Swa36, Swa39, Swe35, Swi38, Tae38, Tay30a, Tay31, Tay35, Tay38, Tea38a, Tea38b, Tea38c, Tee37a, Tee37b, Tho32a, Tho32b, Tho35a, Tho32c, Tho35b, Tho32d, Tho39a, Tho30, Tho32e, Tho34a, Tho38b, Tho39b, Tho39c, Thr39, Tim30, Tim31, Tim32a, Tim32b, Tim33, Tim35, Tim36, Tim37, Tip36, Tip37, Tip39a, Too35, Too36, Too37, Tre38, Tro39, Tru39, Tuc36a, Tuc37, Tuf37, Tul33, Ush33, Vai36, Van35, Van39, Vee36, Vee39, Via37]. Book [Vin30, Wad34, Wai39, Wai36, Wal39a, Wal39b, Wal38, Wal37a, Wal37b, War37, War38a, War34b, War30, War36, Wat30a, Wat38a, Wat36, Wat38b, Wat39a, Wat39b, Web38, Wei32a, Wei32b, Wei39b, Wel36a, Wel39a, Wel35, Wel36b, Wel39b, Wel38b, Wel39c, Whe36a, Whi35, Whi37, Whi39, Wil36a, Wil36b, Wil37a, Wil38b, Wil38c, Wil33a, Wil32b, Wil36c, Wil36d, Wil37c, Wil38d, Wil39b, Wil32c, Wil33b, Wil33c, Wil34g, Wil36e, Wil30c, Wil35b, Wil39c, Win36, Win39, Wit36, WS39, Wol35a, Wol38, Wol35c, Woo37, Wor37, Wor39, Wri31, Wri34b, Wri37, Wri38, Wri39, Ynt30a, Ynt30b, Ynt32, Ynt35a, Ynt35b, Ynt37, Yod33, Yod34, You32, You36b, You36a, Zap31, Zel36, Zim39, dC34, dC37a, dR37, vB39, vH30a, vH30b, vH32, vH34a, vH34b, vH38a, vS31]. Boole [Wil34h]. Borrowed [Rob34, Rob36]. Britain [Pea36c, Tol34]. British [Hot30b]. Broadened [Fox36a]. Budgetary [Han34b]. Building [FW37, Rig33b]. Bull [Kin31g]. Burdened [Kin34a]. Bureau [Che34, CD37, Ano33v, Aus35, D.36, H.30a, H.30b, H.30c, H.31a, H.31b, H.31c, H.32a, H.32b, H.32c, H.33a, H.33b, H.33c, H.34d, H.34b, H.34c, H.34e, H.35a, H.35d, H.35b, H.35c, H.36a, H.36b, H.36c, H.37a, H.37b, H.37c, H.37d, H.38a, H.38b, H.38c, Lin37b, Per36, Sho37]. Bureaus [Mye34]. Burks [Lor35]. Business [Ano34a, Cam31, D.36, DuB36, Gos37, Gos39, H.34d, Han34a, Har34b, HCW^{+}34, Kin30c, Kin34c, MF35, Sho37, Sma33, Smi31, Tip39b, TBSM38, ZDM^{+}34, Hot33a]. By-Laws [Ano30b].

C. [Man45, Ros38]. Cadillac [Ano39n]. Calculating [Li34, Pis32, Sch33a, Whe36b]. Calculation [CD37]. Calculations [MH39]. Calendar [Ano34a, Kuz35a]. California [Stu37b]. Calumet [Tay30b]. Can [Bad31, Kin39c]. Canada [Gol34]. Canadian [Tur39]. Capita [Tho37b]. Capital [Kuz38b, Rob34, Rob36, War38b, Woo38]. Card [Rea34, Vic38, Vic39]. Care [Ano32c]. Case [CJS33]. Cases [Sto33]. Catalytic [Mue34]. Causation [Cou30]. Cause [Dun36, Har34b]. Caused [Kin32e]. Causes [HH31]. Census [Coa31b, D.36, Dav33a, Edg31, Fry30, GB34, Gos37, Gos39, H.31c, H.34d, H.34e, H.35d, Hum37, Joy30b, Joy34a, Kol35, Kol37, McF32, McL32, Tho36a, Tru31, Tru33, Tru35, Ano33v, Aus35, D.36, H.30a, H.30b, H.30c, H.31a, H.31b, H.31c, H.32a, H.32b, H.32c, H.33a, H.33b, H.33c, H.34d, H.34b, H.34c, H.34e, H.35a, H.35d, H.35b, H.35c, H.36a, H.36b, H.36c, H.37a, H.37b, H.37c, H.37d, H.38a, H.38b, H.38c, Lin37b, Sch35, Sho37, SN30, Tho37a]. Censuses [Ano37n, Wol32]. Centenary [Wal34]. Center [Ano30c, Eel30]. Central [Ano33d, Ano34d, Ano35b, Ano36w, Dic35, Kis32, MF35]. Certain [Dor39, Fri30b, GB34, Mal36, SS33, Sub34]. Challenge [Wil34h]. Chance [BMH35]. Change [Cha35b, Kin33e, Ree37b, Sas30, Tru38]. Changed [Bog38b]. Changes [AC34, Cop36, Den36, Fis36a, Fis36b, Kin31d, Lom38, SP34, TT38, Wil34f, Woo38]. Changing [Hom33, Wei38]. Chapter [Ano35c, Ano35d, Ano35e, Ano36a, Ano36b, Ano36c, Ano36d, Ano37c, Ano37d, Ano37e, Ano37f, Ano38b, Ano38c, Ano38d, Ano38e, Ano39b, Ano39c, Ano39d, Ano39e]. Character [Bak30]. Characteristic [Ano36i, MTDD35]. Characteristics [Eze32b, Wai32b]. Charles [Phe39]. Charts [Ano38n, Bur37a, Kuc31b, vH31]. Check [MH39]. Checking [Sch35]. Checks [McI37]. Chi [Ber38, Cam38b, Dem34, Dem38]. Chi-Square [Ber38, Cam38b]. Chi-Test [Dem34]. Chicago [Ano37s, GS36, HD30, Tay30b]. Children [Sch35, Gir39]. China [Wil30b]. Chinese [Wis36]. Chronology [Wil34e]. Churchill [Bel36]. Circles [Cro32b, Hun32]. Cities [Kop32b, Mag33, Ogb35, Tho37b]. City [Ano38t, ZDM^{+}34, Ano36j, Ano36y, Hic39, Wil34d]. Class [SN30]. Classes [Kis32, Sch33a, Sog33, Yat34]. Classification [Ano36i, Eps35, MTDD35, Sne34]. Classifications [Yat34]. Clearing [Men39b]. Climate [CC35]. Cluster [FM31, FM32]. Coal [Ror35]. Code [Che34, Com34, Dun32, Rea34]. Coded [RT33]. Codes [Sel34]. Coefficient [Bak30, COP32a, Cow32, EM34a, For32, GB34, Gri32, HG30, McC34, Mos39, Pea31, Pea32, Sch33b, Sto34a, Wol35b]. Coefficients [Gri30, Li34, Ric31, Wat30b]. Collect [Cop34, Lub34]. Collected [CS39]. Collection [HCW^{+}34, Hur34b, Mit39a, Mye34, Phe34b]. Collegiate [Ano34e]. Combination [Bat35]. Combined [Sto34a]. Comment [Kuz38b, vS30]. Comments [Ano31d, Rie31c]. Commercial [Bar38, HCW^{+}34]. Committee [Ano31a, Ano33a, Ano33e, Ano33f, Ebl38b, Giv34a, Giv34b, Giv34c, Ric33a, Ste35c, Ste36a, Ste37a, Ste38a, Ste39a]. Committees [Ano33g, Ano34f, Ano34x]. Commodities [Hic39]. Commodity [Fis30, Tur39, Woo32a]. Commodore [Ano36j, Ano36y]. Communities [Dor39, Ebe38, Kis32]. Company [Ayr38]. Comparability [Jan38, Ree37b, Wur36]. Comparable [Hor31]. Comparative [Leo32a, Mag33]. Comparison [Fra31]. Comparisons [Cro32b]. Compensation [Big37]. Compilation [Bec38a]. Complete [Kin35b]. Completed [Bur33, Lor35, WH35]. Completeness [Whe34a]. Components [Gir36, War34a]. Composition [Cha31, Lor35, MK37, RG30]. Computation [Sib30, Wis34]. Computing [vH31]. Concentration [Mah30, Zab30]. Concept [Kin39d, Ric37, RvS39a, RvS39b]. Conception [Eel30]. Concepts [Fer36, JW39, Web39]. Concerning [Wil34b]. Conclusion [Sub34]. Condition [Gin30]. Conditions [Bog38b, Fri30b, Han34a]. Conference [Ano36e, Cha37, Wil38a]. Confidence [Ric37]. Confluence [Men39b]. Consistency [San37]. Consolidating [Kin36a]. Consolidation [Dav39, Sho37]. Constants [Hor32, Hor35b, Wau35]. Constitution [Ano30b]. Constructing [BT32]. Construction [Ano38n, Bur37a, Cha35b, Cla34a, FW37, GH37, Gil33b, Hud39, Per34, Tay34]. Constructive [Ano32b]. Consumer [Art39, Bar38, Fos38, Geo39, Hum39, Kin31b, Leo32b, Leo35, Nug38, SP37, Tip39b]. Consumers [Kin34a]. Consumption [KSF36, MK37, Ror35]. Continental [Cha31]. Contingency [PHT^{+}30]. Contrasted [Zub39]. Contribute [Bur32b]. Contribution [Bow35a, Ney38a, Tol34]. Control [Har34b, Hay36b, Kin35b, Ray36]. Convention [Nix38]. Convertibility [Pal39b]. Cooperation [Che34, Hur34b, Phe34b]. Cooperative [Ano36f, Ano36g]. Copper [Leo36a]. Corporate [Fab35, Rob34, Rob36]. Corporation [Slo37]. Corporations [Gou36]. Correcting [Gri30]. Correction [Ano35f, Ano36h, Ano36i, Ano36j, Ano37h, Ano37g, Ano38f, Fri39b, FM32, Hor35b, Mos38, Vic38]. Corrections [EM34a, EM34b]. Correlation [Ano32d, Ano36h, Bak30, Bea30b, COP32a, Cow32, Dav30b, DM37, Eze32b, FM31, FM32, GB34, Gri30, Gri32, Ham33b, HG30, Hun32, Imm37, Li34, McI37, Mor37, Nep34, Pea31, Pea32, Ric31, San34, Sto34a, Sto34c, Sto36, Stu34, Wai32a, Wai32b, WD32, Wal39d, Wat30b, Wat33, Wis34, Won37]. Correlational [GS36]. Corresponding [Lea31]. Cost [Ach34b, Cha35b, Dea37, Hog31c, Hog34b, Kin33d, Tay34, Wil34f, Wri34a]. Cost-of-Living [Hog34b, Wri34a]. Costs [Ano32c, Ayl37, GH37, Mag33, Sch33a]. Cotterman [Man45]. Cotton [CB36]. Council [Ano30a, Ano31b, Ano32e, Ano33b]. Count [BMH35]. Counting [Ben37b, BA37a, BA37b]. Court [Wal36]. Covariance [Bai31]. Coverage [Wel38a]. Created [Kin34a]. Credit [Fos38]. Credits [Rog32]. Crime [Ogb35, Wat31]. Criminal [Mar34b]. Criteria [DS33, Wil35a]. Cubes [Cro32b]. Curb [Kuv30]. Currency [Kin35b]. Current [Llo38, MF35, Sho37]. Curve [Ber36b, Dem34, Dem38, ML31, Mue34, Sch30, Sma33, Wal34]. Curves [BT32, Dea37, Li34, Ste31, Ste32, Sti39, vH33b]. Curvilinear [Li34]. Cycle [Cam31, Har34b, Kin30c, Tip39b]. Cycles [Rig33b]. Cyclical [Kin30e].

D [Phe37b]. Data [Bec38a, CS39, Fry30, Hin38, Hom33, Hum37, Joy30b, Joy34a, Kin30d, Kol35, Kol37, McL32, Mit39a, Mue34, Pis34, San37, Sas30, Ste31, Ste34c, Sto33, Sto34a, Stu37b, T.34, Tru31, Wal39d]. Death [Dow31, Dun36, Kop32b, Lin37a, Wur36]. Deaths [Dor32, Kuc30, Lin37a, Lot30]. Debt [Sny34]. December [Ano38t, Asc37, Ano30e, Ano31e, Ano32g, Ano33h, Ano34l]. Decennial [H.31c]. Decision [Bra36a]. Decline [Rey39, Whe34b]. Declining [Woo32b]. Decrease [Dor39]. Demand [BT32, Cur38, Geo39, Kin39d, Mos39, RvS39a, RvS39b, Sch33b, Sti39]. Denmark [Nyb37]. Density [Wat31]. Denver [Car34]. Dependent [Cow32, HG30, Sch35]. Deposit [Fox36b, Gol38]. Deposits [Cur38]. Depression [CP34, Rog32, SP34, Tho33, Wil38e]. Derivation [Ham33b, Mue34]. Derived [Joy30b, Joy34a, Kol35, Kol37]. Deriving [Wol35b]. Design [Ano38n]. Determination [Ano36i, EM34a, McC34, MTDD35, Mos37, Mos38, Wat30b, Wol35b]. Determining [Wau35, Zub39]. Detroit [Ano39n, Hab32, Mag33]. Developed [Ano31d, Rie31c]. Developing [Pal39b]. Development [Ney35, Sis36, Ste35a]. Developments [Kin31c, Kin33h]. Deviation [HG30, Mav32, Wil38f, vH33c]. Deviations [Imm37]. Diagram [Li34]. Dictionary [Ano36v]. Did [B.39, Bra39, Wei39a]. Difference [Pee37, vH33b, vS30, Wor34]. Differences [Spi34, Zub39]. Different [Yat34]. Differential [CJ32, DS33, SN30]. Differentials [Jaf39]. Difficulties [Ber38]. Digit [DM37]. Dinner [Kin34b]. Discrete [Ano38f, Mue38]. Discussion [Ben37b, BA37b, Cla34a, Cop34, Cop36, Dav33a, Fet37, Har39c, JW39, Kyr37]. Discussions [Cha37]. Disease [Kin31f, Mal36]. Dispersion [Bea34]. Displacement [Wei32c]. Disproportionate [Bra33a, Sne34]. Disputes [Dou32]. Dissimilar [Hor31]. Distinctive [Fer36]. Distortion [Hog31c]. Distributed [Nep34]. Distribution [Ano37h, Ano37g, AC34, Bur33, COP32a, Cru35b, Cur38, Kin30d, Lom38, Lor35, Lot36, ML31, McF32, Mue36, Ric37, Wal36, Ynt33]. Distributions [Ano38f, Asc37, Bak30, Hor31, Lea31, Mue38]. Divergent [Tay32]. Dividends [Dic38]. Division [MF35, Sho37]. Divorce [YD32]. Domestic [HD30, Llo38]. Dr. [Kuc30, Kuz38b, Lot30, War38b]. Durable [Geo39, Hum39, Leo35, Tip39b]. Duration [YD32]. During [Kre34, SP34]. Duty [Bur37b]. Dynamics [Kin39d, RvS39a, RvS39b].

Early [Bow35a]. Earner [Hog34b]. Earning [Car38]. Earnings [Fab35, Rob34, Rob36, Tho38a]. East [Kis32]. Easter [Hom33, Pis34]. Economic [Ayr38, Bow36b, Cla37, CP34, Cop34, Cur38, Ebe38, Edw33, EM34b, Hom33, Kin33e, Kuc31b, Lew32, Lub34, SS33, Stu37b, SP34, Wat30b, Wei38]. Economics [Fis33, Spr34]. Edgar [Kin36b]. Editor [Ano35l, Ano30c]. Editors [Ano36p]. Effect [Ano36h, Dor32, Dun36, Kin35a, Kuv36, Lin37a, Mah30, Ric31, Sto36, Tho33]. Effective [SS33]. Effectiveness [Vol39]. Effects [DF37, Fet37, GB34]. Efficiency [Wei32c]. Efficient [Alt37, Big37]. Effort [Wil30b]. Eighteenth [Wil30a, Wil31]. Eighth [Ano37s]. Elasticity [Kin39d, Mos39, RvS39a, RvS39b, Sch33b]. Element [SS33]. Elementary [Har39c, Kel39, Mos37, Mos38]. Elimination [Kin32g, Wol35b, Won37]. Ellipses [Hun32]. Emergency [Ros34]. Emigration [Gol34]. Empirical [CJS33, Whe36b]. Employees [Cam31, Wil34f]. Employer [Jar37]. Employment [Cop36, Fis36a, Fis36b, Hab32, Hil31, Hog30, Joy34b, Kin30e, Mah30, Sch36b, Ste35b, Tay30b, Wil38a]. Employment-Unemployment [Hog30]. Encountered [Ber38]. Encyclopaedia [Cad35]. End [YD32]. Engel [Hei31]. Engineering [Dav37b, She31]. Entering [Wil34a]. Enumeration [Coa31b, Coc39, Dav33a]. Equal [Ano32d, Hun32]. Equal-Frequency [Ano32d, Hun32]. Equation [Smi34]. Equations [Ham33b]. Equipped [DM37]. Equity [Lom38]. Equivalent [Bur38]. Errata [Ano31d, Ano32d, Ric30a]. Error [Hog30, Mos39, Sch30, Sch33b]. Errors [Mos37, Mos38, NT36, Ste34c]. Erythrocyte [BMH35]. Established [Kin35b]. Estimate [Wil30b]. Estimated [BMH35]. Estimates [Ebe38, Kre34, Lor35, Tay37]. Estimating [Bur33, FW37, Woo34]. European [Mag33]. Evaluating [Ano36h, Eze34, Hor32, Hor35b, Sto36]. Evaluation [Dun36]. Events [Zub39]. Examination [Lot36]. Example [Sto34a]. Exchange [Kuv30]. Exchanges [Kin34d]. Executive [Hay36b, Tit36]. Exercise [Kin35b]. Expansion [Art39, Rey39]. Expectancy [Wil38f]. Expected [Sne34, Sto33]. Expenditure [Kap38]. Expenditures [Bec38a, Bra38, Kyr37, Wil37d, Wri34a]. Expenses [Tho38a]. Experiment [CF39, Giv34c]. Experiments [COP32a, Imm37, Kin35a, Ric30a, Ric30b]. Exponential [Smi34]. Expression [Sas30]. Extension [Wis31]. External [Wal39c].

Factor [Edw36, Wol35b, Wub39]. Factorial [GS36]. Factories [Hab32].

Factors [Cha35b, Cla34b, Lew32, Mit39b, Ogb35, Wil34a]. Factory [Cop36, Fis36a, Fis36b, Tit36]. Families [Kap38, KSF36, MK37, Sto34b, WH35]. Family [Bad33, Bra38, Kyr37, SP34, Tru31, Wil37d, Wri34a]. Farmer [Eze34]. Farms [Ben37b, BA37a, BA37b]. Federal [Sch35, Ano37i, Gol36, Hur34b, Kin34d, Kin35b, Phe34b, Ros34, Wil34f]. Fellowships [Ano32e]. Ferris [Phe39]. Fertility [Kis32, Sto34b, SN30, TT38]. Fertilizer [Bra36b]. Fiducial [Ric37]. Field [Ano32f, Coa31b, CJ32, Imm37, Man34, Per36, Tru38]. Fifteenth [Tru33]. Fifth [Ano34r]. Files [San37]. Finance [Ayr38]. Financial [Dav39, Kin32b, Kin33e]. Financing [Bar38, Nug38, Woo38]. Find [Bad31]. Finding [Bat35]. First [Dav30b]. Fisher [Fri30b, Wis31]. Fit [Ber36b, WHM31]. Fitted [Smi34]. Fitting [Dem34, Dem38, Li34, ML31, Ste31, Ste32]. Flexible [Dun35, Kin39c]. Fluctuations [BMH35, Hab32, HH31, Wai38]. Food [Sti34]. Force [Edw36]. Forecast [Sch30]. Forecaster [Bur37b]. Forecasting [Kin32c, Kin34f, Smi31, Tru38]. Forecasts [Kin30a, Woo32a]. Foreign [Mea35b]. Forest [MG34]. Forestry [Ayl37]. Form [Fri30b]. Formation [Kuz38b, War38b]. Formula [Bat35, PS30, Wis31]. Four [Kre34]. Fourteen [Mag33]. Fraternities [Bur33, Lor35]. Freight [Dun35]. Frequencies [Bra33a, Zub39]. Frequency [Ano32d, Ano38f, Hun32, Lea31, ML31, Mue38, Pee37, Ric37]. Frisch [Sub34]. Front [Ano30d, Ano34g, Ano34h, Ano34i, Ano34j, Ano34k, Ano35g, Ano35h, Ano35i, Ano35j, Ano36k, Ano36l, Ano36m, Ano36n, Ano37j, Ano37k, Ano37l, Ano37m, Ano38g, Ano38h, Ano38i, Ano38j, Ano39f, Ano39g, Ano39h, Ano39i]. Fry [Ros38]. Function [Ayr38, Gol36, Gol38]. Fundamental [Wil34a]. Further [Cam38b, COP32a, Eze32b, Joh36, Pea32, RvS39b, Wai32a]. Future [Gin30, Whe36b].

Gainful [Edw33, EM34b, Smi35c]. Gains [Kin36a]. Garments [Gir39]. General [Ano36f, Eps35, Hor32, Hor35b, PS30, Sch36b, San36a]. General-Purpose [Eps35]. Generations [CS39]. Genetic [San34]. Geographic [Cha34a, Lot36]. Geometric [Dun32, Mor37, vH30c]. Geometrically [RT33]. George [Ste34d]. German [Kop32b]. Germany [TT38, Tol34]. Gets [Gos39]. Gettemy [Phe39]. Given [ML31]. Gold [Bur38, Den36, Rog32]. Goodness [Ber36b, WHM31]. Goods [Art39, Geo39, Hum39, KSF36, Leo32b, Leo35, Rey39, Tip39b]. Gosset [Ney38b]. Government [Ano33e, Ano35a, Ano38w, Che34, DuB36, Gal38, Giv34a, Giv34b, Giv34c, Kin33d, Ric33a, Ril37, Kin34d]. Graduated [Ber36b, Sch33a]. Grants [Ano32e]. Graphic [Ano36o, Cro32b, Eze32b, Mav32, Sib30, Spu37, Sto39a, Wai32a, Wai32b]. Great [Pea36c, Tol34]. Grid [ML31]. Gross [Rob34, Rob36, War34a]. Group [Cow32]. Grouping [Edw33, EM34b, GB34]. Groups [Kin31g, Mal36, Zub39]. Growers [CB36]. Growth [Joh35, Joh36, TBSM38, Val30]. Guidance [Mit39c].

H [Man45]. Haddon [Ano38t]. Hall [Ano38t]. Handbook [Stu37b]. Handling [SS33]. Harmonic [Fer31b]. Harry [Ebl38a]. Health [Bur32b, Ste35b]. Height [Bur37a]. Hemocytometer [BMH35]. High [Ric31]. Highlights [Gos37]. Highway [Sch33a]. Hill [Gol39]. Honor [Kin34b]. Horace [Hot33a]. Hotel [Ano36j, Ano36y, Ano37s, Ano39n]. Hour [Bea34, Mit39c]. Hours [Nix38]. Household [CSS39]. Households [WW38]. Housing [Ach34b]. Human [Ney38a]. Hundredth [Ano39n]. Hypotheses [Wil35a]. Hypothesis [NT36].

Ideal [PS30]. II [Wai32a]. Illustration [HG30]. Implications [Giv35]. Implicit [Fri37a, Fri39b]. Importance [Val30]. Improve [vH31]. Improved [Ano32d, Ano35v, Cow37, Hun32]. Improvement [Ril37, Sne36, Wil34b]. Inadequacy [McL32, San34]. Inadequately [Ano36h, Sto36]. Income [Ano36e, Bad31, Bad33, Cop35, Cru35b, Haa36, Kuz38b, Rob34, Rob36, SP34, Tho37b, War38b, Wil34a, Ynt33]. Incomparability [Hog30]. Increase [B.39, Bra39, HP31, Lot36, Sny34, Wei39a, Wil30b, Woo31, Woo32b]. Increases [Wei32c]. Independent [EM34a, McC34, Ric31]. Index [Ano38w, Art37, Cla34a, Cop36, CD37, Eva36, Fer36, Fis36a, Fis36b, Fri30b, Har36a, Hic39, Hog31c, Hog34b, Joy30b, Joy34a, Kol35, Kol37, McI38b, Per34, PS30, Smi31, SP31, T.34, Tur36, Val30, Wat33, vH30c]. Indexes [Cha35b, Fis30, Gar38, Hud39, Joh38b, Kuv30, Leo32a, Leo32b, Leo35, Pis32, Wel30, Wri34a, ZDM^{+}34]. Individual [McL32]. Individuals [Bur33]. Industrial [Dou32, HCW^{+}34, Mil37a, Mit39c, Ney35, Phe34b, Ray36, Sog33, Tho36b]. Industries [Eps35, Fab35, Fab38, McL32, Rom35, Val30]. Industry [Ano37y, Ken36, Mah30, Pea36c, Sar34, Sel34, Wei32c]. Inequality [Ynt33]. Infant [Cha34a, Nyb37]. Inflation [Kin33f]. Inflexibility [Mea35a]. Influence [Ano36i, Fox36b, Lor35, MTDD35, Pis34, Wat31]. Influenced [Cha35b]. Information [Ano30r, Ano31p, Ano32r, Ano33w, Ano34-27, Ano35w, Ano36-30, Ano37x, Ano38-28, Ano39v, Giv34a, Giv34b, Giv34c, Ric33a, Tit36]. Inherit [Cha33b]. Inheritance [CS39]. Inquiry [Mag33]. Insecticides [CSS39]. Instability [Cla34b]. Instead [Hun32]. Institute [Ano35k, Ano36q, Ano38k, Ano38l, Bra36a, DW35, Wil30a, Wil34d, Kin34b, Wil34c, Wil37b]. Institutes [Com34]. Institution [Cha30b]. Institutional [Vol39]. Insurance [Bow35a, Bow36a, Fox36b, Kin33h, Ror36, Tol34]. Integrated [DuB36]. Intelligence [Ano32f]. Intensity [vH33c, vH38b]. Intercorrelation [Ric31]. Interest [Dic38, Sch36b, Sma33]. Interests [Dem36b]. Internal [Tho37a]. International [Ano38l, Dic38, DW35, Mag33, Nix38, Rog32, Wil30a, Wil34c, Kin34b, Wil34d, Wil37b]. Interpolation [Sto39a, Wis30, Woo32b]. Interpretation [Ber38, Cop35, Cru33, Mit39a, Wal39c]. Interpretations [Cam38b, Ste34c]. Intrinsic [Lot36]. Inventories [Ell38]. Inventory [Gre34, Mit39a]. Investment [Slo37]. Invited [Ano35v]. Involved [CSS39, Ham33b]. Involving [Wil35a]. Issues [Bog38b]. Italy [Zab30].

January [Asc37]. Jerome [Ebl38a]. John [Sch36b]. Joint [Cou30, Mye34]. Joseph [Gol39]. Journal [Ano38m]. Judicial [Mar34b]. June [Ano30f, Ano31f, Ano32h, Ano33i, Ano34m].

Karl [Cam33, Pea36b]. Kelley [Hol37]. Keynes [Sch36b]. Kind [NT36, Tit36].

L. [Man45]. Labor [CD37, Edw36, HD30, Mye34, Per36, Wei38, Wub39]. Laboratory [BMH35]. Lag [vH33b]. Large [Kop32b, Ste39b]. Large-Scale [Ste39b]. Last [Kre34]. Law [Hei31]. Laws [Ano30b]. Lead [Kin31g]. Least [Mos39]. Least-Squares [Mos39]. Legal [Bra36a]. Lending [Gol36, Kin34e]. Leroy [Phe37b]. Letter [Gre32, Har34b, Pen32b, Ric33b, Wil34b]. Letters [Ano35l, Ano36p, SHR^{+}34]. Level [Kin32d, Kin32e]. Leviathan [Mil35]. Life [Bow35a, Wil38f]. Light [Han34a]. Limitations [Hei31, Hud39, Sti39]. Limits [Ric37]. Line [Joh35, Joh36]. Linear [Li34]. Liquid [CSS39]. Liquidation [Kin32b]. List [Pal39b]. Living [Bra38, Hog31c, Hog34b, Mag33, TK33, Wil34f, Wri34a]. Loads [CJS33]. Local [Ebe38]. Localities [Tay32]. Log [Ayl37]. Logarithmic [Ste31]. London [Tuc36b, Wil34c]. Long [Sch39d, Sny34]. Long-Term [Sny34]. Long-Time [Sch39d]. Lotka [Kuc30, Lot30]. Low [Ach34b]. Luther [Ros38].

Machine [Fra31]. Machine-Tool [Fra31]. Machines [Kin33b]. Made [Lin37b]. Mail [Mil37b]. Mail-Order [Mil37b]. Mailed [Mit39b]. Making [Ano37y, Cha30b, Cla34b, Fry30, Kin39c]. Malcolm [Bel36]. Man [Bea34]. Man-Hour [Bea34]. Management [Ken36, Mit39c]. Manual [Ano38n]. Manufacturer [Ell38]. Manufactures [Joy30b, Joy34a, Kol35, Sho37]. Manufacturing [Eps35, Fab35, Kol37]. Map [Ano32d, Hun32, vH38b]. Maps [Kuc31b]. March [Ano30g, Ano32i, Ano31g, Ano33j, Ano34n]. Marginal [Bak30]. Market [Kin31g]. Marketing [Bog38b, Ken36]. Markets [Gou36, Hic39, Tur39]. Marriage [Kop32b, Tay32]. Marriages [YD32]. Married [SN30]. Material [GB34]. Materials [Dod37]. Mathematical [Ano34e, Ano35k, Ano36q, Ano38a, Ano38k, Cam33, Rom35]. Mathematics [Wil30d]. Matter [Ano30d, Ano31c, Ano32a, Ano33c, Ano34b, Ano34c, Ano34g, Ano34h, Ano34i, Ano34j, Ano34k, Ano35g, Ano35h, Ano35i, Ano35j, Ano36k, Ano36l, Ano36m, Ano36n, Ano37b, Ano37j, Ano37k, Ano37l, Ano37m, Ano38g, Ano38h, Ano38i, Ano38j, Ano39a, Ano39f, Ano39g, Ano39h, Ano39i]. Maynard [Sch36b]. Mean [Fer31b, Smi34]. Meaning [NMDW36]. Means [Ber36b, Imm37]. Measurable [Ano36i, MTDD35]. Measure [Wil31, vH33b, vH33c]. Measured [Ano36h, Sto36]. Measurement [DF37, Fab38, Fer37, Fet37, Kin33d, Kuz32b, Lew37, Wub39]. Measures [Giv31b, Hil31, Lot36, TT38, Wei38, Ynt33]. Measuring [Cou30, HG30, Kyr37, Mag39, Spu37, TK33, Wel38a, Wil34f, Wil37d]. Meat [Sar34]. Mechanical [Dun32, vH38b]. Mechanics [Tru35]. Medical [Ano32c]. Mediocrity [Hot33a]. Meet [Fri30b]. Meeting [Ano36j, Ano36q, Ano36y, Ano37s, Ano38t, Ano39n, Kin30d, Kin30e, Kin34b, Kin34c, Ror30, Ste38b, Wil34c, Wil34d, Wil37b, Ano34r]. Member [Ano38o]. Members [Ano30e, Ano30g, Ano30f, Ano30i, Ano30h, Ano31e, Ano31g, Ano31f, Ano31h, Ano32g, Ano32i, Ano32h, Ano32j, Ano33h, Ano33j, Ano33i, Ano33k, Ano34l, Ano34n, Ano34m, Ano34o, Ano35n, Ano35o, Ano35p, Ano36r, Ano36s, Ano36t, Ano36u, Ano37o, Ano37p, Ano37q, Ano37r, Ano38p, Ano38q, Ano38r, Ano39j, Ano39k, Ano39l, Ano39m, Kin34b, Kuz35a]. Menaced [Kin33b]. Mendelian [CS39]. Mental [Mal36]. Merchandise [Kin30d]. Mergers [Kin30b]. Metals [Wil32a]. Meter [vH33a]. Method [Bea30b, Bur33, CD37, Eze32b, Hol37, Hor32, Hor35b, Kin30a, Li34, Lor35, Mag39, Pea36c, Pis32, Rid35, Ror31, San34, SS33, She31, Smi34, Sne34, Spu37, Sto39a, Sto33, Vee30, Wai32a, Wai32b, Wau35, Whe36b, Wol35b, vH31, vS30]. Methods [Ano37n, Dod37, Kem34, Kin31b, Kin32c, Kin32f, Kin33a, Kin34f, Kyr37, Pet37, Ric38a, Ric30a, Ric30b, SP37, Ste32, Tay37, Wil37d]. Mexican [Tay32]. Mexicans [Tay30b]. Mexico [Wil34d]. Migration [Lew32, Tho37a]. Milton [Fos39b]. Minimum [Ano30c, Big37]. Minneapolis [CJS33]. Minutes [Kin34c]. Miscellaneous [Ano30j, Ano30k, Ano30l, Ano30m, Ano31i, Ano31j, Ano31k, Ano32k, Ano32l, Ano33l, Ano33m, Ano33n, Ano33o, Ano34p, Ano34q, Ano35m, DC31, H.34a, PLE^{+}32, Phe32, Phe34a]. Mistaken [Eel30]. Misuse [Coh38]. Mixtures [Ano38f, Mue38]. Model [WD32]. Modern [Kin31c]. Moment [Bak30, Dav30b]. Monday [Ano38t]. Monetary [Kin35a, Kin35b, Mea35a]. Money [Kin35a, Sma33, Sch36b]. Monopolies [Kin34a]. Month [Ano34a]. Monthly [Fis30, Hic39]. Mortality [Cha34a, Dor32, DS33, Dun36, Jan38, Lin37b, Nyb37]. Motor [Sch33a]. Movements [Mil37b]. Mr. [Fri34, Ney38b]. Multiple [Dun32, Gri30, Ham33b, Hor32, Hor35b, Rea34, Sne34, Wau35, Won37, Yat34]. Museum [Kuc31b].

N.R.A. [Kuv36]. Natality [Lin37b]. Nation [Edw36]. National [Ano36e, Bra36b, Cop35, Cru35b, Kin33g, Kuz38b, War34a, War38b]. Native [SN30]. Natural [Dor39, HP31, Lot36]. Nature [CD36, Fer31b]. Necessary [Fri30b]. Needs [Tit36]. Neglected [Sch35]. Negro [Edw36, HP31, Lew32, Woo31]. Net [Gol34, Jaf39, Lot36, Wat30b, Wol35b]. Netherlands [Met36, Met38]. News [Ano35s, Ano35t, Ano35u, Ano36z, Ano36-27, Ano36-28, Ano36-29, Ano37t, Ano37u, Ano37v, Ano37w, Ano38x, Ano38y, Ano38z, Ano38-27, Ano39s, Ano39t, Ano39u, Han39]. Next [Kin31g]. Ninety [Ano36j, Ano36y, Ano37s, Ano38t, Ano34r]. Ninety-Eighth [Ano37s]. Ninety-Fifth [Ano34r]. Ninety-Ninth [Ano38t]. Ninety-Seventh [Ano36j, Ano36y]. Ninth [Ano38t]. Nominating [Ste35c, Ste36a, Ste37a, Ste38a, Ste39a]. Nomograms [Gri30]. Nomographs [Zub39]. Non [Dor32, Fab35]. Non-Manufacturing [Fab35]. Non-Resident [Dor32]. Nonresident [Lin37a]. Normal [Ano32d, Fra31, Ham33b, Hun32, ML31, Wal34]. Normality [Fri37a, Fri39b]. North [Kis32]. Northeast [Dav33a]. Note [Ano30c, Ano32m, Ano32n, Ano33p, Cow32, Fri34, HW35, Hol37, Man45, Rid31, Stu34, Tay32, Vic39, Wil34e, Wis30, Woo34]. Notes [Ano30j, Ano30k, Ano30l, Ano30m, Ano30o, Ano30p, Ano31i, Ano31j, Ano31k, Ano31l, Ano31m, Ano31n, Ano31o, Ano32k, Ano32l, Ano32o, Ano32p, Ano33l, Ano33m, Ano33n, Ano33o, Ano33q, Ano34p, Ano34q, Ano35m, Ano35q, Ano35s, Ano35t, Ano35u, Ano36x, Ano36z, Ano36-27, Ano36-28, Ano36-29, Ano37t, Ano37u, Ano37v, Ano37w, Ano38x, Ano38y, Ano38z, Ano38-27, Ano39s, Ano39t, Ano39u, DC31, H.34a, Han39, LH35, PLE^{+}32, Phe32, Phe34a, Ror35, T.30, Tur36, Wal39c]. NRA [Che34, Kin34a, Sel34]. Number [Cla34a, Fri30b, McI38b, Per34, Wis31, Woo34]. Numbers [Ano38w, Art37, Eva36, Fer36, Har36a, PS30, Sne34, Yat34, vH30c].

Obituaries [Ano34s]. Obituary [Ano32m, Ano33p, Ano33q, Lot33, Ste34d]. Objective [Cow37]. Observed [Ber36b, Pee37]. Obtaining [Hor31]. Occupation [Smi35c]. Occupational [Ano36v]. Occupations [Fra31, Pal39b]. Official [Dor32, Dow31]. One [Ano39n, CS39, Cow32]. One-Hundredth [Ano39n]. Open [Gre32, Har34b, Pen32b, SHR^{+}34]. Operate [Tit36]. Operating [Ach34b]. Operation [Ray36]. Opinion [Wil31]. Opportunities [Ric34]. Order [Mil37b]. Ordered [Ste34c]. Ordinal [Spi34, Wat33]. Ordinals [Spi34]. Organization [Ano36w, Lan31]. Origin [Kin32g]. Other [Cha35b, Cow32]. Ought [Kin34d, Kin34e]. Our [Kin36a]. Outline [Lin37b]. Outlook [Kin32d]. Output [Fab38].

Packing [Sar34]. Paired [Spi34]. Paper [Man45]. Parabolic [Li34, Ste32]. Parole [Wil39a]. Part [Cow32, Gri32]. Partial [Ano36h, Ham33b, Sto34a, Sto36, WD32]. Partition [Yan33]. Past [Ano30n]. Pattern [Kuz32b]. Patterns [Gir39, Kap38]. Payments [Dic38, Tay37]. Payroll [Wil38a]. Pearson [Pea36b, Cam33]. Peavey [Phe37b]. Pendleton [Ste34d]. Percentage [Ste31]. Performance [Cow37]. Period [Ano38w]. Personal [Ano30o, Ano30p, Ano31l, Ano31m, Ano31n, Ano31o, Ano32o, Ano32p, Ano32n, Ano35q, Ano36x, LH35, T.30, Ynt33]. Personals [Ano32q, Ano33r, Ano33s, Ano33t, Ano33u, Ano34t, Ano34u, Ano34v, Ano34w, Ano35r]. Personnel [Ano35a, Ril37]. Persons [WW38, Fos39b]. Pertaining [Stu37b]. Phase [vH33b]. Physical [Fab38, Hud39, Kuv36, Leo32b, Leo35]. Physicists [Dem36b]. Pitfalls [Kis34b, Tru38, Wis34]. Place [Sch39d, TBSM38]. Planning [Cop34, Ebe38, Lub34, Sch39d]. Plans [KSF36]. Point [Ano30c]. Poisson [Ric37, Wal36]. Poland [Lan31, Ney35]. Policy [Hin38, Mea35a, Slo37]. Polish [Ano38m, Ano38s]. Political [HCW^{+}34]. Politics [Ric38a]. Population [Ano30c, Bur33, Cha31, Dor39, Ebe38, Edg31, Eel30, Kin31d, Kis34b, Lot38, Mal36, Met36, Met38, Sti34, T.34, Tru33, Tru35, Wat31, Whe36b, Wil30b, Wol32]. Populations [Ney38a, Woo32b]. Portrayal [Sib30]. Postal [Sis36]. Postponable [Tip39b]. Power [Fer36, Fer37, Lew37]. Practical [Hal36]. Pre [Edg31]. Pre-Census [Edg31]. Precious [Wil32a]. Predicting [CJS33]. Preliminary [Gar38]. Preparation [DM37]. Present [Dem36b, Gin30]. Presentation [Ano36o, Mav32, Mor37, Vol39]. Presented [SP37]. Presidential [Ric33b]. Presidents [Ano30n]. Prevention [Kos39]. Price [Ano32b, Ano36g, AC34, Asc37, Cop36, CD37, Den36, Fer36, Fis30, Fis36a, Fis36b, Gar38, Hum39, Kin32d, Kin32e, Kin39d, Kuv30, McI38b, Mea35a, RvS39a, RvS39b, Wai38]. Prices [Art39, Kin30g, Kin33g, Kin34f, Kin35a, Kin39c, Mil37a, Mil37b, Rey39, Sma33, Tur36, Wil32a, Woo32a]. Primary [Dun36, Hin38, vH33b]. Principal [Car38, Gir36]. Prison [San37]. Probability [Ano37h, Ano37g, Ano38f, Bar36, ML31, Mue36, Mue38, NMDW36, Rom35]. Probable [HG30]. Problem [Ayl37, CB36, Jar37, McI38b, Wel38a, Wil34h]. Problems [Bea30b, Bec38a, Bow36b, CSS39, Cla34a, Cla37, Fab38, Har39c, Hud39, Jon38, Kel39, Kin31e, Nep34, Pal39b, Per34, Sar34, SP37, Sel34, Sto34c]. Procedure [Ano34x]. Procedures [CJS33, San36a]. Proceedings [Ano36j, Ano36y, Ano37s, Ano38t, Ano39n]. Processes [Han34a]. Producer [Leo32b, Leo35, Rey39]. Product [Bak30, War34a]. Product-Moment [Bak30]. Production [Hud39, Joh38b, Joy30b, Joy34a, Kin39c, Kol35, Kol37, Kuv36, Leo32a, Leo32b, Leo35]. Productivity [B.39, Bea34, Bra39, Mag39, Mil37a, Wei38, Wei39a, Wub39]. Products [Sib30]. Professor [Hol37]. Program [Ach34b, Ano32b, Bec38a, DuB36, Fox36a, Giv35, Ror30, Ros34, Sch39d]. Programs [DC37b]. Progress [Ano33v, D.36, Gin30, H.30a, H.30b, H.30c, H.31a, H.31b, H.31c, H.32a, H.32b, H.32c, H.33a, H.33b, H.33c, H.34d, H.34b, H.34c, H.34e, H.35a, H.35d, H.35b, H.35c, H.36a, H.36b, H.36c, H.37a, H.37b, H.37c, H.37d, H.38a, H.38b, H.38c, Joy34b, Rid35, San36a]. Property [Car34, GH37, Gre34]. Proposal [DuB36, For32]. Prospective [Kin31d]. Prosperity [Bad33, Kin30f]. Protection [Ano37h, Ano37g, Mue36]. Public [Ano33f, Bur32b, Fos38, Har36a, Hin38, Kin31e, Sch39d, Tho34b, Wil31]. Publications [Ano30q, Ano38u, Ano38v, Ano39o, Ano39p, Ano39q, Ano39r]. Pulmonary [CC35]. Punched [Vic38, Vic39]. Punching [Dun32, Rea34]. Purchases [SP37]. Purchasing [Fer36, Fer37, Lew37, Tip39b]. Purchasing-Power [Fer36]. Purpose [Eps35, Mag39, PS30, San36a, Wil39a]. Purposes [Sch33a].

Quantitative [Ric38a]. Quantity [Sma33]. Quarterly [Ell38, Fis30]. Question [Kuz35a]. Questionnaires [Mit39b]. Quotients [Sib30].

R. [HW35]. Radio [Wel38a]. Railroad [Cam31, Dav39]. Railroads [Kin30f]. Railway [Dun35]. Random [Bur33, Stu34, Wor34]. random-difference [Wor34]. Randomly [CS39]. Ranked [Wal39d]. Ranks [Fri37a, Fri39b]. Rapidity [Whe34b]. Rate [CJ32, HH31, Jaf39, Mit39b, Sas30, Whe34b, Woo31, Woo32b]. Rates [Cha34a, Dow31, Dun35, Kop32b, Lin37a, Mal36, Mue34, Pat39, Sma33, Tol34, Vee30, Wur36]. Ratio [Bra34, Wal39d, Wis34]. Ratios [Rob34, Rob36]. Readjustment [Kin33e]. Real [Car34, GH37, Gre34, Tuc36b]. Received [Ano38u, Ano38v, Ano39o, Ano39p, Ano39q, Ano39r]. Receiving [WW38]. Recently [Ano31d, Rie31c]. Recession [Art39]. Reciprocal [DF37, Fet37]. Recommendations [Cha37]. Recording [San37]. Records [Edg31]. Recovery [Cop34, Kin32b, Kin33g, Lub34]. Reduce [Nyb37]. Reduction [Hum39]. Reference [Bak30, Bra36a, YD32]. Referendum [Gal38]. Reform [Kuz35a]. Regarding [Fri30b]. Region [Tay30b]. Regional [ZDM^{+}34]. Registration [Whe34a]. Regression [Hor32, Hor35b, Ric31, Wau35, Wau36]. Regressions [Mos37, Mos38]. Regulated [Kin34d]. Rejoinder [Fis36b, Kin39d, Lot30, Wai32a, War38b, Wei39a]. Relating [Gou36]. Relation [Bar38, BMH35, Cop36, CC35, Fis36a, Fis36b, Lot36, vH30c]. Relationships [Tru38]. Relative [Asc37, Val30]. Release [San36a]. Reliability [Gar38, MG34, ZDM^{+}34]. Relief [Ano36f, Bec38a, CJS33, Man34, Ros34, Sto34b, WW38]. Relieving [Kin33a]. Religious [Jaf39]. Remarks [Eze32b, Joh36, Wai32a]. Rental [Ach34b, SP34]. Rents [Cla34a, Per34]. Repealing [Wil31]. Replacement [Geo39]. Reply [B.39, Eze32b, Fer37, Fri34, Kuc30, Lot30, RvS39b]. Report [Fis30]. Reporting [DuB36, Jar37]. Reports [Ano33f, Ano34y, Fox36a, Tru33]. Representative [Ste39b]. Represented [Bur33]. Reproduction [Jaf39]. Reproductivity [Lot36]. Requirements [Alt37, Big37, DC37b, Mea35a, Sti34]. Research [Ano30a, Ano31b, Ano32e, Ano33b, Ano34x, Ano36g, Ano37y, Kin32f, Ros34, San34, Sch39d, Wis36]. Reserve [Gol36]. Resident [Dor32]. Residential [FW37]. Residual [Tru38]. Residuals [BT32]. Respect [MK37]. Responsibilities [Ric34]. Responsiveness [Den36]. Results [Ano37h, Ano37g, Lot38, Mue36, Pea32, Pee37, Wat30b]. Resume [Kin34e]. Retail [Kin33g]. Return [Mit39b]. Returns [SN30]. Revealed [Kin33h]. Review [Abb38, Abe34, Abr31, Abr35, Ach30, Ach34a, Ada33, Ada39, Ale37a, Ale37b, AE38, Alt35, Alt36, Amo37, Amo38, And35, And39, And30, Ang30, Ang32, Ang33, Ang39, Arn31, Art36, Bac31, Bal36, Bar37, Bar33, Bau36, Bau33, Bea31a, Bea30a, Bea31b, Bea35, Bec37, Bec38b, Bel31, Ben37a, Ber36a, Bez39, Bid34, Bla34, Bla37, Bla38, Bla36, Bla33, Bli32, Blo31, Blo37, Bob38, Bod39, Bog35, Bog36, Bog37, Bog38a, Boi36, Bor38, Bow35b, Boy33, Bra30, Bre39, Bri35, Bri31, Bro31, Bro38, Bro39, Bro36, Bru36, Bru37, Bru35, Buc38a, Buc38b, Bue38a, Bue38b, Bue38c, Bur32a, Bur30, Bur36a, Bur31, Bur36b, Bur39, Cam38a, Can39, Car36, Car30, Car35, Car31, Cav37, Cha30a, Cha33a, Cha39, Cha34b, Cha35a, Cha38, Cio37, Cio38, Cio39, Cla31].

Review [Cla32, Cla33, Cla39, Cli39, Coa31a, Col37, Col38, Com33, Cop30, Cop32b, Cor34, Cov37, Cov38, Cox36, Cox38, Cox39a, Cox39b, Cox39c, Cra35, Cra36, Cra37, Cre38, Cro38a, Cro32a, Cro30, Cro33, Cro36, Cro38b, Cro38c, Cru31, Cru35a, Cru36, Cun33, Cus36, Cus39, D.37, Dam37, Dau35, Dau36, Dau39, Dav30a, Dav32, Dav34a, Dav33b, Dav34b, Dav37a, Dea30, Dea32, Dem36a, Dem37, Dem39, Den37, Der39, Des38, Dew35, Dew36, Dew38, Die36, Dod36, Dod38, Dol39, Dou38a, Dou30, Dou38b, Dow38, Dre39b, Dre39a, Dru36, Dru38, DuB35, Dub31, Dul31, Dul34, Dul36a, Dul36b, Dul37, Dun39, Dun37, Ede37, Edi31, Edw39, Eld35, Eld36a, Eld36b, Ell37, Els36, Emm31, Eng38, Eps33, Eps34b, Eps34a, Eps34c, Eps36, Eva33, Eze32a, Eze37, Fab36, Fal38, Fal37]. Review [Fel30, Fel37, Fel36, Fel39, Fer31a, Fer32a, Fer32b, Fer34, Fer39, Fic35, Fis32, Fis37a, Fis37b, Fis36c, Fle35, Fle38a, Fle38b, Fol35, For39, Fos35b, Fos35a, Fos39a, Fra30, Fre38, Fri30a, Fri39a, Fri39c, Fri36a, Fri36b, Fri37b, Fri38, Fri39d, Fri39e, Fry36, Gay31a, Gay31b, Gay31c, Geh31, Geh37, Geo37, Ger39, Gev35, Gid33, Gid39, Gil33a, Gil34, Gil35, Gil36a, Gil36b, Gil37, Giv31a, Gla30, Glo31, Gok36, Goo34, Gra38, Gre36a, Gre37, Gre36b, Gre38, Gre39, Gri36, Gri35, Had38, Hag37, Hag39, Hai32, Hai33, Hal33, Hal37, Hal38, Ham33a, Ham35, Ham36, Han30, Han33, Han35, Han37, Har37a, Har37b, Har30, Har32a, Har34a, Har39a, Har32b, Har36b, Har39b, Har39d, Har37c, Hau37, Hay36a, Hed36, Hei39, Hen32, Hen39, Her30, Hew30, Hil38]. Review [Hin39, Hin32, His32, Hob39, Hof34a, Hof34b, Hog31a, Hog31b, Hog34a, Hol32, Hol36, Hom34, Hoo33, Hoo37, Hoo38a, Hoo38b, Hoo39, Hor35a, Hot30a, Hot32a, Hot32b, Hot33b, Hot33a, Hot35a, Hot35b, Hot35c, Hot35d, Hot37a, Hot37b, Hot38, Hot39, Hou39, How36a, How36b, How37, Hoy39, Hub37, Hur34a, Ing39, Inm39, Isa34, Isa36, Isa38, Ise36, Jam35, JP36, Jen34, Jer31, Jer32, Jer36, Joh38a, Joh39, Jon37, Joy30a, Joy33, Kah39, Kel38, Kel33, Kel34, Kel36, Kes36, Kil38, Kin39a, Kin31a, Kin32a, Kin33c, Kin39b, Kis34a, Kis36, Kni37, Kni38, Kni39, Koe36, Kop30, Kop31, Kop32a, Kre31, Kre32, Kro34, Kuc30, Kuc31a, Kul34, Kul38, Kuz30a, Kuz30b, Kuz31, Kuz32a, Kuz33, Kuz35b, Kuz38a, Kuz38b, Kyr30, Kyr33, Lan33, Lan35, Lat30, Lat34, Lea38]. Review [Lea39, Led36, Leo36b, Leo37, Ler38, Lev30, Lev37, Lew38b, Lew38a, Lew39, Lin32, Lin39, Liv39, Liv38, Loc36, Lor39, Lot30, Lou39, Lov36, Lub30, Lun34, Lun35, Lun36a, Lun37, Lun38, Lun39, Lun36b, Lut33, Lut37, Lut39, Lyo31, Lyo36, Mac33, Mac38, Mac39a, Mac39b, Mad38, Mal35, Mar30, Mar33, Mar34a, Mar35a, Mar35b, Mar39a, Mar39b, Mar39c, Mas33, Mas35, Mas36, Mas37, Mas38, McC36a, McC36b, McD38, McD39a, McD39b, McI36, McI38a, McK36, McL36, McL38, McL39, Mea31, Mea33, Mea37, Men39a, Men30, Men32, Men35, Mer33, Mey38, Mil38, Mil30, Min30, Min33, Min35, Mit30, Mit32, Mit36, Mon38, Moo37, Mor39, Mor36a, Mor36b, Mud37, Mud39, Mus33, Mye39, Nad39, Nag39, Nat35, Nat37, Nat39, Nel39, New38, Ney39, Nic39]. Review [Nys34, Orc30, Orc31a, Orc31b, Orc33a, Orc33b, Orc34, Orc35, Orc36, Orc37, Ost31, Pal38a, Pal39a, Pal36, Pal38b, Pal39c, Pal39d, Par37a, Par39, Par38, Par37b, Pat30, Pat34a, Pat34b, Pat37, Pea38, Pea36a, Pea36d, Pec31, Pec33, Pen31, Pen32a, Per37, Per31, Pet33, Pet39, Phe31, Phe37a, Pri35, Pro39a, Pro39b, Puf39, Ree33, Ree35, Ree37a, Ree39, Ric38b, Rid37, Rie32, Rie30, Rie31a, Rie31b, Rie37, Rig33a, Rig39, Ril38, Rob33, Rob38, Rob39, Rod32, Roe30, Roe38, Rog34, Rol30a, Rol30b, Rol31, Roo36a, Roo36b, Rub30, Rub33, San36b, San39, San31, Sch39a, Sch39b, Sch38a, Sch38b, Sch39c, Sch31, Sch36a, Sch39e, Sch36b, Sch36c, Sch39f, Sco37, Sco39, Sco38, Sea36, Sel38, Sel32, Sel33, She39, Shi37a, Shi37b, Shi38, Shl39, Sho33]. Review [Sho35, Sho36, Shr39, Sie39, Sim37, Sim38, Slo30, Slo31, Sma34, Smi32, Smi33, Smi35a, Smi35b, Smi36, Smi37b, Smi37a, Smi39, Sne39, Sou30, Sou34, Sou36, Spa35, Spe30, Spe31, Spe39, Spi39, Ste34a, Ste34b, Ste35d, Ste36b, Ste36c, Ste37b, Ste39c, Sto32, Sto31, Sto37, Sto38, Sto39b, Str37, Str39, Stu37a, Swa36, Swa39, Swe35, Swi38, Tae38, Tay30a, Tay31, Tay35, Tay38, Tea38a, Tea38b, Tea38c, Tee37a, Tee37b, Tho32a, Tho32b, Tho35a, Tho32c, Tho35b, Tho32d, Tho39a, Tho30, Tho32e, Tho34a, Tho38b, Tho39b, Tho39c, Thr39, Tim30, Tim31, Tim32a, Tim32b, Tim33, Tim35, Tim36, Tim37, Tip36, Tip37, Tip39a, Too35, Too36, Too37, Tre38, Tro39, Tru39, Tuc36a, Tuc37, Tuf37, Tul33, Ush33, Vai36, Van35, Van39, Vee36]. Review [Vee39, Via37, Vin30, Wad34, Wai39, Wai36, Wal39a, Wal39b, Wal38, Wal37a, Wal37b, War37, War38a, War38b, War34b, War30, War36, Wat30a, Wat38a, Wat36, Wat38b, Wat39a, Wat39b, Web38, Wei32a, Wei32b, Wei39b, Wel36a, Wel39a, Wel35, Wel36b, Wel39b, Wel38b, Wel39c, Whe36a, Whi35, Whi37, Whi39, Wil36a, Wil36b, Wil37a, Wil38b, Wil38c, Wil33a, Wil32b, Wil36c, Wil36d, Wil37c, Wil38d, Wil39b, Wil32c, Wil33b, Wil33c, Wil34g, Wil36e, Wil30c, Wil35b, Wil39c, Win36, Win39, Wit36, WS39, Wol35a, Wol38, Wol35c, Woo37, Wor37, Wor39, Wri31, Wri34b, Wri37, Wri38, Wri39, Ynt30a, Ynt30b, Ynt32, Ynt35a, Ynt35b, Ynt37, Yod33, Yod34, You32, You36b, You36a, Zap31, Zel36, Zim39, dC34, dC37a, dR37, vB39, vH30a, vH30b, vH32, vH34a, vH34b, vH38a]. Review [vS31]. Revised [CD37]. Revising [Hog34b]. Revision [Ano38w, SP31]. Revival [Han34a]. Richards [vH33a]. Roope [vH33a]. Rorty [Bel36]. Rule [Ano34z, Ano35f, Ano35v]. Rural [Man34].

S [Ney38b]. Sales [Ayr38, Cow37, Hum39]. Sample [Bra33b, Han34b, Rid31]. Samples [Ano31d, Bat35, Cru33, Rie31c, Woo34]. Sampling [Bow36b, Bur33, COP32a, Coa31b, Coc39, Gal38, Kis34b, Man34, Ney38a, SP37, Ste34c, Ste39b, Vic38, Wil38f]. Savings [Sis36]. Scale [Ste39b]. Scales [Tol34, TK33, vH31]. Scatter [Li34]. Scatter-Diagram [Li34]. Schedules [Sch33a]. Science [Ano30a, Ano31b, Ano32e, Ano33b, Coa39]. Sciences [Cad35]. Scientific [Gir39, Ror31]. Scores [Hor31, RT33]. Seasonal [Kuz32b, Mah30, Pis32, Spu37, Wel30, WF34, vH33c, vS30]. Seasonality [Yod38]. Second [NT36]. Secondary [Lom38, vH33b]. Secrist [Hot33a]. Secular [Wel30]. Securities [Lom38]. Security [Alt37, Cla37, DC37b, Giv35, Gou36, Jar37, Kin34d]. Selection [Dun36, DM37]. September [Ano30h, Ano31h, Ano32j, Ano33k, Ano34o]. Series [Ano38n, Ber36b, Ell38, Hom33, Joh35, Joh36, Lea31, Sas30, Sib30, Spi34, Stu34, ZDM^{+}34, Zub39, Wor34]. Service [Ano32f, Fis33, Giv34c]. Services [Giv34a, Giv34b, Giv34c, KSF36, Ric33a, Zab30]. Session [Wil30a]. Setting [Bow36a, Pat39]. Seventh [Ano36j, Ano36y, Fis30]. Several [Ano38f, Lot36, Mue38, Wil35a]. Shading [vH38b]. Shall [Cha33b, Fri30b]. Shape [Hor31]. Share [Rob34, Rob36]. Short [Kin30a, Kuz32b]. Short-Term [Kin30a]. Short-Time [Kuz32b]. Should [Cop34, Kin35b, Lub34]. Sickness [CP34]. Side [Hal36]. Significance [Bak30, Bra33b, Car38, Cha34a, Fry38, Gol36, Pea31, Pea32, Sto33, Woo34, Zub39]. Simple [CS39, Gri30]. Simplest [Sas30]. Simplified [Bea30b, Rea34, Wau35]. Simultaneous [Mos37, Mos38]. Since [Ano30e, Ano31g, Ano31f, Ano31h, Ano32g, Ano32h, Ano32j, Ano33h, Ano33j, Ano33i, Ano33k, Ano34l, Ano34o, Bea34, Kin32c, Ano30g, Ano30f, Ano30h, Ano31e, Ano32i, Ano34n, Ano34m, Wil30b]. Single [Ano38f, Mue38]. Size [GB34, WH35]. Sizes [Bur33, Gir39, Lor35]. Skewness [Bat35]. Slide [Ano34z, Ano35f, Ano35v]. Slumps [Kin32e]. Small [Ano31d, Cha34a, Rid31, Rie31c]. Snyder [Man45]. Social [Alt37, Ano30a, Ano31b, Ano32e, Ano32f, Ano33b, Cad35, Cla37, Dav37b, DC37b, Edw33, EM34b, Giv31b, Giv35, HW35, Hur34b, Jar37, Kis32, Ste34c, SN30, Wil38e, Kuc31b]. Social-Economic [Edw33, EM34b]. Societies [Wil34e]. Society [Ano38m, Ano38s, Coa39, Dod37]. Sociological [Bow36b, Sto33, Sto34a]. Sociology [Sto34c]. Some [Ber38, Cla34a, Dem38, Eze32b, Har39c, Hay36b, Hum37, Jon38, Kel39, Kem34, Lot38, Nep34, Pea32, Per34, Ric30a, Ric30b, Wai32b, Wei38, Wil34a]. Sources [Hog30, Stu37b]. Soviet [Rom35]. Space [Ste34c]. Spain [Edg31]. Spatially [Nep34]. Special [Bak30, Bra36a, Haa36, Ste38b, YD32]. Specific [Val30]. Spent [Bad31]. Spurious [Ric31]. Square [Ber38, Cam38b]. Squares [Cro32b, Mos39]. Stability [Den36, For32]. Stabilized [HP31, Kin39c]. Stabilizing [Fox36b, Mea35a]. Stage [H.31c]. Standard [Fra31, Gir39, Imm37, Mav32, Mos37, Mos38, Mos39, Sch30, Sch33b, TK33, Wil38f, vH33c]. Standardization [CSS39]. Standards [BMH35, Cow37]. State [Bow36a, Hur34b, Mye34, Phe34b, Sch39d, YD32, Mal36, Sch35]. Statements [Slo37, Ste31]. States [CD37, Kis32, SN30, Ben37b, BA37a, BA37b, Cam31, Cha31, Edw33, EM34b, Gol34, Kin31d, Kin33d, Kin34e, Leo36a, Lot36, Mea35b, Rig33b, Sny34, Sog33, Sti34, Whe34a, Wil34a]. Statistical [Alt37, Ano30b, Ano30i, Ano30q, Ano33d, Ano34d, Ano34z, Ano35b, Ano35f, Ano35v, Ano35s, Ano35t, Ano35u, Ano36w, Ano36z, Ano36-27, Ano36-28, Ano36-29, Ano37t, Ano37u, Ano37v, Ano37w, Ano38l, Ano38m, Ano38s, Ano38x, Ano38y, Ano38z, Ano38-27, Ano39s, Ano39t, Ano39u, Ayl37, Bar36, Big37, Bra36a, Bra36b, Bur37b, Bur38, Bur33, Cha37, Cla37, CD36, CC35, Dea37, Dic35, Dod37, DC37b, DW35, FM31, FM32, Giv31b, Giv35, Han39, Hin38, Kem34, Kin31c, Kin31f, Kin32f, Kin34b, Kos39, Kuz35a, MF35, MH39, Ney35, Pat39, Pea36c, RG30, Ric34, Ric30a, Ric30b, Rid35, Ros34, San37, Sar34, Sel34, She31, Sne36, Ste35a, Ste38b, Sti34, Sti39, Tay37, Tru38, Vol39, Wil35a, Wil30a, Wil34c, Wil34d, Wil34e, Wil37b, Zab30, ZDM^{+}34, vH31]. Statisticians [Ano37i, Hot30b]. Statistics [Ano32b, Ano33e, Ano33f, Ano34e, Ano35k, Ano36f, Ano36q, Ano37a, Ano37y, Ano38a, Ano38k, Aus35, Bow35a, Bur32b, Cam33, Coh38, CF39, Cop34, Dav37b, Dem36b, Dor32, DuB36, Dun36, Fis33, Gil33b, Gin30, Giv34a, Giv34b, Giv34c, Gou36, Haa36, Hal36, HW35, Har39c, Hay36b, Hil31, HCW^{+}34, Hot30b, Hur34b, Jan38, Joy34b, Kel39, Ken36, Kin33h, Lan31, Leo36a, Llo38, Lub34, Mar34b, MF35, Mea35b, Mil35, Mye34, Nix38, Ogb32, Pet37, Phe34b, Ric33a, Rom35, Ror31, RT33, SS33, Spr34, Tay34, Tho34b, Tho36b, Wal39c, Wil38a, Wil39a, Wil34b, Wil30d, Wis36, CD37, Ebl38b, Per36, Sho37]. Stevens [Ano37s]. Stimulus [Hum39]. Stock [Kin31g, Kin34f, Kuv30, McI38b, Tur39]. Straw [Cru33]. Straw-Vote [Cru33]. Strike [Pet37]. Strikes [Yod38]. Structure [Fox36b]. Student’s [NT36]. Studies [Car34, Haa36, Han34b]. Study [CJ32, CC35, Kis34b, KSF36, Leo32a, Mag33, SP37, SN30, Wil38e]. Subclass [Sne34]. Subgroups [Fab35]. Subramanian [Fri34]. Subsets [Mos37, Mos38, Wau36]. Successfully [Kin32c]. Sufficient [Fri30b]. Sugar [Bra36a]. Suggestions [Ano35v]. Sum [Spi34]. Summation [Li34, Mue34, MH39, Ste32]. Supervision [Fos38, Jon38]. Supreme [Wal36]. Surface [Ano32d, Hun32]. Survey [San36a]. Surveys [Hog30, JW39, Mit39c, MG34, Ste39b, Web39]. Swedish [Tho36a, Tho37a]. Sydenstricker [Kin36b]. Symmetrical [Sas30]. System [Met36, Met38, Sis36]. Systems [Big37].

Table [Bra33a]. Tables [DM37, Sne34]. Tabulation [Dun32, Tru31, Tru35]. Tabulations [Lin37b]. Tabulator [DM37]. Taming [WF34]. Tangent [vH33a]. Tax [Sch33a]. Teaching [CF39, Har39c, Kel39, WD32]. Technical [Hud39, Kin34f]. Technique [Vic38, Vic39]. Techniques [Sne36]. Technological [Cha35b, Wub39]. Tendencies [Nug38]. Term [Kin30a, Sny34]. Test [Ano37h, Ano37g, Ber36b, Ber38, Bra33b, Cam38b, Dem34, Dem38, Fry38, Mue36, Pea31, Pea32, Pee37, Wil35a]. Testing [Dod37, NT36, Wat30b]. Tests [CS39, Fri30b, Sto33]. Textile [Cha37]. Their [Car38, Kin35a, Kin39c, Kuz32b, Mos37, Mos38, Wei38, Wei32c]. Them [Kin32e]. Theory [Ano31d, FM31, FM32, Kin31c, Kin35a, Ney35, Ney38a, PHT^{+}30, Rid31, Rie31c, Rom35, Sch36b, Sma33]. there [Kin30c]. Third [SP31]. Thoughts [Dem38]. Three [Wis31]. Time [Ano38n, Fra31, Kuz32b, Lea31, Pis34, Sas30, Sch39d, SS33, Ste34c, Tay32, Wor34]. Time-Series [Ano38n]. Today [Hot30b]. Tool [Fra31]. Total [Bec38a]. Tract [GB34]. Trade [Bra36a, Com34, DF37, Fet37, Hal36, Joh38b, Llo38, Mea35b, Sar34, SP31]. Training [Ano34e, Per36]. Transient [Leo35]. Travel [Ano30c]. Treatment [Vol39]. Trend [Dou32, Joh35, Joh36, Wel30]. Trends [BT32, Car38, Smi35c, Tho38a, Tur39]. Trigonometrical [vH31]. Triumph [Hot33a]. Tuberculosis [CC35, Dow31]. Turnover [Vee30]. Twenties [B.39, Bra39, Wei39a]. Two [AC34, Bat35, Bur37a, CS39, Pee37, Spi34, Tay32, Zub39, vH30c, vH33b]. Two-Way [AC34]. Type [MK37]. Types [FM31, FM32, Kis32, Tay32].

Unemployed [Kin33a]. Unemployment [Big37, Bow36a, Giv31b, Hil31, Hog30, Hum37, JW39, Kin32g, Kre34, Ror36, Tol34, Web39]. Unequal [Yat34]. Union [Rom35]. Unique [Bra33b]. Unit [HCW^{+}34]. United [Ben37b, BA37a, BA37b, Cam31, Cha31, CD37, Edw33, EM34b, Gol34, Kin31d, Kin33d, Kin34e, Leo36a, Lot36, Mea35b, Rig33b, Sny34, Sog33, Sti34, SN30, Whe34a, Wil34a]. untitled [Abb38, Abe34, Abr31, Abr35, Ach30, Ach34a, Ada33, Ada39, Ale37a, Ale37b, AE38, Alt35, Alt36, Amo37, Amo38, And35, And39, And30, Ang30, Ang32, Ang33, Ang39, Arn31, Art36, Bac31, Bal36, Bar37, Bar33, Bau36, Bau33, Bea31a, Bea30a, Bea31b, Bea35, Bec37, Bec38b, Bel31, Ben37a, Ber36a, Bez39, Bid34, Bla34, Bla37, Bla38, Bla36, Bla33, Bli32, Blo31, Blo37, Bob38, Bod39, Bog35, Bog36, Bog37, Bog38a, Boi36, Bor38, Bow35b, Boy33, Bra30, Bre39, Bri35, Bri31, Bro31, Bro38, Bro39, Bro36, Bru36, Bru37, Bru35, Buc38a, Buc38b, Bue38a, Bue38b, Bue38c, Bur32a, Bur30, Bur36a, Bur31, Bur36b, Bur39, Cam38a, Can39, Car36, Car30, Car35, Car31, Cav37, Cha30a, Cha33a, Cha39, Cha34b, Cha35a, Cha38, Cio37, Cio38, Cio39, Cla31]. untitled [Cla32, Cla33, Cla39, Cli39, Coa31a, Col37, Col38, Com33, Cop30, Cop32b, Cor34, Cov37, Cov38, Cox36, Cox38, Cox39a, Cox39b, Cox39c, Cra35, Cra36, Cra37, Cre38, Cro38a, Cro32a, Cro30, Cro33, Cro36, Cro38b, Cro38c, Cru31, Cru35a, Cru36, Cun33, Cus36, Cus39, D.37, Dam37, Dau35, Dau36, Dau39, Dav30a, Dav32, Dav34a, Dav33b, Dav34b, Dav37a, Dea30, Dea32, Dem36a, Dem37, Dem39, Den37, Der39, Des38, Dew35, Dew36, Dew38, Die36, Dod36, Dod38, Dol39, Dou38a, Dou30, Dou38b, Dow38, Dre39b, Dre39a, Dru36, Dru38, DuB35, Dub31, Dul31, Dul34, Dul36a, Dul36b, Dul37, Dun39, Dun37, Ede37, Edi31, Edw39, Eld35, Eld36a, Eld36b, Ell37, Els36, Emm31, Eng38, Eps33, Eps34b, Eps34a, Eps34c, Eps36, Eva33, Eze32a, Eze37, Fab36, Fal38, Fal37]. untitled [Fel30, Fel37, Fel36, Fel39, Fer31a, Fer32a, Fer32b, Fer34, Fer39, Fic35, Fis32, Fis37a, Fis37b, Fis36c, Fle35, Fle38a, Fle38b, Fol35, For39, Fos35b, Fos35a, Fos39a, Fra30, Fre38, Fri30a, Fri39a, Fri39c, Fri36a, Fri36b, Fri37b, Fri38, Fri39d, Fri39e, Fry36, Gay31a, Gay31b, Gay31c, Geh31, Geh37, Geo37, Ger39, Gev35, Gid33, Gid39, Gil33a, Gil34, Gil35, Gil36a, Gil36b, Gil37, Giv31a, Gla30, Glo31, Gok36, Goo34, Gra38, Gre36a, Gre37, Gre36b, Gre38, Gre39, Gri36, Gri35, Had38, Hag37, Hag39, Hai32, Hai33, Hal33, Hal37, Hal38, Ham33a, Ham35, Ham36, Han30, Han33, Han35, Han37, Har37a, Har37b, Har30, Har32a, Har34a, Har39a, Har32b, Har36b, Har39b, Har39d, Har37c, Hau37, Hay36a, Hed36, Hei39, Hen32, Hen39, Her30, Hew30, Hil38]. untitled [Hin39, Hin32, His32, Hob39, Hof34a, Hof34b, Hog31a, Hog31b, Hog34a, Hol32, Hol36, Hom34, Hoo33, Hoo37, Hoo38a, Hoo38b, Hoo39, Hor35a, Hot30a, Hot32a, Hot32b, Hot33b, Hot35a, Hot35b, Hot35c, Hot35d, Hot37a, Hot37b, Hot38, Hot39, Hou39, How36a, How36b, How37, Hoy39, Hub37, Hur34a, Ing39, Inm39, Isa34, Isa36, Isa38, Ise36, Jam35, JP36, Jen34, Jer31, Jer32, Jer36, Joh38a, Joh39, Jon37, Joy30a, Joy33, Kah39, Kel38, Kel33, Kel34, Kel36, Kes36, Kil38, Kin39a, Kin31a, Kin32a, Kin33c, Kin39b, Kis34a, Kis36, Kni37, Kni38, Kni39, Koe36, Kop30, Kop31, Kop32a, Kre31, Kre32, Kro34, Kuc31a, Kul34, Kul38, Kuz30a, Kuz30b, Kuz31, Kuz32a, Kuz33, Kuz35b, Kuz38a, Kyr30, Kyr33, Lan33, Lan35, Lat30, Lat34, Lea38, Lea39, Led36, Leo36b]. untitled [Leo37, Ler38, Lev30, Lev37, Lew38b, Lew38a, Lew39, Lin32, Lin39, Liv39, Liv38, Loc36, Lor39, Lou39, Lov36, Lub30, Lun34, Lun35, Lun36a, Lun37, Lun38, Lun39, Lun36b, Lut33, Lut37, Lut39, Lyo31, Lyo36, Mac33, Mac38, Mac39a, Mac39b, Mad38, Mal35, Mar30, Mar33, Mar34a, Mar35a, Mar35b, Mar39a, Mar39b, Mar39c, Mas33, Mas35, Mas36, Mas37, Mas38, McC36a, McC36b, McD38, McD39a, McD39b, McI36, McI38a, McK36, McL36, McL38, McL39, Mea31, Mea33, Mea37, Men39a, Men30, Men32, Men35, Mer33, Mey38, Mil38, Mil30, Min30, Min33, Min35, Mit30, Mit32, Mit36, Mon38, Moo37, Mor39, Mor36a, Mor36b, Mud37, Mud39, Mus33, Mye39, Nad39, Nag39, Nat35, Nat37, Nat39, Nel39, New38, Ney39, Nic39, Nys34, Orc30, Orc31a, Orc31b]. untitled [Orc33a, Orc33b, Orc34, Orc35, Orc36, Orc37, Ost31, Pal38a, Pal39a, Pal36, Pal38b, Pal39c, Pal39d, Par37a, Par39, Par38, Par37b, Pat30, Pat34a, Pat34b, Pat37, Pea38, Pea36a, Pea36d, Pec31, Pec33, Pen31, Pen32a, Per37, Per31, Pet33, Pet39, Phe31, Phe37a, Pri35, Pro39a, Pro39b, Puf39, Ree33, Ree35, Ree37a, Ree39, Ric38b, Rid37, Rie32, Rie30, Rie31a, Rie31b, Rie37, Rig33a, Rig39, Ril38, Rob33, Rob38, Rob39, Rod32, Roe30, Roe38, Rog34, Rol30a, Rol30b, Rol31, Roo36a, Roo36b, Rub30, Rub33, San36b, San39, San31, Sch39a, Sch39b, Sch38a, Sch38b, Sch39c, Sch31, Sch36a, Sch39e, Sch36c, Sch39f, Sco37, Sco39, Sco38, Sea36, Sel38, Sel32, Sel33, She39, Shi37a, Shi37b, Shi38, Shl39, Sho33, Sho35, Sho36, Shr39, Sie39, Sim37]. untitled [Sim38, Slo30, Slo31, Sma34, Smi32, Smi33, Smi35a, Smi35b, Smi36, Smi37b, Smi37a, Smi39, Sne39, Sou30, Sou34, Sou36, Spa35, Spe30, Spe31, Spe39, Spi39, Ste34a, Ste34b, Ste35d, Ste36b, Ste36c, Ste37b, Ste39c, Sto32, Sto31, Sto37, Sto38, Sto39b, Str37, Str39, Stu37a, Swa36, Swa39, Swe35, Swi38, Tae38, Tay30a, Tay31, Tay35, Tay38, Tea38a, Tea38b, Tea38c, Tee37a, Tee37b, Tho32a, Tho32b, Tho35a, Tho32c, Tho35b, Tho32d, Tho39a, Tho30, Tho32e, Tho34a, Tho38b, Tho39b, Tho39c, Thr39, Tim30, Tim31, Tim32a, Tim32b, Tim33, Tim35, Tim36, Tim37, Tip36, Tip37, Tip39a, Too35, Too36, Too37, Tre38, Tro39, Tru39, Tuc36a, Tuc37, Tuf37, Tul33, Ush33, Vai36, Van35, Van39, Vee36, Vee39, Via37, Vin30, Wad34, Wai39, Wai36]. untitled [Wal39a, Wal39b, Wal38, Wal37a, Wal37b, War37, War38a, War34b, War30, War36, Wat30a, Wat38a, Wat36, Wat38b, Wat39a, Wat39b, Web38, Wei32a, Wei32b, Wei39b, Wel36a, Wel39a, Wel35, Wel36b, Wel39b, Wel38b, Wel39c, Whe36a, Whi35, Whi37, Whi39, Wil36a, Wil36b, Wil37a, Wil38b, Wil38c, Wil33a, Wil32b, Wil36c, Wil36d, Wil37c, Wil38d, Wil39b, Wil32c, Wil33b, Wil33c, Wil34g, Wil36e, Wil30c, Wil35b, Wil39c, Win36, Win39, Wit36, WS39, Wol35a, Wol38, Wol35c, Woo37, Wor37, Wor39, Wri31, Wri34b, Wri37, Wri38, Wri39, Ynt30a, Ynt30b, Ynt32, Ynt35a, Ynt35b, Ynt37, Yod33, Yod34, You32, You36b, You36a, Zap31, Zel36, Zim39, dC34, dC37a, dR37, vB39, vH30a, vH30b, vH32, vH34a, vH34b, vH38a, vS31]. Upon [Dor32, Dun36, GB34, Kin35a, Tho33, SN30]. Urban [Kap38, SP37]. Use [Ano37y, BT32, Coc39, Fer31b, Fri37a, Fri39b, Wri34a, Fry30, Wat30b, Wor34]. Used [JW39, Kin32c, Pet37, Web39]. Uses [Hay36b, Wei38]. Using [Hun32, Mye34]. Utility [Har36a, Kin31e].

Validity [Wat30b]. Valuation [Har36a]. Value [HG30, Smi34, Spi34, War34a]. Values [Ano38f, GH37, Mue38, Yan33]. Variable [Cow32, HG30]. Variables [Ach34b, Ano36h, Cow32, Mos37, Mos38, Nep34, Ric31, Sto36, Wau36, Wil35a, Wis31, Won37]. Variance [Bra33a, Coc39, Fri37a, Fri39b, SS33]. Variate [vS30]. Variates [Men39b]. Variation [Mah30, Ogb35, Spu37, Wel30, YD32, vS30]. Variations [Asc37, Bea30b, Bra38, Kin30e, Kuz32b, Kyr37, Tho37b, Wil37d]. Various [Kin33a, Kis32, ZDM^{+}34]. Vehicles [Sch33a]. Velocity [Tru38]. Versus [Ree37b]. Vienna [Kuc31b]. View [CD36]. Viewpoint [She31]. Visiting [Kin34b]. Visualization [vH31]. Vital [Aus35, Bow35a]. Volume [Ano30r, Ano31p, Ano32r, Ano33w, Ano34-27, Ano35w, Ano36-30, Ano37x, Ano38-28, Ano39v, Ayl37, Fab38, FW37, Kuv36, Leo32b, Leo35, SP31]. Vote [Cru33, GS36]. vs [HCW^{+}34].

W [Man45, Ney38b]. Wage [Hog34b, Mit39c, Pat39]. Wages [HD30, Nix38, Tuc36b]. Wanted [Tho36b]. Warburton [Kuz38b, War38b]. Warren [Fos39b]. Washington [Wil34f]. Watkins [Ste34d]. Way [AC34, Gos39]. Wealth [Ano36e, Cop35, Wil34a, Ynt33]. Weekly [Fis30, Pis32, WF34]. Weighted [Art37]. Weighting [Bra34]. Welfare [Ano32f, Ano33f, Hur34b, Sch39d]. Westergaard [Sto33]. Westerner [Wil30b]. which [Fri30b, YD32]. White [SN30, HW35]. Whither [Kin30g]. Who [Cha33b, Cop34, Lub34]. Whole [Bur37b]. Wholesale [Cop36, CD37, Fis30, Fis36a, Fis36b, Hic39]. whose [Woo32b]. Wildlife [Ano37n]. will [Kin31g]. Wisconsin [Ebl38b, YD32]. without [Sib30]. Women [SN30]. Work [Ano33v, Bec38a, Bra36b, D.36, H.30a, H.30b, H.30c, H.31a, H.31b, H.31c, H.32a, H.32b, H.32c, H.33a, H.33b, H.33c, H.34d, H.34b, H.34c, H.34e, H.35a, H.35d, H.35b, H.35c, H.36a, H.36b, H.36c, H.37a, H.37b, H.37c, H.37d, H.38a, H.38b, H.38c, Nix38, Tho34b]. Worker [Mag33]. Workers [Edw33, EM34b, Smi35c, Wei32c]. Working [Fra31]. World [Cha37, Wil32a].

Yarding [Ayl37]. Years [Kre34, SN30]. Yield [Bea30b]. Yields [Tho33].

York [Ano36j, Ano36y, Hic39, Kuv30, Mal36, Sch35].

References

Abbott:1938:BRB

[Abb38] Charles Cortez Abbott. Book review: [untitled]. Journal of the American Statistical Association, 33(204):777–779, December 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279093.

Abel:1934:BRB

[Abe34] Theodore Abel. Book review: [untitled]. Journal of the American Statistical Association, 29(186):235–236, June 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278311.

Abrahamson:1931:BRB

[Abr31] Albert Abrahamson. Book review: [untitled]. Journal of the American Statistical Association, 26(175):370–371, September 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277929.

Abrams:1935:BRB

[Abr35] Ray H. Abrams. Book review: [untitled]. Journal of the American Statistical Association, 30(189):133–134, March 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278548.

Arthur:1934:TWD

[AC34] Henry B. Arthur and Lawrence W. Conant. A two-way distribution of price changes. Journal of the American Statistical Association, 29(186):175–179, June 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278287.

Achinstein:1930:BRB

[Ach30] Asher Achinstein. Book review: [untitled]. Journal of the American Statistical Association, 25(171):370–371, September 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278208.

Achinstein:1934:BRB

[Ach34a] Asher Achinstein. Book review: [untitled]. Journal of the American Statistical Association, 29(185):117–119, March 1934. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278483.

Achinstein:1934:OCV

[Ach34b] Asher Achinstein. Operating cost variables in a low rental housing program. Journal of the American Statistical Association, 29(185): 33–37, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277799.

Adams:1933:BRB

[Ada33] Arthur B. Adams. Book review: [untitled]. Journal of the American Statistical Association, 28(182):257–258, June 1933. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278057.

Adams:1939:BRB

[Ada39] Harold E. Adams. Book review: [untitled]. Journal of the American Statistical Association, 34(206):461, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278898.

Alexander:1938:BRB

[AE38] R. S. Alexander and R. Parker Eastwood. Book review: [untitled]. Journal of the American Statistical Association, 33(202):476–478, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278938.

Alexander:1937:BRBa

[Ale37a] R. S. Alexander. Book review: [untitled]. Journal of the American Statistical Association, 32(198):435–436, June 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279604.

Alexander:1937:BRBb

[Ale37b] R. S. Alexander. Book review: [untitled]. Journal of the American Statistical Association, 32(200):790–791, December 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279397.

Altschul:1935:BRB

[Alt35] Eugen Altschul. Book review: [untitled]. Journal of the American Statistical Association, 30(191):643–644, September 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278123.

Altschul:1936:BRB

[Alt36] Eugen Altschul. Book review: [untitled]. Journal of the American Statistical Association, 31(194):432–433, June 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278587.

Altmeyer:1937:SRS

[Alt37] Arthur J. Altmeyer. Statistical requirements of the social security board for efficient administration. Journal of the American Statistical Association, 32(197):15–18, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279854.

Amos:1937:BRB

[Amo37] J. Ellwood Amos. Book review: [untitled]. Journal of the American Statistical Association, 32(198):422–425, June 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279596.

Amos:1938:BRB

[Amo38] J. Ellwood Amos. Book review: [untitled]. Journal of the American Statistical Association, 33(201):302, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279154.

Andrews:1930:BRB

[And30] John B. Andrews. Book review: [untitled]. Journal of the American Statistical Association, 25(172):503–504, December 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277995.

Anderson:1935:BRB

[And35] Thomas J. Anderson, Jr. Book review: [untitled]. Journal of the American Statistical Association, 30(192):770–771, December 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277748.

Anderson:1939:BRB

[And39] Ira D. Anderson. Book review: [untitled]. Journal of the American Statistical Association, 34(207):600–601, September 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279502.

Angell:1930:BRB

[Ang30] James W. Angell. Book review: [untitled]. Journal of the American Statistical Association, 25(170):232–233, June 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277643.

Angell:1932:BRB

[Ang32] James W. Angell. Book review: [untitled]. Journal of the American Statistical Association, 27(180):445–447, December 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278650.

Angell:1933:BRB

[Ang33] James W. Angell. Book review: [untitled]. Journal of the American Statistical Association, 28(181):113–114, March 1933. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278019.

Angell:1939:BRB

[Ang39] James W. Angell. Book review: [untitled]. Journal of the American Statistical Association, 34(207):616–618, September 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279518.

Anonymous:1930:ASS

[Ano30a] Anonymous. Awards of the social science research council. Journal of the American Statistical Association, 25(171):334–336, September 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2278199.

Anonymous:1930:CLA

[Ano30b] Anonymous. Constitution and by-laws of the American statistical association. Journal of the American Statistical Association, 25(172):4– 5, December 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278419.

Anonymous:1930:ENC

[Ano30c] Anonymous. Editor’s note on the center of population and point of minimum travel. Journal of the American Statistical Association, 25(172):447–452, December 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277970.

Anonymous:1930:FM

[Ano30d] Anonymous. Front matter. Journal of the American Statistical Association, 25(172):1–2, December 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278417.

Anonymous:1930:MASa

[Ano30e] Anonymous. Members added since December, 1929. Journal of the American Statistical Association, 25(169):101–103, March 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278507.

Anonymous:1930:MASc

[Ano30f] Anonymous. Members added since June, 1930. Journal of the American Statistical Association, 25(171):360–362, September 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278205.

Anonymous:1930:MASb

[Ano30g] Anonymous. Members added since march, 1930. Journal of the American Statistical Association, 25(170):224–228, June 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277640.

Anonymous:1930:MASe

[Ano30h] Anonymous. Members added since September, 1930. Journal of the American Statistical Association, 25(172):477–478, December 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277979.

Anonymous:1930:MASd

[Ano30i] Anonymous. [members of the American Statistical Association]. Journal of the American Statistical Association, 25(172):6–64, December 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278420.

Anonymous:1930:MNa

[Ano30j] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 25(169):93–101, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278506.

Anonymous:1930:MNb

[Ano30k] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 25(170):213–223, June 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277638.

Anonymous:1930:MNc

[Ano30l] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 25(171):347–356, September 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278203.

Anonymous:1930:MNd

[Ano30m] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 25(172):471–476, December 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277977.

Anonymous:1930:PP

[Ano30n] Anonymous. Past presidents. Journal of the American Statistical Association, 25(172):3, December 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278418.

Anonymous:1930:PNa

[Ano30o] Anonymous. Personal notes. Journal of the American Statistical Association, 25(170):223–224, June 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277639.

Anonymous:1930:PNb

[Ano30p] Anonymous. Personal notes. Journal of the American Statistical Association, 25(172):476–477, December 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277978.

Anonymous:1930:PAS

[Ano30q] Anonymous. [publications of the American statistical association]. Journal of the American Statistical Association, 25(172):65–98, December 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278421.

Anonymous:1930:VI

[Ano30r] Anonymous. Volume information. Journal of the American Statistical Association, 25(169):1–9, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278488.

Anonymous:1931:ACA

[Ano31a] Anonymous. Additional committee appointments. Journal of the American Statistical Association, 26(174):213–214, June 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277775.

Anonymous:1931:ASS

[Ano31b] Anonymous. Awards of the social science research council. Journal of the American Statistical Association, 26(175):333–336, September 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277912.

Anonymous:1931:BM

[Ano31c] Anonymous. Back matter. Journal of the American Statistical Association, 26(174):??, June 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277793.

Anonymous:1931:ECA

[Ano31d] Anonymous. Errata: Comments on applications of recently developed theory of small samples. Journal of the American Statistical Association, 26(175):336, September 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277913.

Anonymous:1931:MASa

[Ano31e] Anonymous. Members added since December, 1930. Journal of the American Statistical Association, 26(173):85–87, March 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278267.

Anonymous:1931:MASc

[Ano31f] Anonymous. Members added since June, 1931. Journal of the American Statistical Association, 26(175):348–349, September 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277916.

Anonymous:1931:MASb

[Ano31g] Anonymous. Members added since March, 1931. Journal of the American Statistical Association, 26(174):214–219, June 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277776.

Anonymous:1931:MASd

[Ano31h] Anonymous. Members added since September, 1931. Journal of the American Statistical Association, 26(176):475–476, December 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277953.

Anonymous:1931:MNa

[Ano31i] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 26(173):74–84, March 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278265.

Anonymous:1931:MNb

[Ano31j] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 26(174):204–212, June 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277773.

Anonymous:1931:MNc

[Ano31k] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 26(175):336–347, September 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277914.

Anonymous:1931:PNa

[Ano31l] Anonymous. Personal notes. Journal of the American Statistical Association, 26(173):84–85, March 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278266.

Anonymous:1931:PNb

[Ano31m] Anonymous. Personal notes. Journal of the American Statistical Association, 26(174):212–213, June 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277774.

Anonymous:1931:PNc

[Ano31n] Anonymous. Personal notes. Journal of the American Statistical Association, 26(175):347–348, September 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277915.

Anonymous:1931:PNd

[Ano31o] Anonymous. Personal notes. Journal of the American Statistical Association, 26(176):474–475, December 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277952.

Anonymous:1931:VI

[Ano31p] Anonymous. Volume information. Journal of the American Statistical Association, 26(173):1–9, March 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278252.

Anonymous:1932:BM

[Ano32a] Anonymous. Back matter. Journal of the American Statistical Association, 27(179):??, September 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278187.

Anonymous:1932:CPP

[Ano32b] Anonymous. A constructive program for price statistics. Journal of the American Statistical Association, 27(177):74–78, March 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277884.

Anonymous:1932:CMC

[Ano32c] Anonymous. The costs of medical care. Journal of the American Statistical Association, 27(178):190–191, June 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277668.

Anonymous:1932:EIE

[Ano32d] Anonymous. Errata: An improved equal-frequency map of the normal correlation surface. Journal of the American Statistical Association, 27(180):431, December 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278643.

Anonymous:1932:FGS

[Ano32e] Anonymous. Fellowships and grants of the social science research council, 1932-33. Journal of the American Statistical Association, 27(178):188–190, June 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277667.

Anonymous:1932:ISS

[Ano32f] Anonymous. An intelligence service in the social welfare field. Journal of the American Statistical Association, 27(178):187–188, June 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277666.

Anonymous:1932:MASa

[Ano32g] Anonymous. Members added since December, 1931. Journal of the American Statistical Association, 27(177):99–101, March 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277890.

Anonymous:1932:MASc

[Ano32h] Anonymous. Members added since June, 1932. Journal of the American Statistical Association, 27(179):333–334, September 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278172.

Anonymous:1932:MASb

[Ano32i] Anonymous. Members added since march, 1932. Journal of the American Statistical Association, 27(178):212–214, June 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277675.

Anonymous:1932:MASd

[Ano32j] Anonymous. Members added since September, 1932. Journal of the American Statistical Association, 27(180):444, December 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278649.

Anonymous:1932:MNa

[Ano32k] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 27(178):204–211, June 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277673.

Anonymous:1932:MNb

[Ano32l] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 27(180):438–443, December 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278646.

Anonymous:1932:ON

[Ano32m] Anonymous. Obituary note. Journal of the American Statistical Association, 27(180):444, December 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278648.

Anonymous:1932:PNc

[Ano32n] Anonymous. Personal note. Journal of the American Statistical Association, 27(179):333, September 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278171.

Anonymous:1932:PNa

[Ano32o] Anonymous. Personal notes. Journal of the American Statistical Association, 27(177):99, March 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277889.

Anonymous:1932:PNb

[Ano32p] Anonymous. Personal notes. Journal of the American Statistical Association, 27(178):212, June 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277674.

Anonymous:1932:P

[Ano32q] Anonymous. Personals. Journal of the American Statistical Association, 27(180):443–444, December 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278647.

Anonymous:1932:VI

[Ano32r] Anonymous. Volume information. Journal of the American Statistical Association, 27(177):1–10, March 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277874.

Anonymous:1933:ACA

[Ano33a] Anonymous. Additional committee appointments. Journal of the American Statistical Association, 28(183):356–357, September 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278437.

Anonymous:1933:ASS

[Ano33b] Anonymous. Awards of the social science research council, 1933-34. Journal of the American Statistical Association, 28(182):208–210, June 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278037.

Anonymous:1933:BM

[Ano33c] Anonymous. Back matter. Journal of the American Statistical Association, 28(181):??, March 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278026.

Anonymous:1933:CSB

[Ano33d] Anonymous. The central statistical board. Journal of the American Statistical Association, 28(184):438, December 1933. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278134.

Anonymous:1933:CGS

[Ano33e] Anonymous. Committee on government statistics. Journal of the American Statistical Association, 28(184):439, December 1933. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278135.

Anonymous:1933:CRS

[Ano33f] Anonymous. Committee on reports and statistics of the American public welfare association. Journal of the American Statistical Association, 28(183):336–338, September 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278431.

Anonymous:1933:C

[Ano33g] Anonymous. Committees. Journal of the American Statistical Association, 28(182):235, June 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278045.

Anonymous:1933:MASa

[Ano33h] Anonymous. Members added since December, 1932. Journal of the American Statistical Association, 28(181):101–102, March 1933. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278012.

Anonymous:1933:MASc

[Ano33i] Anonymous. Members added since June, 1933. Journal of the American Statistical Association, 28(183):357, September 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278438.

Anonymous:1933:MASb

[Ano33j] Anonymous. Members added since March, 1933. Journal of the American Statistical Association, 28(182):235–237, June 1933. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278046.

Anonymous:1933:MASd

[Ano33k] Anonymous. Members added since September, 1933. Journal of the American Statistical Association, 28(184):457–458, December 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278141.

Anonymous:1933:MNa

[Ano33l] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 28(181):93–100, March 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278009.

Anonymous:1933:MNb

[Ano33m] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 28(182):223–233, June 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278042.

Anonymous:1933:MNc

[Ano33n] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 28(183):347–355, September 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278435.

Anonymous:1933:MNd

[Ano33o] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 28(184):448–455, December 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278138.

Anonymous:1933:ONa

[Ano33p] Anonymous. Obituary note. Journal of the American Statistical Association, 28(181):100–101, March 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278011.

Anonymous:1933:ONb

[Ano33q] Anonymous. Obituary notes. Journal of the American Statistical Association, 28(182):234–235, June 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278044.

Anonymous:1933:Pa

[Ano33r] Anonymous. Personals. Journal of the American Statistical Association, 28(181):100, March 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278010.

Anonymous:1933:Pb

[Ano33s] Anonymous. Personals. Journal of the American Statistical Association, 28(182):234, June 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278043.

Anonymous:1933:Pc

[Ano33t] Anonymous. Personals. Journal of the American Statistical Association, 28(183):355–356, September 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278436.

Anonymous:1933:Pd

[Ano33u] Anonymous. Personals. Journal of the American Statistical Association, 28(184):455–456, December 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278139.

Anonymous:1933:PWC

[Ano33v] Anonymous. Progress of work in the Census Bureau. Journal of the American Statistical Association, 28(181):91–93, March 1933. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278008.

Anonymous:1933:VI

[Ano33w] Anonymous. Volume information. Journal of the American Statistical Association, 28(181):1–10, March 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277997.

Anonymous:1934:MBC

[Ano34a] Anonymous. The 13-month business calendar. Journal of the American Statistical Association, 29(185):100–104, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277812.

Anonymous:1934:BMa

[Ano34b] Anonymous. Back matter. Journal of the American Statistical Association, 29(185):??, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278487.

Anonymous:1934:BMb

[Ano34c] Anonymous. Back matter. Journal of the American Statistical Association, 29(187):??, September 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278094.

Anonymous:1934:CSB

[Ano34d] Anonymous. The central statistical board. Journal of the American Statistical Association, 29(186):192–194, June 1934. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278293.

Anonymous:1934:CTM

[Ano34e] Anonymous. Collegiate training in mathematical statistics. Journal of the American Statistical Association, 29(188):425, December 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278620.

Anonymous:1934:C

[Ano34f] Anonymous. Committees for 1934. Journal of the American Statistical Association, 29(185):238–240, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277835.

Anonymous:1934:FMa

[Ano34g] Anonymous. Front matter. Journal of the American Statistical Association, 29(185):1–4, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277794.

Anonymous:1934:FMb

[Ano34h] Anonymous. Front matter. Journal of the American Statistical Association, 29(185):??, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278454.

Anonymous:1934:FMc

[Ano34i] Anonymous. Front matter. Journal of the American Statistical Association, 29(186):??, June 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278281.

Anonymous:1934:FMd

[Ano34j] Anonymous. Front matter. Journal of the American Statistical Association, 29(187):??, September 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278060.

Anonymous:1934:FMe

[Ano34k] Anonymous. Front matter. Journal of the American Statistical Association, 29(188):??, December 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278605.

Anonymous:1934:MASa

[Ano34l] Anonymous. Members added since December, 1933. Journal of the American Statistical Association, 29(185):100–102, March 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278472.

Anonymous:1934:MASc

[Ano34m] Anonymous. Members added since June, 1934. Journal of the American Statistical Association, 29(187):337–338, September 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278083.

Anonymous:1934:MASb

[Ano34n] Anonymous. Members added since March, 1934. Journal of the American Statistical Association, 29(186):219–221, June 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278301.

Anonymous:1934:MASd

[Ano34o] Anonymous. Members added since September, 1934. Journal of the American Statistical Association, 29(188):437–438, December 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278624.

Anonymous:1934:MNa

[Ano34p] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 29(187):328–335, September 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278081.

Anonymous:1934:MNb

[Ano34q] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 29(188):428–437, December 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278622.

Anonymous:1934:NFA

[Ano34r] Anonymous. Ninety-Fifth Annual Meeting. Journal of the American Statistical Association, 29(185):1–13, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277795.

Anonymous:1934:O

[Ano34s] Anonymous. Obituaries. Journal of the American Statistical Association, 29(186):219, June 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278300.

Anonymous:1934:Pa

[Ano34t] Anonymous. Personals. Journal of the American Statistical Association, 29(185):99, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278470.

Anonymous:1934:Pb

[Ano34u] Anonymous. Personals. Journal of the American Statistical Association, 29(186):218–219, June 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278299.

Anonymous:1934:Pc

[Ano34v] Anonymous. Personals. Journal of the American Statistical Association, 29(187):336, September 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278082.

Anonymous:1934:Pd

[Ano34w] Anonymous. Personals. Journal of the American Statistical Association, 29(188):437, December 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278623.

Anonymous:1934:PRC

[Ano34x] Anonymous. Procedure for research committees. Journal of the American Statistical Association, 29(188):425, December 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278619.

Anonymous:1934:R

[Ano34y] Anonymous. Reports. Journal of the American Statistical Association, 29(185):198–237, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277834.

Anonymous:1934:SSR

[Ano34z] Anonymous. A statistical slide rule. Journal of the American Statistical Association, 29(187):318, September 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278074.

Anonymous:1934:VI

[Ano34-27] Anonymous. Volume information. Journal of the American Statistical Association, 29(185):1–11, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278453.

Anonymous:1935:BGP

[Ano35a] Anonymous. Better government personnel. Journal of the American Statistical Association, 30(190):446–448, June 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277695.

Anonymous:1935:CSB

[Ano35b] Anonymous. The central statistical board. Journal of the American Statistical Association, 30(192):714–716, December 1935. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277730.

Anonymous:1935:CAa

[Ano35c] Anonymous. Chapter activities. Journal of the American Statistical Association, 30(190):453–456, June 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277698.

Anonymous:1935:CAb

[Ano35d] Anonymous. Chapter activities. Journal of the American Statistical Association, 30(191):601–608, September 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278110.

Anonymous:1935:CAc

[Ano35e] Anonymous. Chapter activities. Journal of the American Statistical Association, 30(192):723–725, December 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277736.

Anonymous:1935:CSS

[Ano35f] Anonymous. Correction: a statistical slide rule. Journal of the American Statistical Association, 30(192):722, December 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277735.

Anonymous:1935:FMa

[Ano35g] Anonymous. Front matter. Journal of the American Statistical Association, 30(189):??, March 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278524.

Anonymous:1935:FMb

[Ano35h] Anonymous. Front matter. Journal of the American Statistical Association, 30(190):??, June 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277687.

Anonymous:1935:FMc

[Ano35i] Anonymous. Front matter. Journal of the American Statistical Association, 30(191):??, September 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278095.

Anonymous:1935:FMd

[Ano35j] Anonymous. Front matter. Journal of the American Statistical Association, 30(192):??, December 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277717.

Anonymous:1935:IMS

[Ano35k] Anonymous. The institute of mathematical statistics. Journal of the American Statistical Association, 30(192):710–711, December 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277728.

Anonymous:1935:LE

[Ano35l] Anonymous. Letters to the editor. Journal of the American Statistical Association, 30(192):718–721, December 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277733.

Anonymous:1935:MN

[Ano35m] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 30(189):106–116, March 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278537.

Anonymous:1935:NMa

[Ano35n] Anonymous. New members. Journal of the American Statistical Association, 30(190):471–472, June 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277701.

Anonymous:1935:NMb

[Ano35o] Anonymous. New members. Journal of the American Statistical Association, 30(191):628–629, September 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278113.

Anonymous:1935:NMc

[Ano35p] Anonymous. New members. Journal of the American Statistical Association, 30(192):754–755, December 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277739.

Anonymous:1935:PN

[Ano35q] Anonymous. Personal notes. Journal of the American Statistical Association, 30(191):627, September 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278112.

Anonymous:1935:P

[Ano35r] Anonymous. Personals. Journal of the American Statistical Association, 30(190):470, June 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277700.

Anonymous:1935:SNNa

[Ano35s] Anonymous. Statistical news and notes. Journal of the American Statistical Association, 30(190):456–470, June 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277699.

Anonymous:1935:SNNb

[Ano35t] Anonymous. Statistical news and notes. Journal of the American Statistical Association, 30(191):608–627, September 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278111.

Anonymous:1935:SNNc

[Ano35u] Anonymous. Statistical news and notes. Journal of the American Statistical Association, 30(192):726–751, December 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277737.

Anonymous:1935:SII

[Ano35v] Anonymous. Suggestions invited for an improved statistical slide rule. Journal of the American Statistical Association, 30(192):722, December 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277734.

Anonymous:1935:VI

[Ano35w] Anonymous. Volume information. Journal of the American Statistical Association, 30(189):i–x, March 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278523.

Anonymous:1936:CAa

[Ano36a] Anonymous. Chapter activities. Journal of the American Statistical Association, 31(193):144–152, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278342.

Anonymous:1936:CAb

[Ano36b] Anonymous. Chapter activities. Journal of the American Statistical Association, 31(194):391–397, June 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278572.

Anonymous:1936:CAc

[Ano36c] Anonymous. Chapter activities. Journal of the American Statistical Association, 31(195):563–577, September 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278390.

Anonymous:1936:CAd

[Ano36d] Anonymous. Chapter activities. Journal of the American Statistical Association, 31(196):739–740, December 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278678.

Anonymous:1936:CNI

[Ano36e] Anonymous. Conference on national income and wealth. Journal of the American Statistical Association, 31(193):130–131, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278338.

Anonymous:1936:CAG

[Ano36f] Anonymous. A cooperative arrangement for general relief statistics. Journal of the American Statistical Association, 31(196):736–737, December 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278676.

Anonymous:1936:CRP

[Ano36g] Anonymous. Cooperative research on price. Journal of the American Statistical Association, 31(193):128–130, March 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278337.

Anonymous:1936:CEE

[Ano36h] Anonymous. A correction: Evaluating the effect of inadequately measured variables in partial correlation analysis. Journal of the American Statistical Association, 31(195):562, September 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278389.

Anonymous:1936:CIC

[Ano36i] Anonymous. A correction: On the influence of classification on the determination of a measurable characteristic. Journal of the American Statistical Association, 31(193):126, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278334.

Anonymous:1936:CPN

[Ano36j] Anonymous. A correction: Proceedings of the Ninety-Seventh Annual Meeting Commodore Hotel, New York City. Journal of the American Statistical Association, 31(194):391, June 1936. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278571.

Anonymous:1936:FMa

[Ano36k] Anonymous. Front matter. Journal of the American Statistical Association, 31(193):xiii–xv, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278316.

Anonymous:1936:FMb

[Ano36l] Anonymous. Front matter. Journal of the American Statistical Association, 31(194):??, June 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278555.

Anonymous:1936:FMc

[Ano36m] Anonymous. Front matter. Journal of the American Statistical Association, 31(195):??, September 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278369.

Anonymous:1936:FMd

[Ano36n] Anonymous. Front matter. Journal of the American Statistical Association, 31(196):??, December 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278662.

Anonymous:1936:GP

[Ano36o] Anonymous. Graphic presentation. Journal of the American Statistical Association, 31(195):560–561, September 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278387.

Anonymous:1936:LE

[Ano36p] Anonymous. Letters to the editors. Journal of the American Statistical Association, 31(195):561–562, September 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278388.

Anonymous:1936:MIM

[Ano36q] Anonymous. Meeting of the institute of mathematical statistics. Journal of the American Statistical Association, 31(193):128, March 1936.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278336.

Anonymous:1936:NMa

[Ano36r] Anonymous. New members. Journal of the American Statistical Association, 31(193):169–172, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278345.

Anonymous:1936:NMb

[Ano36s] Anonymous. New members. Journal of the American Statistical Association, 31(194):415–416, June 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278575.

Anonymous:1936:NMc

[Ano36t] Anonymous. New members. Journal of the American Statistical Association, 31(195):605–607, September 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278393.

Anonymous:1936:NMd

[Ano36u] Anonymous. New members. Journal of the American Statistical Association, 31(196):755–756, December 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278680.

Anonymous:1936:OD

[Ano36v] Anonymous. An occupational dictionary. Journal of the American Statistical Association, 31(194):389–390, June 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278569.

Anonymous:1936:OCS

[Ano36w] Anonymous. Organization of the central statistical board. Journal of the American Statistical Association, 31(194):386–387, June 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278567.

Anonymous:1936:PN

[Ano36x] Anonymous. Personal notes. Journal of the American Statistical Association, 31(193):169, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278344.

Anonymous:1936:PNS

[Ano36y] Anonymous. Proceedings of the Ninety-Seventh Annual Meeting Commodore Hotel, New York City. Journal of the American Statistical Association, 31(193):173–211, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278346.

Anonymous:1936:SNNa

[Ano36z] Anonymous. Statistical news and notes. Journal of the American Statistical Association, 31(193):153–169, March 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278343.

Anonymous:1936:SNNb

[Ano36-27] Anonymous. Statistical news and notes. Journal of the American Statistical Association, 31(194):397–410, June 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278573.

Anonymous:1936:SNNc

[Ano36-28] Anonymous. Statistical news and notes. Journal of the American Statistical Association, 31(195):578–602, September 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278391.

Anonymous:1936:SNNd

[Ano36-29] Anonymous. Statistical news and notes. Journal of the American Statistical Association, 31(196):740–755, December 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278679.

Anonymous:1936:VI

[Ano36-30] Anonymous. Volume information. Journal of the American Statistical Association, 31(193):i–xi, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278315.

Anonymous:1937:AS

[Ano37a] Anonymous. Accident statistics. Journal of the American Statistical Association, 32(200):715–716, December 1937. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279375.

Anonymous:1937:BM

[Ano37b] Anonymous. Back matter. Journal of the American Statistical Association, 32(197):??, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279897.

Anonymous:1937:CAa

[Ano37c] Anonymous. Chapter activities. Journal of the American Statistical Association, 32(197):164–171, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279877.

Anonymous:1937:CAb

[Ano37d] Anonymous. Chapter activities. Journal of the American Statistical Association, 32(198):400–408, June 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279589.

Anonymous:1937:CAc

[Ano37e] Anonymous. Chapter activities. Journal of the American Statistical Association, 32(199):565–577, September 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278777.

Anonymous:1937:CAd

[Ano37f] Anonymous. Chapter activities. Journal of the American Statistical Association, 32(200):746–752, December 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279379.

Anonymous:1937:CPDb

[Ano37g] Anonymous. A correction: The probability distribution of protection test results. Journal of the American Statistical Association, 32(197): 137, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279875.

Anonymous:1937:CPDa

[Ano37h] Anonymous. Correction: The probability distribution of protection test results. Journal of the American Statistical Association, 32(197): iii, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279850.

Anonymous:1937:FAS

[Ano37i] Anonymous. Federal accident statisticians. Journal of the American Statistical Association, 32(198):376–377, June 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279586.

Anonymous:1937:FMa

[Ano37j] Anonymous. Front matter. Journal of the American Statistical Association, 32(197):??, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279851.

Anonymous:1937:FMb

[Ano37k] Anonymous. Front matter. Journal of the American Statistical Association, 32(198):??, June 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279566.

Anonymous:1937:FMc

[Ano37l] Anonymous. Front matter. Journal of the American Statistical Association, 32(199):??, September 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278763.

Anonymous:1937:FMd

[Ano37m] Anonymous. Front matter. Journal of the American Statistical Association, 32(200):??, December 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279366.

Anonymous:1937:MWC

[Ano37n] Anonymous. Methods in wildlife censuses. Journal of the American Statistical Association, 32(199):537–542, September 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278774.

Anonymous:1937:NMa

[Ano37o] Anonymous. New members. Journal of the American Statistical Association, 32(197):172–176, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279878.

Anonymous:1937:NMb

[Ano37p] Anonymous. New members. Journal of the American Statistical Association, 32(198):408–412, June 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279590.

Anonymous:1937:NMc

[Ano37q] Anonymous. New members. Journal of the American Statistical Association, 32(199):577–579, September 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278778.

Anonymous:1937:NMd

[Ano37r] Anonymous. New members. Journal of the American Statistical Association, 32(200):753–754, December 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279381.

Anonymous:1937:PNE

[Ano37s] Anonymous. Proceedings: Ninety-Eighth Annual Meeting: Stevens Hotel, Chicago. Journal of the American Statistical Association, 32(197):177–216, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279879.

Anonymous:1937:SNNa

[Ano37t] Anonymous. Statistical news and notes. Journal of the American Statistical Association, 32(197):138–163, March 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279876.

Anonymous:1937:SNNb

[Ano37u] Anonymous. Statistical news and notes. Journal of the American Statistical Association, 32(198):380–399, June 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279588.

Anonymous:1937:SNNc

[Ano37v] Anonymous. Statistical news and notes. Journal of the American Statistical Association, 32(199):546–564, September 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278776.

Anonymous:1937:SNNd

[Ano37w] Anonymous. Statistical news and notes. Journal of the American Statistical Association, 32(200):721–745, December 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279378.

Anonymous:1937:VI

[Ano37x] Anonymous. Volume information. Journal of the American Statistical Association, 32(197):i–xii, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279849.

Anonymous:1937:WUI

[Ano37y] Anonymous. What use is industry making of research and statistics? Journal of the American Statistical Association, 32(198):368– 371, June 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279582.

Anonymous:1938:AMS

[Ano38a] Anonymous. The annals of mathematical statistics. Journal of the American Statistical Association, 33(203):582–583, September 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279700.

Anonymous:1938:CAa

[Ano38b] Anonymous. Chapter activities. Journal of the American Statistical Association, 33(201):257–264, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279132.

Anonymous:1938:CAb

[Ano38c] Anonymous. Chapter activities. Journal of the American Statistical Association, 33(202):440–451, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278923.

Anonymous:1938:CAc

[Ano38d] Anonymous. Chapter activities. Journal of the American Statistical Association, 33(203):596–602, September 1938. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279705.

Anonymous:1938:CAd

[Ano38e] Anonymous. Chapter activities. Journal of the American Statistical Association, 33(204):738–740, December 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279070.

Anonymous:1938:CDF

[Ano38f] Anonymous. A correction: Discrete frequency distributions arising from mixtures of several single probability values. Journal of the American Statistical Association, 33(203):584, September 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279703.

Anonymous:1938:FMa

[Ano38g] Anonymous. Front matter. Journal of the American Statistical Association, 33(201):??, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279107.

Anonymous:1938:FMb

[Ano38h] Anonymous. Front matter. Journal of the American Statistical Association, 33(202):??, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278903.

Anonymous:1938:FMc

[Ano38i] Anonymous. Front matter. Journal of the American Statistical Association, 33(203):??, September 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279687.

Anonymous:1938:FMd

[Ano38j] Anonymous. Front matter. Journal of the American Statistical Association, 33(204):??, December 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279056.

Anonymous:1938:IMS

[Ano38k] Anonymous. Institute of mathematical statistics. Journal of the American Statistical Association, 33(201):229, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279127.

Anonymous:1938:ISI

[Ano38l] Anonymous. International statistical institute. Journal of the American Statistical Association, 33(201):229, March 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279126.

Anonymous:1938:JPS

[Ano38m] Anonymous. Journal of the Polish statistical society. Journal of the American Statistical Association, 33(203):583–584, September 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279701.

Anonymous:1938:MDC

[Ano38n] Anonymous. A manual of design and construction of time-series charts. Journal of the American Statistical Association, 33(203):584, September 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279702.

Anonymous:1938:NMb

[Ano38o] Anonymous. New member. Journal of the American Statistical Association, 33(202):452–454, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278924.

Anonymous:1938:NMa

[Ano38p] Anonymous. New members. Journal of the American Statistical Association, 33(201):265–268, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279133.

Anonymous:1938:NMc

[Ano38q] Anonymous. New members. Journal of the American Statistical Association, 33(203):603–604, September 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279706.

Anonymous:1938:NMd

[Ano38r] Anonymous. New members. Journal of the American Statistical Association, 33(204):742–743, December 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279072.

Anonymous:1938:PSS

[Ano38s] Anonymous. Polish statistical society. Journal of the American Statistical Association, 33(201):229, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279128.

Anonymous:1938:PNN

[Ano38t] Anonymous. Proceedings: Ninety-Ninth Annual Meeting: Haddon Hall, Atlantic City, Monday, December 27, 1937. Journal of the American Statistical Association, 33(201):179–225, March 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279124.

Anonymous:1938:PRa

[Ano38u] Anonymous. Publications received. Journal of the American Statistical Association, 33(203):650–656, September 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279733.

Anonymous:1938:PRb

[Ano38v] Anonymous. Publications received. Journal of the American Statistical Association, 33(204):792–796, December 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279105.

Anonymous:1938:RBP

[Ano38w] Anonymous. Revision of the base period for government index numbers. Journal of the American Statistical Association, 33(202):411– 415, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278920.

Anonymous:1938:SNNa

[Ano38x] Anonymous. Statistical news and notes. Journal of the American Statistical Association, 33(201):233–256, March 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279131.

Anonymous:1938:SNNb

[Ano38y] Anonymous. Statistical news and notes. Journal of the American Statistical Association, 33(202):418–439, June 1938. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278922.

Anonymous:1938:SNNc

[Ano38z] Anonymous. Statistical news and notes. Journal of the American Statistical Association, 33(203):585–595, September 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279704.

Anonymous:1938:SNNd

[Ano38-27] Anonymous. Statistical news and notes. Journal of the American Statistical Association, 33(204):725–737, December 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279069.

Anonymous:1938:VI

[Ano38-28] Anonymous. Volume information. Journal of the American Statistical Association, 33(201):iii–xii, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279106.

Anonymous:1939:BM

[Ano39a] Anonymous. Back matter. Journal of the American Statistical Association, 34(207):??, September 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279532.

Anonymous:1939:CAa

[Ano39b] Anonymous. Chapter activities. Journal of the American Statistical Association, 34(205):136–141, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279172.

Anonymous:1939:CAb

[Ano39c] Anonymous. Chapter activities. Journal of the American Statistical Association, 34(206):392–404, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278861.

Anonymous:1939:CAc

[Ano39d] Anonymous. Chapter activities. Journal of the American Statistical Association, 34(207):577–582, September 1939. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279490.

Anonymous:1939:CAd

[Ano39e] Anonymous. Chapter activities. Journal of the American Statistical Association, 34(208):744–748, December 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279818.

Anonymous:1939:FMa

[Ano39f] Anonymous. Front matter. Journal of the American Statistical Association, 34(205):??, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279156.

Anonymous:1939:FMb

[Ano39g] Anonymous. Front matter. Journal of the American Statistical Association, 34(206):??, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278840.

Anonymous:1939:FMc

[Ano39h] Anonymous. Front matter. Journal of the American Statistical Association, 34(207):??, September 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279480.

Anonymous:1939:FMd

[Ano39i] Anonymous. Front matter. Journal of the American Statistical Association, 34(208):??, December 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279806.

Anonymous:1939:NMa

[Ano39j] Anonymous. New members. Journal of the American Statistical Association, 34(205):141–143, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279173.

Anonymous:1939:NMb

[Ano39k] Anonymous. New members. Journal of the American Statistical Association, 34(206):404–407, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278862.

Anonymous:1939:NMc

[Ano39l] Anonymous. New members. Journal of the American Statistical Association, 34(207):582–587, September 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279491.

Anonymous:1939:NMd

[Ano39m] Anonymous. New members. Journal of the American Statistical Association, 34(208):748–750, December 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279819.

Anonymous:1939:POH

[Ano39n] Anonymous. Proceedings: One-Hundredth Annual Meeting: Book– Cadillac Hotel, Detroit. Journal of the American Statistical Association, 34(205):144–185, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279174.

Anonymous:1939:PRa

[Ano39o] Anonymous. Publications received. Journal of the American Statistical Association, 34(205):235–238, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279211.

Anonymous:1939:PRb

[Ano39p] Anonymous. Publications received. Journal of the American Statistical Association, 34(206):465–468, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278902.

Anonymous:1939:PRc

[Ano39q] Anonymous. Publications received. Journal of the American Statistical Association, 34(207):636–640, September 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279531.

Anonymous:1939:PRd

[Ano39r] Anonymous. Publications received. Journal of the American Statistical Association, 34(208):784–786, December 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279848.

Anonymous:1939:SNNa

[Ano39s] Anonymous. Statistical news and notes. Journal of the American Statistical Association, 34(205):112–135, March 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279171.

Anonymous:1939:SNNb

[Ano39t] Anonymous. Statistical news and notes. Journal of the American Statistical Association, 34(206):369–391, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278860.

Anonymous:1939:SNNc

[Ano39u] Anonymous. Statistical news and notes. Journal of the American Statistical Association, 34(208):716–743, December 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279817.

Anonymous:1939:VI

[Ano39v] Anonymous. Volume information. Journal of the American Statistical Association, 34(205):v–xvi, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279155.

Arner:1931:BRB

[Arn31] G. B. L. Arner. Book review: [untitled]. Journal of the American Statistical Association, 26(176):496–497, December 1931. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277960.

Arthur:1936:BRB

[Art36] Henry B. Arthur. Book review: [untitled]. Journal of the American Statistical Association, 31(194):436–438, June 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278590.

Arthur:1937:WAI

[Art37] H. B. Arthur. Weighted aggregates and index numbers. Journal of the American Statistical Association, 32(198):362–364, June 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279579.

Arthur:1939:PCG

[Art39] Henry B. Arthur. Prices of consumers’ goods in expansion and recession. Journal of the American Statistical Association, 34(205):41–48, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279160.

Ascher:1937:VPR

[Asc37] Leonard Ascher. Variations in price relative distributions, January 1927 to December 1936. Journal of the American Statistical Association, 32(198):271–280, June 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279569.

Austin:1935:VSC

[Aus35] W. L. Austin. Vital statistics in the Census Bureau. Journal of the American Statistical Association, 30(191):597–599, September 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278107.

Aylesworth:1937:SPF

[Ayl37] Evelyn Aylesworth. A statistical problem in forestry: Allocation of log yarding costs according to log volume. Journal of the American Statistical Association, 32(198):365–368, June 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279581.

Ayres:1938:EFS

[Ayr38] Milan V. Ayres. The economic function of the sales finance company. Journal of the American Statistical Association, 33(201):59–70, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279114.

B:1939:DPI

[B.39] E. C. B. Did productivity increase in the twenties?: Reply. Journal of the American Statistical Association, 34(206):334, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278854.

Black:1937:CFUa

[BA37a] John D. Black and R. H. Allen. The counting of farms in the United States. Journal of the American Statistical Association, 32 (199):439–463, September 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278764.

Black:1937:CFUb

[BA37b] John D. Black and R. H. Allen. The counting of farms in the United States: Discussion. Journal of the American Statistical Association, 32(199):469–470, September 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278766.

Bacon:1931:BRB

[Bac31] Dorothy C. Bacon. Book review: [untitled]. Journal of the American Statistical Association, 26(175):373–374, September 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277932.

Bader:1931:CWF

[Bad31] Louis Bader. Can we find out how the American income is spent? Journal of the American Statistical Association, 26(175):285–294, September 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277906.

Bader:1933:AFI

[Bad33] Louis Bader. The American family income and prosperity. Journal of the American Statistical Association, 28(183):303–311, September 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2278426.

Bailey:1931:AC

[Bai31] A. L. Bailey. The analysis of covariance. Journal of the American Statistical Association, 26(176):424–435, December 1931. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277943.

Baker:1930:SPM

[Bak30] George A. Baker. The significance of the product-moment coefficient of correlation with special reference to the character of the marginal distributions. Journal of the American Statistical Association, 25(172):387–396, December 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277964.

Balderston:1936:BRB

[Bal36] C. Candy Balderston. Book review: [untitled]. Journal of the American Statistical Association, 31(196):813–815, December 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278715.

Barnett:1933:BRB

[Bar33] George E. Barnett. Book review: [untitled]. Journal of the American Statistical Association, 28(181):103–104, March 1933. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278013.

Bartlett:1936:SP

[Bar36] M. S. Bartlett. Statistical probability. Journal of the American Statistical Association, 31(195):553–555, September 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278384.

Barnes:1937:BRB

[Bar37] Irston R. Barnes. Book review: [untitled]. Journal of the American Statistical Association, 32(200):798–799, December 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279402.

Barry:1938:CFR

[Bar38] David C. Barry. Consumer financing and its relation to the commercial bank. Journal of the American Statistical Association, 33(201): 51–58, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279113.

Baten:1935:FFS

[Bat35] William Dowell Baten. A formula for finding the skewness of the combination of two or more samples. Journal of the American Statistical Association, 30(189):95–98, March 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278531.

Bauer:1933:BRB

[Bau33] John Bauer. Book review: [untitled]. Journal of the American Statistical Association, 28(182):254–257, June 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278056.

Baude:1936:BRB

[Bau36] Walter A. Baude. Book review: [untitled]. Journal of the American Statistical Association, 31(194):434–435, June 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278588.

Bean:1930:BRB

[Bea30a] L. H. Bean. Book review: [untitled]. Journal of the American Statistical Association, 25(172):493–495, December 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277986.

Bean:1930:ASM

[Bea30b] Louis H. Bean. Application of a simplified method of correlation to problems in acreage and yield variations. Journal of the American Statistical Association, 25(172):428–439, December 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277968.

Beach:1931:BRB

[Bea31a] W. Edwards Beach. Book review: [untitled]. Journal of the American Statistical Association, 26(175):363–364, September 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277924.

Bean:1931:BRB

[Bea31b] Louis H. Bean. Book review: [untitled]. Journal of the American Statistical Association, 26(176):481–484, December 1931. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277955.

Beal:1934:DMH

[Bea34] Arthur F. Beal. Dispersion in man-hour productivity since 1929. Journal of the American Statistical Association, 29(185):66–71, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277805.

Bean:1935:BRB

[Bea35] Louis H. Bean. Book review: [untitled]. Journal of the American Statistical Association, 30(190):484–487, June 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277706.

Beckhart:1937:BRB

[Bec37] Benjamin Haggott Beckhart. Book review: [untitled]. Journal of the American Statistical Association, 32(198):417–419, June 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279593.

Beck:1938:PCD

[Bec38a] Dorothy Fahs Beck. Problems in the compilation of data on total relief and work program expenditures. Journal of the American Statistical Association, 33(202):353–362, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278911.

Beckhart:1938:BRB

[Bec38b] Benjamin Haggott Beckhart. Book review: [untitled]. Journal of the American Statistical Association, 33(203):626–628, September 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279720.

Bell:1931:BRB

[Bel31] Spurgeon Bell. Book review: [untitled]. Journal of the American Statistical Association, 26(175):368, September 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277927.

Belcher:1936:MCR

[Bel36] Donald R. Belcher. Malcolm Churchill Rorty. Journal of the American Statistical Association, 31(195):603–604, September 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278392.

Benedict:1937:BRB

[Ben37a] M. R. Benedict. Book review: [untitled]. Journal of the American Statistical Association, 32(199):596–600, September 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278789.

Benedict:1937:CFU

[Ben37b] M. R. Benedict. The counting of farms in the United States: Discussion. Journal of the American Statistical Association, 32 (199):463–468, September 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278765.

Berglund:1936:BRB

[Ber36a] Abraham Berglund. Book review: [untitled]. Journal of the American Statistical Association, 31(194):429–430, June 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278585.

Berkson:1936:TGF

[Ber36b] Joseph Berkson. A test for the goodness of fit of a curve graduated to a series of observed means. Journal of the American Statistical Association, 31(196):723–725, December 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278670.

Berkson:1938:SDI

[Ber38] Joseph Berkson. Some difficulties of interpretation encountered in the application of the chi-square test. Journal of the American Statistical Association, 33(203):526–536, September 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279690.

Bezanson:1939:BRB

[Bez39] Anne Bezanson. Book review: [untitled]. Journal of the American Statistical Association, 34(208):774–775, December 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279838.

Bidwell:1934:BRB

[Bid34] P. W. Bidwell. Book review: [untitled]. Journal of the American Statistical Association, 29(185):113–114, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278480.

Bigge:1937:MSR

[Big37] George E. Bigge. Minimum statistical requirements for efficient administration of unemployment compensation systems. Journal of the American Statistical Association, 32(197):19–27, March 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279855.

Blanchard:1933:BRB

[Bla33] Ralph H. Blanchard. Book review: [untitled]. Journal of the American Statistical Association, 28(181):116, March 1933. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278022.

Black:1934:BRB

[Bla34] John D. Black. Book review: [untitled]. Journal of the American Statistical Association, 29(186):236–238, June 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278312.

Blair:1936:BRB

[Bla36] Jacob J. Blair. Book review: [untitled]. Journal of the American Statistical Association, 31(194):441–442, June 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278593.

Black:1937:BRB

[Bla37] John D. Black. Book review: [untitled]. Journal of the American Statistical Association, 32(198):428–429, June 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279598.

Black:1938:BRB

[Bla38] John D. Black. Book review: [untitled]. Journal of the American Statistical Association, 33(202):461–463, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278929.

Bliss:1932:BRB

[Bli32] Charles A. Bliss. Book review: [untitled]. Journal of the American Statistical Association, 27(177):109–110, March 1932. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277895.

Bloch:1931:BRB

[Blo31] Louis Bloch. Book review: [untitled]. Journal of the American Statistical Association, 26(173):105–107, March 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278274.

Blough:1937:BRB

[Blo37] Roy Blough. Book review: [untitled]. Journal of the American Statistical Association, 32(199):609–610, September 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278795.

Berkson:1935:LSR

[BMH35] Joseph Berkson, Thomas B. Magath, and Margaret Hurn. Laboratory standards in relation to chance fluctuations of the erythrocyte count as estimated with the hemocytometer. Journal of the American Statistical Association, 30(190):414–426, June 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277692.

Bober:1938:BRB

[Bob38] M. M. Bober. Book review: [untitled]. Journal of the American Statistical Association, 33(204):761–762, December 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279084.

Bode:1939:BRB

[Bod39] Karl F. Bode. Book review: [untitled]. Journal of the American Statistical Association, 34(207):610–611, September 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279511.

Bogen:1935:BRB

[Bog35] Jules I. Bogen. Book review: [untitled]. Journal of the American Statistical Association, 30(192):763–765, December 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277744.

Bogen:1936:BRB

[Bog36] Jules I. Bogen. Book review: [untitled]. Journal of the American Statistical Association, 31(196):809, December 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278711.

Bogen:1937:BRB

[Bog37] Jules I. Bogen. Book review: [untitled]. Journal of the American Statistical Association, 32(200):774–775, December 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279390.

Bogen:1938:BRB

[Bog38a] Jules I. Bogen. Book review: [untitled]. Journal of the American Statistical Association, 33(203):628, September 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279721.

Bogen:1938:CCM

[Bog38b] Jules I. Bogen. Changed conditions in the marketing of new issues. Journal of the American Statistical Association, 33(201):31–41, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279111.

Bois:1936:BRB

[Boi36] Eugene F. Du Bois. Book review: [untitled]. Journal of the American Statistical Association, 31(196):765–766, December 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278687.

Borton:1938:BRB

[Bor38] Hugh Borton. Book review: [untitled]. Journal of the American Statistical Association, 33(204):774–776, December 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279091.

Bowerman:1935:ECL

[Bow35a] Walter G. Bowerman. An early contribution by life insurance to vital statistics. Journal of the American Statistical Association, 30 (191):584–585, September 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278103.

Bowman:1935:BRB

[Bow35b] Raymond T. Bowman. Book review: [untitled]. Journal of the American Statistical Association, 30(192):768–770, December 1935. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277747.

Bowers:1936:SSA

[Bow36a] Glenn A. Bowers. Setting up state administration for unemployment insurance. Journal of the American Statistical Association, 31(193): 67–72, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278326.

Bowley:1936:ASE

[Bow36b] A. L. Bowley. The application of sampling to economic and sociological problems. Journal of the American Statistical Association, 31(195):474–480, September 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278371.

Boyle:1933:BRB

[Boy33] James E. Boyle. Book review: [untitled]. Journal of the American Statistical Association, 28(184):470–472, December 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278147.

Bratt:1930:BRB

[Bra30] Elmer Clark Bratt. Book review: [untitled]. Journal of the American Statistical Association, 25(171):383–385, September 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278217.

Brandt:1933:AVT

[Bra33a] A. E. Brandt. The analysis of variance in a “2 × s” table with disproportionate frequencies. Journal of the American Statistical Association, 28(182):164–173, June 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278030.

Brandt:1933:TSU

[Bra33b] A. E. Brandt. A test for significance in a unique sample. Journal of the American Statistical Association, 28(184):434–437, December 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278133.

Bratt:1934:WAR

[Bra34] Elmer C. Bratt. On weighting the activity ratio. Journal of the American Statistical Association, 29(186):188–189, June 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278291.

Brand:1936:LAS

[Bra36a] Charles J. Brand. Legal aspects of statistical activities of trade associations with special reference to the sugar institute decision. Journal of the American Statistical Association, 31(194):367–375, June 1936.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278564.

Brand:1936:SWN

[Bra36b] Charles J. Brand. The statistical work of the national fertilizer association. Journal of the American Statistical Association, 31 (195):552–553, September 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278383.

Brady:1938:VFL

[Bra38] Dorothy S. Brady. Variations in family living expenditures. Journal of the American Statistical Association, 33(202):385–389, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278915.

Bratt:1939:DPI

[Bra39] Elmer C. Bratt. Did productivity increase in the twenties? Journal of the American Statistical Association, 34(206):326–332, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278852.

Bremer:1939:BRB

[Bre39] C. Daniel Bremer. Book review: [untitled]. Journal of the American Statistical Association, 34(208):770–771, December 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279836.

Britten:1931:BRB

[Bri31] Rollo H. Britten. Book review: [untitled]. Journal of the American Statistical Association, 26(175):357–358, September 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277920.

Brissenden:1935:BRB

[Bri35] P. F. Brissenden. Book review: [untitled]. Journal of the American Statistical Association, 30(190):493–494, June 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277711.

Brown:1931:BRB

[Bro31] Douglass V. Brown. Book review: [untitled]. Journal of the American Statistical Association, 26(174):238–239, June 1931. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277789.

Browne:1936:BRB

[Bro36] H. F. Browne. Book review: [untitled]. Journal of the American Statistical Association, 31(195):650–651, September 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278415.

Brown:1938:BRB

[Bro38] J. Douglas Brown. Book review: [untitled]. Journal of the American Statistical Association, 33(203):644–645, September 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279728.

Brown:1939:BRB

[Bro39] Douglass V. Brown. Book review: [untitled]. Journal of the American Statistical Association, 34(206):464–465, June 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278901.

Brunsman:1935:BRB

[Bru35] Howard G. Brunsman. Book review: [untitled]. Journal of the American Statistical Association, 30(192):773–774, December 1935. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277750.

Brumbaugh:1936:BRB

[Bru36] M. A. Brumbaugh. Book review: [untitled]. Journal of the American Statistical Association, 31(193):238–239, March 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278363.

Brumbaugh:1937:BRB

[Bru37] M. A. Brumbaugh. Book review: [untitled]. Journal of the American Statistical Association, 32(200):809–811, December 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279409.

Bean:1932:UTR

[BT32] L. H. Bean and G. B. Thorne. The use of “trends in residuals” in constructing demand curves. Journal of the American Statistical Association, 27(177):61–67, March 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277881.

Buchanan:1938:BRB

[Buc38a] Norman S. Buchanan. Book review: [untitled]. Journal of the American Statistical Association, 33(204):764–766, December 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279086.

Buck:1938:BRB

[Buc38b] Solon J. Buck. Book review: [untitled]. Journal of the American Statistical Association, 33(202):499–501, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278950.

Buehler:1938:BRBa

[Bue38a] Alfred G. Buehler. Book review: [untitled]. Journal of the American Statistical Association, 33(201):294–296, March 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279149.

Buehler:1938:BRBb

[Bue38b] Alfred G. Buehler. Book review: [untitled]. Journal of the American Statistical Association, 33(202):458–459, June 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278927.

Buehler:1938:BRBc

[Bue38c] Alfred G. Buehler. Book review: [untitled]. Journal of the American Statistical Association, 33(204):780, December 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279095.

Burgess:1930:BRB

[Bur30] Robert W. Burgess. Book review: [untitled]. Journal of the American Statistical Association, 25(172):504, December 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277996.

Burns:1931:BRB

[Bur31] Arthur F. Burns. Book review: [untitled]. Journal of the American Statistical Association, 26(173):88–96, March 1931. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278268.

Burch:1932:BRB

[Bur32a] Guy Irving Burch. Book review: [untitled]. Journal of the American Statistical Association, 27(179):345–346, September 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278179.

Burgess:1932:WSC

[Bur32b] May Ayres Burgess. What statistics contribute to public health. Journal of the American Statistical Association, 27(178):193–195, June 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277670.

Burks:1933:SME

[Bur33] Barbara S. Burks. A statistical method for estimating the distribution of sizes of completed fraternities in a population represented by a random sampling of individuals. Journal of the American Statistical Association, 28(184):388–394, December 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278128.

Burgess:1936:BRB

[Bur36a] Robert W. Burgess. Book review: [untitled]. Journal of the American Statistical Association, 31(196):795–798, December 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278704.

Burns:1936:BRB

[Bur36b] Arthur Robert Burns. Book review: [untitled]. Journal of the American Statistical Association, 31(196):779–780, December 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278696.

Burgess:1937:CTH

[Bur37a] May Ayres Burgess. The construction of two height charts. Journal of the American Statistical Association, 32(198):290–310, June 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279571.

Burgess:1937:WDS

[Bur37b] Robert W. Burgess. The whole duty of the statistical forecaster. Journal of the American Statistical Association, 32(200):636–642, December 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279368.

Burgess:1938:SEG

[Bur38] W. Randolph Burgess. The statistical equivalent of gold. Journal of the American Statistical Association, 33(201):1–11, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279108.

Burton:1939:BRB

[Bur39] F. W. Burton. Book review: [untitled]. Journal of the American Statistical Association, 34(207):624–625, September 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279525.

Cadmore:1935:ESS

[Cad35] S. A. Cadmore. The encyclopaedia of the social sciences. Journal of the American Statistical Association, 30(190):449–450, June 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277696.

Campbell:1931:BCA

[Cam31] C. D. Campbell. The business cycle and accidents to railroad employees in the United States. Journal of the American Statistical Association, 26(175):295–302, September 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277907.

Camp:1933:KPM

[Cam33] Burton H. Camp. Karl Pearson and mathematical statistics. Journal of the American Statistical Association, 28(184):395–401, December 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2278129.

Camp:1938:BRB

[Cam38a] Burton H. Camp. Book review: [untitled]. Journal of the American Statistical Association, 33(202):480–483, June 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278940.

Camp:1938:FIC

[Cam38b] Burton H. Camp. Further interpretations of the chi-square test. Journal of the American Statistical Association, 33(203):537–542, September 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279691.

Canning:1939:BRB

[Can39] John B. Canning. Book review: [untitled]. Journal of the American Statistical Association, 34(208):757–758, December 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279825.

Carpenter:1930:BRB

[Car30] Niles Carpenter. Book review: [untitled]. Journal of the American Statistical Association, 25(172):491–493, December 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277985.

Carver:1931:BRB

[Car31] Harry C. Carver. Book review: [untitled]. Journal of the American Statistical Association, 26(175):350–353, September 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277917.

Carmichael:1934:DRP

[Car34] F. L. Carmichael. Denver real property studies. Journal of the American Statistical Association, 29(185):111–117, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277814.

Carroll:1935:BRB

[Car35] Mollie Ray Carroll. Book review: [untitled]. Journal of the American Statistical Association, 30(190):496, June 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277713.

Carmichael:1936:BRB

[Car36] F. L. Carmichael. Book review: [untitled]. Journal of the American Statistical Association, 31(193):234–235, March 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278360.

Carson:1938:TPE

[Car38] William J. Carson. Trends of principal earning assets and their significance. Journal of the American Statistical Association, 33(202):311– 318, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278905.

Cavin:1937:BRB

[Cav37] J. P. Cavin. Book review: [untitled]. Journal of the American Statistical Association, 32(200):787–789, December 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279396.

Cox:1936:CGA

[CB36] A. B. Cox and L. H. Bean. The A.A.A., the cotton growers, and the agricultural problem. Journal of the American Statistical Association, 31(194):295–317, June 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278559.

Cowles3rd:1935:SSC

[CC35] Alfred Cowles 3rd and Edward N. Chapman. A statistical study of climate in relation to pulmonary tuberculosis. Journal of the American Statistical Association, 30(191):517–536, September 1935. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278097.

Cohen:1936:SVN

[CD36] Morris R. Cohen and C. J. Ducasse. The statistical view of nature. Journal of the American Statistical Association, 31(194):327– 347, June 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278561.

Cutts:1937:RMC

[CD37] Jesse M. Cutts and Samuel J. Dennis. Revised method of calculation of the wholesale price index of the United States Bureau of Labor Statistics. Journal of the American Statistical Association, 32(200):663–674, December 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279371.

Cohen:1939:ETS

[CF39] Jerome B. Cohen and John M. Firestone. On an experiment in the teaching of statistics. Journal of the American Statistical Association, 34(208):714–715, December 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279816.

Chaddock:1930:BRB

[Cha30a] Robert E. Chaddock. Book review: [untitled]. Journal of the American Statistical Association, 25(170):236–238, June 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277646.

Chapin:1930:IM

[Cha30b] F. Stuart Chapin. An institution in the making. Journal of the American Statistical Association, 25(172):464–465, December 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277973.

Chaddock:1931:CPC

[Cha31] Robert E. Chaddock. Composition of the population of continental United States. Journal of the American Statistical Association, 26(176):449–454, December 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277946.

Chaddock:1933:BRB

[Cha33a] Robert E. Chaddock. Book review: [untitled]. Journal of the American Statistical Association, 28(183):375–376, September 1933. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278452.

Chaddock:1933:WSI

[Cha33b] Robert E. Chaddock. Who shall inherit america? Journal of the American Statistical Association, 28(183):342–345, September 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278433.

Chaddock:1934:SIM

[Cha34a] Robert E. Chaddock. Significance of infant mortality rates for small geographic areas. Journal of the American Statistical Association, 29(187):243–249, September 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278061.

Chamberlin:1934:BRB

[Cha34b] Edward Chamberlin. Book review: [untitled]. Journal of the American Statistical Association, 29(188):445–447, December 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278629.

Chapin:1935:BRB

[Cha35a] F. Stuart Chapin. Book review: [untitled]. Journal of the American Statistical Association, 30(191):631–632, September 1935. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278115.

Chawner:1935:CCI

[Cha35b] Lowell J. Chawner. Construction cost indexes as influenced by technological change and other factors. Journal of the American Statistical Association, 30(191):561–576, September 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278100.

Chalmers:1937:SDR

[Cha37] W. Ellison Chalmers. Statistical discussions and recommendations of the world textile conference. Journal of the American Statistical Association, 32(198):371–373, June 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279583.

Chawner:1938:BRB

[Cha38] Lowell J. Chawner. Book review: [untitled]. Journal of the American Statistical Association, 33(204):747–748, December 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279074.

Chaddock:1939:BRB

[Cha39] Robert E. Chaddock. Book review: [untitled]. Journal of the American Statistical Association, 34(206):439–441, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278881.

Cheney:1934:CAC

[Che34] Roy A. Cheney. The code authority in cooperation with NRA or a government bureau. Journal of the American Statistical Association, 29(185):177–179, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277829.

Ciocco:1937:BRB

[Cio37] Antonio Ciocco. Book review: [untitled]. Journal of the American Statistical Association, 32(200):819–820, December 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279415.

Ciocco:1938:BRB

[Cio38] Antonio Ciocco. Book review: [untitled]. Journal of the American Statistical Association, 33(202):484–485, June 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278942.

Ciocco:1939:BRB

[Cio39] Antonio Ciocco. Book review: [untitled]. Journal of the American Statistical Association, 34(207):588–589, September 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279493.

Conrad:1932:FSD

[CJ32] Herbert S. Conrad and Harold E. Jones. A field study of the differential birth rate. Journal of the American Statistical Association, 27(178):153–159, June 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277656.

Chapin:1933:PRC

[CJS33] F. Stuart Chapin, Ernst Jacobson, and Sarah Stone. Predicting relief case loads for Minneapolis by empirical procedures, 1932–1933. Journal of the American Statistical Association, 28(184):414–422, December 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278131.

Clark:1931:BRB

[Cla31] J. M. Clark. Book review: [untitled]. Journal of the American Statistical Association, 26(173):101–103, March 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278271.

Clark:1932:BRB

[Cla32] W. C. Clark. Book review: [untitled]. Journal of the American Statistical Association, 27(179):341–343, September 1932. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278176.

Clark:1933:BRB

[Cla33] J. M. Clark. Book review: [untitled]. Journal of the American Statistical Association, 28(183):360–362, September 1933. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278440.

Clague:1934:SPC

[Cla34a] Ewan Clague. Some problems in the construction of an index number of rents: Discussion. Journal of the American Statistical Association, 29(185):29–32, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277798.

Clark:1934:FMI

[Cla34b] J. M. Clark. Factors making for instability. Journal of the American Statistical Association, 29(185):72–74, March 1934. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277806.

Clague:1937:SEP

[Cla37] Ewan Clague. Statistical and economic problems in the administration of social security. Journal of the American Statistical Association, 32(199):509–516, September 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278770.

Clark:1939:BRB

[Cla39] J. M. Clark. Book review: [untitled]. Journal of the American Statistical Association, 34(206):431–432, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278876.

Cline:1939:BRB

[Cli39] Denzel C. Cline. Book review: [untitled]. Journal of the American Statistical Association, 34(206):452–454, June 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278891.

Coats:1931:BRB

[Coa31a] R. H. Coats. Book review: [untitled]. Journal of the American Statistical Association, 26(174):226–229, June 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277780.

Coats:1931:ESF

[Coa31b] R. H. Coats. Enumeration and sampling in the field of the census. Journal of the American Statistical Association, 26(175):270–284, September 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277905.

Coats:1939:SS

[Coa39] R. H. Coats. Science and society. Journal of the American Statistical Association, 34(205):1–26, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279157.

Cochran:1939:UAV

[Coc39] W. G. Cochran. The use of the analysis of variance in enumeration by sampling. Journal of the American Statistical Association, 34(207):492–510, September 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279483.

Cohen:1938:MS

[Coh38] Jerome B. Cohen. The misuse of statistics. Journal of the American Statistical Association, 33(204):657–674, December 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279057.

Colm:1937:BRB

[Col37] Gerhard Colm. Book review: [untitled]. Journal of the American Statistical Association, 32(197):233–235, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279889.

Colm:1938:BRB

[Col38] Gerhard Colm. Book review: [untitled]. Journal of the American Statistical Association, 33(203):624–626, September 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279719.

Commons:1933:BRB

[Com33] John R. Commons. Book review: [untitled]. Journal of the American Statistical Association, 28(183):358–360, September 1933. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278439.

Compton:1934:TIC

[Com34] Wilson Compton. The trade institutes and code authorities. Journal of the American Statistical Association, 29(185):171–176, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277828.

Copeland:1930:BRB

[Cop30] Morris A. Copeland. Book review: [untitled]. Journal of the American Statistical Association, 25(171):377–379, September 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278212.

Chesire:1932:FES

[COP32a] Leone Chesire, Elena Oldis, and Egon S. Pearson. Further experiments on the sampling distribution of the correlation coefficient. Journal of the American Statistical Association, 27(178):121–128, June 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277653.

Copeland:1932:BRB

[Cop32b] Morris A. Copeland. Book review: [untitled]. Journal of the American Statistical Association, 27(180):456, December 1932. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278657.

Copeland:1934:WSC

[Cop34] Morris A. Copeland. Who should collect economic statistics for recovery and planning?:discussion. Journal of the American Statistical Association, 29(185):184–185, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277831.

Copeland:1935:NWI

[Cop35] Morris A. Copeland. National wealth and income — an interpretation. Journal of the American Statistical Association, 30(190):377– 386, June 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277689.

Copeland:1936:CWP

[Cop36] Morris A. Copeland. Changes in the wholesale price index in relation to factory employment: Discussion. Journal of the American Statistical Association, 31(195):503–505, September 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278375.

Corstvet:1934:BRB

[Cor34] Emma Corstvet. Book review: [untitled]. Journal of the American Statistical Association, 29(186):230–231, June 1934. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278308.

Court:1930:MJC

[Cou30] Andrew T. Court. Measuring joint causation. Journal of the American Statistical Association, 25(171):245–254, September 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278188.

Cover:1937:BRB

[Cov37] John H. Cover. Book review: [untitled]. Journal of the American Statistical Association, 32(198):433–434, June 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279602.

Cover:1938:BRB

[Cov38] John H. Cover. Book review: [untitled]. Journal of the American Statistical Association, 33(201):282, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279141.

Cowan:1932:NCP

[Cow32] Donald R. G. Cowan. A note on the coefficient of part correlation and of correlation of a dependent variable with all but one of a group of other variables. Journal of the American Statistical Association, 27(178):177–179, June 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277660.

Cowan:1937:IOS

[Cow37] Donald R. G. Cowan. Improved objective standards of sales performance. Journal of the American Statistical Association, 32(197):

75–82, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279864.

Cox:1936:BRB

[Cox36] A. B. Cox. Book review: [untitled]. Journal of the American Statistical Association, 31(195):638–640, September 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278409.

Cox:1938:BRB

[Cox38] Reavis Cox. Book review: [untitled]. Journal of the American Statistical Association, 33(203):646–647, September 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279730.

Cox:1939:BRBa

[Cox39a] A. B. Cox. Book review: [untitled]. Journal of the American Statistical Association, 34(205):228–229, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279205.

Cox:1939:BRBb

[Cox39b] Reavis Cox. Book review: [untitled]. Journal of the American Statistical Association, 34(207):599–600, September 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279501.

Cox:1939:BRBc

[Cox39c] Reavis Cox. Book review: [untitled]. Journal of the American Statistical Association, 34(207):601–602, September 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279503.

Collins:1934:EDS

[CP34] Selwyn D. Collins and G. St. J. Perrott. The economic depression and sickness. Journal of the American Statistical Association, 29(185):47– 51, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277802.

Craig:1935:BRB

[Cra35] Allen T. Craig. Book review: [untitled]. Journal of the American Statistical Association, 30(190):478–480, June 1935. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277703.

Crathorne:1936:BRB

[Cra36] A. R. Crathorne. Book review: [untitled]. Journal of the American Statistical Association, 31(194):417–418, June 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278576.

Crathorne:1937:BRB

[Cra37] A. R. Crathorne. Book review: [untitled]. Journal of the American Statistical Association, 32(197):217–218, March 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279880.

Cressey:1938:BRB

[Cre38] George B. Cressey. Book review: [untitled]. Journal of the American Statistical Association, 33(203):647–649, September 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279731.

Croxton:1930:BRB

[Cro30] Frederick E. Croxton. Book review: [untitled]. Journal of the American Statistical Association, 25(170):241, June 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277650.

Cross:1932:BRB

[Cro32a] Cecil F. Cross. Book review: [untitled]. Journal of the American Statistical Association, 27(178):220–222, June 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277678.

Croxton:1932:GCB

[Cro32b] Frederick E. Croxton. Graphic comparisons by bars, squares, circles, and cubes. Journal of the American Statistical Association, 27(177): 54–60, March 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277880.

Croxton:1933:BRB

[Cro33] Frederick E. Croxton. Book review: [untitled]. Journal of the American Statistical Association, 28(183):372, September 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278449.

Croxton:1936:BRB

[Cro36] Frederick E. Croxton. Book review: [untitled]. Journal of the American Statistical Association, 31(195):625–627, September 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278402.

Crosby:1938:BRB

[Cro38a] Edwin L. Crosby. Book review: [untitled]. Journal of the American Statistical Association, 33(204):786–787, December 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279100.

Croxton:1938:BRBa

[Cro38b] Frederick E. Croxton. Book review: [untitled]. Journal of the American Statistical Association, 33(201):278–279, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279138.

Croxton:1938:BRBb

[Cro38c] Frederick E. Croxton. Book review: [untitled]. Journal of the American Statistical Association, 33(202):483–484, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278941.

Crum:1931:BRB

[Cru31] W. L. Crum. Book review: [untitled]. Journal of the American Statistical Association, 26(176):477–481, December 1931. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277954.

Crum:1933:AIS

[Cru33] W. L. Crum. On analytical interpretation of straw-vote samples. Journal of the American Statistical Association, 28(182):152–163, June 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278029.

Crum:1935:BRB

[Cru35a] W. L. Crum. Book review: [untitled]. Journal of the American Statistical Association, 30(190):473–478, June 1935. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277702.

Crum:1935:NID

[Cru35b] W. L. Crum. The national income and its distribution. Journal of the American Statistical Association, 30(189):35–46, March 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278527.

Crum:1936:BRB

[Cru36] W. L. Crum. Book review: [untitled]. Journal of the American Statistical Association, 31(196):767–770, December 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278690.

Cotterman:1939:TSM

[CS39] Chas. W. Cotterman and Laurence H. Snyder. Tests of simple Mendelian inheritance in randomly collected data of one and two generations. Journal of the American Statistical Association, 34 (207):511–523, September 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/ stable/2279484. See note [Man45].

Campbell:1939:BPI

[CSS39] F. L. Campbell, George W. Snedecor, and W. A. Simanton. Biostatistical problems involved in the standardization of liquid household insecticides. Journal of the American Statistical Association, 34(205): 62–70, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279163.

Cunningham:1933:BRB

[Cun33] William J. Cunningham. Book review: [untitled]. Journal of the American Statistical Association, 28(183):366–369, September 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278445.

Currie:1938:EDD

[Cur38] Lauchlin Currie. The economic distribution of demand deposits. Journal of the American Statistical Association, 33(202):319–326, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278906.

Cusick:1936:BRB

[Cus36] James F. Cusick. Book review: [untitled]. Journal of the American Statistical Association, 31(193):223–224, March 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278352.

Cusick:1939:BRB

[Cus39] James F. Cusick. Book review: [untitled]. Journal of the American Statistical Association, 34(205):203–204, March 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279188.

D:1936:PWC

[D.36] C. L. D. Progress of work in the Census Bureau: The 1935 census of business. Journal of the American Statistical Association, 31(196):737–738, December 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278677.

D:1937:BRB

[D.37] W. E. D. Book review: [untitled]. Journal of the American Statistical Association, 32(199):588, September 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278784.

Dameron:1937:BRB

[Dam37] Kenneth Dameron. Book review: [untitled]. Journal of the American Statistical Association, 32(200):792–794, December 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279399.

Daugherty:1935:BRB

[Dau35] Carroll R. Daugherty. Book review: [untitled]. Journal of the American Statistical Association, 30(190):496–498, June 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277714.

Daugherty:1936:BRB

[Dau36] Carroll R. Daugherty. Book review: [untitled]. Journal of the American Statistical Association, 31(195):648–650, September 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278414.

Daugherty:1939:BRB

[Dau39] Carroll R. Daugherty. Book review: [untitled]. Journal of the American Statistical Association, 34(208):771–773, December 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279837.

Davenport:1930:BRB

[Dav30a] D. H. Davenport. Book review: [untitled]. Journal of the American Statistical Association, 25(172):495–496, December 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277988.

Davies:1930:FMC

[Dav30b] George R. Davies. First moment correlation. Journal of the American Statistical Association, 25(172):413–427, December 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277967.

Davenport:1932:BRB

[Dav32] D. H. Davenport. Book review: [untitled]. Journal of the American Statistical Association, 27(178):222–223, June 1932. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277679.

Davis:1933:DAA

[Dav33a] I. G. Davis. A discussion of the accuracy of agricultural census enumeration in the northeast. Journal of the American Statistical Association, 28(183):272–285, September 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278423.

Davis:1933:BRB

[Dav33b] Joseph S. Davis. Book review: [untitled]. Journal of the American Statistical Association, 28(182):239–241, June 1933. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278048.

Davenport:1934:BRB

[Dav34a] Donald Davenport. Book review: [untitled]. Journal of the American Statistical Association, 29(186):228–229, June 1934. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278306.

Davis:1934:BRB

[Dav34b] Joseph S. Davis. Book review: [untitled]. Journal of the American Statistical Association, 29(185):103, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278473.

Davis:1937:BRB

[Dav37a] H. S. Davis. Book review: [untitled]. Journal of the American Statistical Association, 32(197):240–242, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279893.

Davis:1937:SSE

[Dav37b] Joseph S. Davis. Statistics and social engineering. Journal of the American Statistical Association, 32(197):1–7, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279852.

Davis:1939:FAR

[Dav39] Shelby Cullom Davis. The financial aspects of railroad consolidation. Journal of the American Statistical Association, 34(206):269– 276, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278844.

Dublin:1931:MN

[DC31] Louis I. Dublin and Robert E. Chaddock. Miscellaneous notes. Journal of the American Statistical Association, 26(176):468–474, December 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277951.

deChazeau:1934:BRB

[dC34] Melvin G. de Chazeau. Book review: [untitled]. Journal of the American Statistical Association, 29(188):447–449, December 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278630.

deChazeau:1937:BRB

[dC37a] Melvin G. de Chazeau. Book review: [untitled]. Journal of the American Statistical Association, 32(198):432–433, June 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279601.

DuBrul:1937:SRS

[DC37b] Stephen M. DuBrul and Ewan Clague. Statistical requirements of social security programs. Journal of the American Statistical Association, 32(197):32–34, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279857.

Deardorff:1930:BRB

[Dea30] Neva R. Deardorff. Book review: [untitled]. Journal of the American Statistical Association, 25(171):366–369, September 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278207.

Deardorff:1932:BRB

[Dea32] Neva R. Deardorff. Book review: [untitled]. Journal of the American Statistical Association, 27(179):339–341, September 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278175.

Dean:1937:SCC

[Dea37] Joel Dean. Statistical cost curves. Journal of the American Statistical Association, 32(197):83–89, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279865.

Deming:1934:CTC

[Dem34] W. Edwards Deming. The chi-test and curve fitting. Journal of the American Statistical Association, 29(188):372–382, December 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278607.

Deming:1936:BRB

[Dem36a] W. Edwards Deming. Book review: [untitled]. Journal of the American Statistical Association, 31(195):622–623, September 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278400.

Deming:1936:PIP

[Dem36b] W. Edwards Deming. The present interests of physicists in statistics. Journal of the American Statistical Association, 31(193):124– 126, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278333.

Deming:1937:BRB

[Dem37] W. Edwards Deming. Book review: [untitled]. Journal of the American Statistical Association, 32(199):587–588, September 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278783.

Deming:1938:STC

[Dem38] W. Edwards Deming. Some thoughts on curve fitting and the chi test. Journal of the American Statistical Association, 33(203):543–551, September 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279692.

Deming:1939:BRB

[Dem39] W. Edwards Deming. Book review: [untitled]. Journal of the American Statistical Association, 34(206):421–422, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278867.

Densen:1936:PSR

[Den36] Paul M. Densen. Price stability and responsiveness to changes in the price of gold. Journal of the American Statistical Association, 31(193):85–87, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278328.

Densen:1937:BRB

[Den37] Paul M. Densen. Book review: [untitled]. Journal of the American Statistical Association, 32(199):588–591, September 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278785.

Derksen:1939:BRB

[Der39] J. B. D. Derksen. Book review: [untitled]. Journal of the American Statistical Association, 34(208):762–763, December 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279829.

Despres:1938:BRB

[Des38] Emile Despres. Book review: [untitled]. Journal of the American Statistical Association, 33(201):296–298, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279150.

Dewhurst:1935:BRB

[Dew35] Frederic Dewhurst. Book review: [untitled]. Journal of the American Statistical Association, 30(191):646–648, September 1935. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278125.

Dewhurst:1936:BRB

[Dew36] Frederic Dewhurst. Book review: [untitled]. Journal of the American Statistical Association, 31(196):785–787, December 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278699.

Dewhurst:1938:BRB

[Dew38] J. Frederic Dewhurst. Book review: [untitled]. Journal of the American Statistical Association, 33(201):284–287, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279143.

Durand:1937:MER

[DF37] E. Dana Durand and Frank Whitson Fetter. Measurement of effects of reciprocal trade agreements. Journal of the American Statistical Association, 32(197):50–61, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279861.

Dickinson:1935:CSB

[Dic35] John Dickinson. The central statistical board. Journal of the American Statistical Association, 30(191):590–593, September 1935. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278105.

Dickens:1938:IDB

[Dic38] Paul D. Dickens. Interest and dividends in the balance of international payments. Journal of the American Statistical Association, 33(204):675–686, December 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279058.

Dietrich:1936:BRB

[Die36] Ethel B. Dietrich. Book review: [untitled]. Journal of the American Statistical Association, 31(193):224–225, March 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278353.

Dwyer:1937:PCT

[DM37] Paul S. Dwyer and Alan D. Meacham. The preparation of correlation tables on a tabulator equipped with digit selection. Journal of the American Statistical Association, 32(200):654–662, December 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279370.

Dodd:1936:BRB

[Dod36] Edward L. Dodd. Book review: [untitled]. Journal of the American Statistical Association, 31(196):758–759, December 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278682.

Dodge:1937:ASM

[Dod37] H. F. Dodge. Application of statistical methods in the American society for testing materials. Journal of the American Statistical Association, 32(198):373–374, June 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279584.

Dodd:1938:BRB

[Dod38] Paul A. Dodd. Book review: [untitled]. Journal of the American Statistical Association, 33(202):501–504, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278951.

Dolley:1939:BRB

[Dol39] James C. Dolley. Book review: [untitled]. Journal of the American Statistical Association, 34(207):621–622, September 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279522.

Dorn:1932:EAN

[Dor32] Harold F. Dorn. The effect of allocation of non-resident deaths upon official mortality statistics. Journal of the American Statistical Association, 27(180):401–412, December 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278639.

Dorn:1939:NDP

[Dor39] Harold F. Dorn. The natural decrease of population in certain American communities. Journal of the American Statistical Association, 34(205):106–109, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279168.

Douglas:1930:BRB

[Dou30] Paul H. Douglas. Book review: [untitled]. Journal of the American Statistical Association, 25(169):117–118, March 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278513.

Douty:1932:TID

[Dou32] H. M. Douty. The trend of industrial disputes, 1922–1930. Journal of the American Statistical Association, 27(178):168–172, June 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277658.

Dougherty:1938:BRB

[Dou38a] Carroll R. Dougherty. Book review: [untitled]. Journal of the American Statistical Association, 33(203):645–646, September 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279729.

Douglas:1938:BRB

[Dou38b] Paul H. Douglas. Book review: [untitled]. Journal of the American Statistical Association, 33(204):760–761, December 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279083.

Downes:1931:AOT

[Dow31] Jean Downes. The accuracy of official tuberculosis death rates. Journal of the American Statistical Association, 26(176):393–406, December 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277940.

Dowrie:1938:BRB

[Dow38] George W. Dowrie. Book review: [untitled]. Journal of the American Statistical Association, 33(202):463–465, June 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278930.

deRoover:1937:BRB

[dR37] Raymond de Roover. Book review: [untitled]. Journal of the American Statistical Association, 32(199):591–592, September 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278786.

Dresch:1939:BRBb

[Dre39a] F. W. Dresch. Book review: [untitled]. Journal of the American Statistical Association, 34(208):768–769, December 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279834.

Dresch:1939:BRBa

[Dre39b] Francis W. Dresch. Book review: [untitled]. Journal of the American Statistical Association, 34(207):596–597, September 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279499.

Drummond:1936:BRB

[Dru36] W. M. Drummond. Book review: [untitled]. Journal of the American Statistical Association, 31(193):225–227, March 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278354.

Drummond:1938:BRB

[Dru38] W. M. Drummond. Book review: [untitled]. Journal of the American Statistical Association, 33(204):782–785, December 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279098.

Dorn:1933:CDM

[DS33] Harold F. Dorn and Samuel A. Stouffer. Criteria of differential mortality. Journal of the American Statistical Association, 28 (184):402–413, December 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278130.

Dublin:1931:BRB

[Dub31] Louis I. Dublin. Book review: [untitled]. Journal of the American Statistical Association, 26(174):220–221, June 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277777.

DuBrul:1935:BRB

[DuB35] S. M. DuBrul. Book review: [untitled]. Journal of the American Statistical Association, 30(192):775–776, December 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277751.

DuBrul:1936:PIP

[DuB36] S. M. DuBrul. A proposal for an integrated program for the reporting of business statistics to government agencies. Journal of the American Statistical Association, 31(193):53–57, March 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278323.

Dulles:1931:BRB

[Dul31] Eleanor Lansing Dulles. Book review: [untitled]. Journal of the American Statistical Association, 26(173):96–98, March 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278269.

Dulles:1934:BRB

[Dul34] Eleanor Lansing Dulles. Book review: [untitled]. Journal of the American Statistical Association, 29(186):238–240, June 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278313.

Dulles:1936:BRBa

[Dul36a] Eleanor Lansing Dulles. Book review: [untitled]. Journal of the American Statistical Association, 31(194):435–436, June 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278589.

Dulles:1936:BRBb

[Dul36b] Eleanor Lansing Dulles. Book review: [untitled]. Journal of the American Statistical Association, 31(195):633–635, September 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278406.

Dulles:1937:BRB

[Dul37] Eleanor Lansing Dulles. Book review: [untitled]. Journal of the American Statistical Association, 32(199):592–594, September 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278787.

Dunn:1932:ANG

[Dun32] Halbert L. Dunn. Adaptation of new geometric code to multiple punching in mechanical tabulation. Journal of the American Statistical Association, 27(179):279–286, September 1932. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278161.

Duncan:1935:FRF

[Dun35] Julian S. Duncan. Flexible railway freight rates. Journal of the American Statistical Association, 30(191):537–548, September 1935. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278098.

Dunn:1936:EEU

[Dun36] Halbert L. Dunn. The evaluation of the effect upon mortality statistics of the selection of the primary cause of death. Journal of the American Statistical Association, 31(193):113–123, March 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278332.

Dunn:1937:BRB

[Dun37] Halbert L. Dunn. Book review: [untitled]. Journal of the American Statistical Association, 32(199):586–587, September 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278782.

Duncan:1939:BRB

[Dun39] Acheson J. Duncan. Book review: [untitled]. Journal of the American Statistical Association, 34(207):618–619, September 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279519.

Durand:1935:ISI

[DW35] E. Dana Durand and Walter F. Willcox. The international statistical institute. Journal of the American Statistical Association, 30(191):600, September 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278108.

Eberle:1938:PEL

[Ebe38] George J. Eberle. Population estimates of local communities and economic planning. Journal of the American Statistical Association, 33(204):694–704, December 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279060.

Eblin:1938:HJ

[Ebl38a] Walter H. Eblin. Harry Jerome. Journal of the American Statistical Association, 33(204):741–742, December 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279071.

Ebling:1938:WCS

[Ebl38b] Walter H. Ebling. The Wisconsin Committee on Statistics. Journal of the American Statistical Association, 33(203):573–575, September 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279697.

Eder:1937:BRB

[Ede37] George Jackson Eder. Book review: [untitled]. Journal of the American Statistical Association, 32(197):237–240, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279892.

Edge:1931:PCP

[Edg31] P. Granville Edge. Pre-census population records of Spain. Journal of the American Statistical Association, 26(176):416–423, December 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277942.

Edie:1931:BRB

[Edi31] Lionel D. Edie. Book review: [untitled]. Journal of the American Statistical Association, 26(173):112–113, March 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278279.

Edwards:1933:SEG

[Edw33] Alba M. Edwards. A social-economic grouping of the gainful workers of the United States. Journal of the American Statistical Association, 28(184):377–387, December 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278127.

Edwards:1936:NFN

[Edw36] Alba M. Edwards. The Negro as a factor in the nation’s labor force. Journal of the American Statistical Association, 31(195):529–540, September 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278379.

Edwards:1939:BRB

[Edw39] George W. Edwards. Book review: [untitled]. Journal of the American Statistical Association, 34(205):195–196, March 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279181.

Eells:1930:MCC

[Eel30] Walter Crosby Eells. A mistaken conception of the center of population. Journal of the American Statistical Association, 25(169):33–40, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278491.

Elder:1935:BRB

[Eld35] John D. Elder. Book review: [untitled]. Journal of the American Statistical Association, 30(191):634–635, September 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278118.

Elder:1936:BRBa

[Eld36a] John D. Elder. Book review: [untitled]. Journal of the American Statistical Association, 31(196):759–760, December 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278683.

Elder:1936:BRBb

[Eld36b] John D. Elder. Book review: [untitled]. Journal of the American Statistical Association, 31(196):760, December 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278684.

Ellis:1937:BRB

[Ell37] Howard S. Ellis. Book review: [untitled]. Journal of the American Statistical Association, 32(198):415–417, June 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279592.

Elliott:1938:QSM

[Ell38] D. C. Elliott. A quarterly series of manufacturers’ inventories. Journal of the American Statistical Association, 33(202):349–352, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278910.

Elston:1936:BRB

[Els36] James S. Elston. Book review: [untitled]. Journal of the American Statistical Association, 31(196):766, December 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278688.

Edwards:1934:CCI

[EM34a] Alba M. Edwards and T. C. McCormick. Corrections: a coefficient of independent determination. Journal of the American Statistical Association, 29(187):319, September 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278076.

Edwards:1934:CSE

[EM34b] Alba M. Edwards and T. C. McCormick. Corrections: a socialeconomic grouping of the gainful workers of the United States. Journal of the American Statistical Association, 29(187):318–319, September 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278075.

Emmet:1931:BRB

[Emm31] Boris Emmet. Book review: [untitled]. Journal of the American Statistical Association, 26(173):108–110, March 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278276.

Engle:1938:BRB

[Eng38] N. H. Engle. Book review: [untitled]. Journal of the American Statistical Association, 33(202):471–472, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278935.

Epstein:1933:BRB

[Eps33] Ralph C. Epstein. Book review: [untitled]. Journal of the American Statistical Association, 28(183):364–365, September 1933. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278443.

Epstein:1934:BRBb

[Eps34a] Lillian Epstein. Book review: [untitled]. Journal of the American Statistical Association, 29(187):353, September 1934. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278093.

Epstein:1934:BRBa

[Eps34b] Ralph C. Epstein. Book review: [untitled]. Journal of the American Statistical Association, 29(187):348, September 1934. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278090.

Epstein:1934:BRBc

[Eps34c] Ralph C. Epstein. Book review: [untitled]. Journal of the American Statistical Association, 29(188):444–445, December 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278628.

Epstein:1935:GPC

[Eps35] Ralph C. Epstein. A general-purpose classification for manufacturing industries. Journal of the American Statistical Association, 30(189): 47–57, March 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278528.

Epstein:1936:BRB

[Eps36] Ralph C. Epstein. Book review: [untitled]. Journal of the American Statistical Association, 31(196):770, December 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278691.

Evans:1933:BRB

[Eva33] George Heberton Evans, Jr. Book review: [untitled]. Journal of the American Statistical Association, 28(181):117–119, March 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278024.

Evans:1936:INM

[Eva36] George Heberton Evans, Jr. The index numbers, A.M.I and A.M.II. Journal of the American Statistical Association, 31(196):726–728, December 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278671.

Ezekiel:1932:BRB

[Eze32a] Mordecai Ezekiel. Book review: [untitled]. Journal of the American Statistical Association, 27(179):338–339, September 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278174.

Ezekiel:1932:FRG

[Eze32b] Mordecai Ezekiel. Further remarks on the graphic method of correlation: I. A reply to “Some Characteristics of the Graphic Method of Correlation”. Journal of the American Statistical Association, 27(178):183–185, June 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277662.

Ezekiel:1934:EF

[Eze34] Mordecai Ezekiel. Evaluating 1933 for the farmer. Journal of the American Statistical Association, 29(186):137–146, June 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278283.

Ezekiel:1937:BRB

[Eze37] Mordecai Ezekiel. Book review: [untitled]. Journal of the American Statistical Association, 32(200):799–802, December 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279403.

Fabricant:1935:CES

[Fab35] Solomon Fabricant. Corporate earnings, 1929-31: Subgroups in nonmanufacturing industries. Journal of the American Statistical Association, 30(192):662–670, December 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277719.

Fabricant:1936:BRB

[Fab36] Solomon Fabricant. Book review: [untitled]. Journal of the American Statistical Association, 31(193):212–214, March 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278347.

Fabricant:1938:PMP

[Fab38] Solomon Fabricant. Problems in the measurement of the physical volume of output, by industries. Journal of the American Statistical Association, 33(203):564–570, September 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279695.

Falk:1937:BRB

[Fal37] I. S. Falk. Book review: [untitled]. Journal of the American Statistical Association, 32(200):821–824, December 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279417.

Falconer:1938:BRB

[Fal38] John I. Falconer. Book review: [untitled]. Journal of the American Statistical Association, 33(203):649–650, September 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279732.

Feldman:1930:BRB

[Fel30] Herman Feldman. Book review: [untitled]. Journal of the American Statistical Association, 25(172):497–498, December 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277990.

Feldstein:1936:BRB

[Fel36] Marc J. Feldstein. Book review: [untitled]. Journal of the American Statistical Association, 31(196):820, December 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278718.

Feldman:1937:BRB

[Fel37] Herman Feldman. Book review: [untitled]. Journal of the American Statistical Association, 32(199):621–622, September 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278803.

Fellner:1939:BRB

[Fel39] William Fellner. Book review: [untitled]. Journal of the American Statistical Association, 34(206):430–431, June 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278875.

Ferger:1931:BRB

[Fer31a] Wirth F. Ferger. Book review: [untitled]. Journal of the American Statistical Association, 26(174):229–230, June 1931. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277781.

Ferger:1931:NUH

[Fer31b] Wirth F. Ferger. The nature and use of the harmonic mean. Journal of the American Statistical Association, 26(173):36–40, March 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278257.

Ferger:1932:BRBa

[Fer32a] Wirth F. Ferger. Book review: [untitled]. Journal of the American Statistical Association, 27(178):217–220, June 1932. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277677.

Ferger:1932:BRBb

[Fer32b] Wirth F. Ferger. Book review: [untitled]. Journal of the American Statistical Association, 27(180):457–458, December 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278659.

Ferger:1934:BRB

[Fer34] Wirth F. Ferger. Book review: [untitled]. Journal of the American Statistical Association, 29(186):224–225, June 1934. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278303.

Ferger:1936:DCP

[Fer36] Wirth F. Ferger. Distinctive concepts of price and purchasingpower index numbers. Journal of the American Statistical Association, 31(194):258–272, June 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278557.

Ferger:1937:MPP

[Fer37] Wirth F. Ferger. The measurement of purchasing power: Reply. Journal of the American Statistical Association, 32(198):343–348, June 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279575.

Fertig:1939:BRB

[Fer39] John W. Fertig. Book review: [untitled]. Journal of the American Statistical Association, 34(205):212–213, March 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279196.

Fetter:1937:MER

[Fet37] Frank Whitson Fetter. Measurement of effects of reciprocal trade agreements: Discussion. Journal of the American Statistical Association, 32(197):61–64, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279862.

Ficek:1935:BRB

[Fic35] Karel Ficek. Book review: [untitled]. Journal of the American Statistical Association, 30(191):648–650, September 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278126.

Fisher:1930:WCP

[Fis30] Irving Fisher. Wholesale commodity price indexes — weekly, monthly, quarterly, and annual: Seventh report. Journal of the American Statistical Association, 25(171):308–315, September 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278194.

Fisher:1932:BRB

[Fis32] Irving Fisher. Book review: [untitled]. Journal of the American Statistical Association, 27(180):455–456, December 1932. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278656.

Fisher:1933:SSE

[Fis33] Irving Fisher. Statistics in the service of economics. Journal of the American Statistical Association, 28(181):1–13, March 1933. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277998.

Fisher:1936:CWPa

[Fis36a] Irving Fisher. Changes in the wholesale price index in relation to factory employment. Journal of the American Statistical Association, 31(195):496–502, September 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278374.

Fisher:1936:CWPb

[Fis36b] Irving Fisher. Changes in the wholesale price index in relation to factory employment: Rejoinder. Journal of the American Statistical Association, 31(195):505–506, September 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278376.

Fiske:1936:BRB

[Fis36c] Wyman P. Fiske. Book review: [untitled]. Journal of the American Statistical Association, 31(194):450–452, June 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278601.

Fisher:1937:BRBa

[Fis37a] Irving Fisher. Book review: [untitled]. Journal of the American Statistical Association, 32(197):221–222, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279883.

Fisher:1937:BRBb

[Fis37b] Irving Fisher. Book review: [untitled]. Journal of the American Statistical Association, 32(200):805–807, December 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279406.

Fletcher:1935:BRB

[Fle35] Ralph Carr Fletcher. Book review: [untitled]. Journal of the American Statistical Association, 30(192):782–784, December 1935. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277757.

Fletcher:1938:BRBa

[Fle38a] Ralph Carr Fletcher. Book review: [untitled]. Journal of the American Statistical Association, 33(201):289–291, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279146.

Fletcher:1938:BRBb

[Fle38b] Ralph Carr Fletcher. Book review: [untitled]. Journal of the American Statistical Association, 33(202):492–494, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278947.

Frisch:1931:SCT

[FM31] Ragnar Frisch and Bruce D. Mudgett. Statistical correlation and the theory of cluster types. Journal of the American Statistical Association, 26(176):375–392, December 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277939.

Frisch:1932:CSC

[FM32] Ragnar Frisch and Bruce D. Mudgett. A correction: Statistical correlation and the theory of cluster types. Journal of the American Statistical Association, 27(178):187, June 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277665.

Folts:1935:BRB

[Fol35] Franklin E. Folts. Book review: [untitled]. Journal of the American Statistical Association, 30(192):756–758, December 1935. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277740.

Forsyth:1932:PCS

[For32] C. H. Forsyth. Proposal of a coefficient of stability. Journal of the American Statistical Association, 27(178):173–176, June 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277659.

Forbes:1939:BRB

[For39] T. W. Forbes. Book review: [untitled]. Journal of the American Statistical Association, 34(205):230–231, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279207.

Foster:1935:BRBb

[Fos35a] LeBaron R. Foster. Book review: [untitled]. Journal of the American Statistical Association, 30(192):766–768, December 1935. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277746.

Foster:1935:BRBa

[Fos35b] William Trufant Foster. Book review: [untitled]. Journal of the American Statistical Association, 30(189):129–130, March 1935. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278545.

Foster:1938:PSC

[Fos38] William Trufant Foster. Public supervision of consumer credit. Journal of the American Statistical Association, 33(201):71–80, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279115.

Foster:1939:BRB

[Fos39a] LeBaron R. Foster. Book review: [untitled]. Journal of the American Statistical Association, 34(205):201–202, March 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279186.

Foster:1939:WMP

[Fos39b] William Trufant Foster. Warren Milton Persons. Journal of the American Statistical Association, 34(206):411–415, June 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278864.

Fox:1936:BPB

[Fox36a] Mortimer J. Fox, Jr. A broadened program of bank reports. Journal of the American Statistical Association, 31(195):549–551, September 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2278382.

Fox:1936:DII

[Fox36b] Mortimer J. Fox, Jr. Deposit insurance as an influence for stabilizing the banking structure. Journal of the American Statistical Association, 31(193):103–112, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278331.

Frain:1930:BRB

[Fra30] H. LaRue Frain. Book review: [untitled]. Journal of the American Statistical Association, 25(170):240–241, June 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277649.

Frain:1931:CAN

[Fra31] H. La Rue Frain. A comparison of actual and normal working time in standard machine-tool occupations. Journal of the American Statistical Association, 26(173):30–35, March 1931. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278256.

Freidman:1938:BRB

[Fre38] Lillian Freidman. Book review: [untitled]. Journal of the American Statistical Association, 33(201):280–281, March 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279140.

Friedman:1930:BRB

[Fri30a] Gladys R. Friedman. Book review: [untitled]. Journal of the American Statistical Association, 25(170):242–243, June 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277652.

Frisch:1930:NSC

[Fri30b] Ragnar Frisch. Necessary and sufficient conditions regarding the form of an index number which shall meet certain of Fisher’s tests. Journal of the American Statistical Association, 25(172):397–406, December 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277965.

Frisch:1934:RMS

[Fri34] Ragnar Frisch. Reply to Mr. Subramanian’s note. Journal of the American Statistical Association, 29(187):317, September 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278073.

Fritz:1936:BRBa

[Fri36a] Wilbert G. Fritz. Book review: [untitled]. Journal of the American Statistical Association, 31(194):446–447, June 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278597.

Fritz:1936:BRBb

[Fri36b] Wilbert G. Fritz. Book review: [untitled]. Journal of the American Statistical Association, 31(196):806–809, December 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278710.

Friedman:1937:URA

[Fri37a] Milton Friedman. The use of ranks to avoid the assumption of normality implicit in the analysis of variance. Journal of the American Statistical Association, 32(200):675–701, December 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279372.

Fritz:1937:BRB

[Fri37b] Wilbert G. Fritz. Book review: [untitled]. Journal of the American Statistical Association, 32(198):430–432, June 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279600.

Fritz:1938:BRB

[Fri38] Wilbert G. Fritz. Book review: [untitled]. Journal of the American Statistical Association, 33(202):475–476, June 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278937.

Friedman:1939:BRB

[Fri39a] Milton Friedman. Book review: [untitled]. Journal of the American Statistical Association, 34(205):224–225, March 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279203.

Friedman:1939:CUR

[Fri39b] Milton Friedman. A correction: The use of ranks to avoid the assumption of normality implicit in the analysis of variance. Journal of the American Statistical Association, 34(205):109, March 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279169.

Frisbee:1939:BRB

[Fri39c] Ira N. Frisbee. Book review: [untitled]. Journal of the American Statistical Association, 34(207):595–596, September 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279498.

Fritz:1939:BRBa

[Fri39d] Wilbert G. Fritz. Book review: [untitled]. Journal of the American Statistical Association, 34(206):419–420, June 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278866.

Fritz:1939:BRBb

[Fri39e] Wilbert G. Fritz. Book review: [untitled]. Journal of the American Statistical Association, 34(208):759–760, December 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279827.

Fry:1930:MUC

[Fry30] C. Luther Fry. Making use of census data. Journal of the American Statistical Association, 25(170):129–138, June 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277629.

Fry:1936:BRB

[Fry36] Thornton C. Fry. Book review: [untitled]. Journal of the American Statistical Association, 31(196):757–758, December 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278681.

Fry:1938:TS

[Fry38] Thornton C. Fry. The χ^{2}-test of significance. Journal of the American Statistical Association, 33(203):513–525, September 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279689.

Foster:1937:EVR

[FW37] R. R. Foster and David L. Wickens. Estimating the volume of residential building construction. Journal of the American Statistical Association, 32(197):97–104, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279867.

Gallup:1938:GSR

[Gal38] George Gallup. Government and the sampling referendum. Journal of the American Statistical Association, 33(201):131–142, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279120.

Garver:1938:RPP

[Gar38] Walter B. Garver. The reliability of preliminary price indexes. Journal of the American Statistical Association, 33(202):380–384, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2278914.

Gayer:1931:BRBa

[Gay31a] Arthur D. Gayer. Book review: [untitled]. Journal of the American Statistical Association, 26(173):103–104, March 1931. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278272.

Gayer:1931:BRBb

[Gay31b] Arthur D. Gayer. Book review: [untitled]. Journal of the American Statistical Association, 26(174):233–234, June 1931. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277784.

Gayer:1931:BRBc

[Gay31c] Arthur D. Gayer. Book review: [untitled]. Journal of the American Statistical Association, 26(176):498–499, December 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277962.

Gehlke:1934:CEG

[GB34] C. E. Gehlke and Katherine Biehl. Certain effects of grouping upon the size of the correlation coefficient in census tract material. Journal of the American Statistical Association, 29(185):169–170, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277827.

Gehlke:1931:BRB

[Geh31] C. E. Gehlke. Book review: [untitled]. Journal of the American Statistical Association, 26(174):221–223, June 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277778.

Gehlke:1937:BRB

[Geh37] C. E. Gehlke. Book review: [untitled]. Journal of the American Statistical Association, 32(200):814–817, December 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279413.

George:1937:BRB

[Geo37] Edwin B. George. Book review: [untitled]. Journal of the American Statistical Association, 32(200):791–792, December 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279398.

George:1939:RDC

[Geo39] Edwin B. George. Replacement demand for consumers’ durable goods. Journal of the American Statistical Association, 34(206):261– 268, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278843.

Gerald:1939:BRB

[Ger39] Norman D. Fitz Gerald. Book review: [untitled]. Journal of the American Statistical Association, 34(207):612–613, September 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279513.

Gevorkiantz:1935:BRB

[Gev35] S. R. Gevorkiantz. Book review: [untitled]. Journal of the American Statistical Association, 30(192):780–781, December 1935. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277755.

Garfield:1937:CCR

[GH37] Frank R. Garfield and William M. Hoad. Construction costs and real property values. Journal of the American Statistical Association, 32(200):643–653, December 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279369.

Gideonse:1933:BRB

[Gid33] Harry D. Gideonse. Book review: [untitled]. Journal of the American Statistical Association, 28(181):116–117, March 1933. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278023.

Gideonse:1939:BRB

[Gid39] Max Gideonse. Book review: [untitled]. Journal of the American Statistical Association, 34(205):233–235, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279210.

Gilboy:1933:BRB

[Gil33a] Elizabeth Waterman Gilboy. Book review: [untitled]. Journal of the American Statistical Association, 28(184):459–462, December 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278142.

Gill:1933:CS

[Gil33b] Corrington Gill. Construction statistics. Journal of the American Statistical Association, 28(181):31–54, March 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278000.

Gillman:1934:BRB

[Gil34] Joseph M. Gillman. Book review: [untitled]. Journal of the American Statistical Association, 29(186):231–233, June 1934. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278309.

Gillman:1935:BRB

[Gil35] Joseph M. Gillman. Book review: [untitled]. Journal of the American Statistical Association, 30(189):139–140, March 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278553.

Gillman:1936:BRBa

[Gil36a] Joseph M. Gillman. Book review: [untitled]. Journal of the American Statistical Association, 31(195):651–652, September 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278416.

Gillman:1936:BRBb

[Gil36b] Joseph M. Gillman. Book review: [untitled]. Journal of the American Statistical Association, 31(196):789–791, December 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278702.

Gillman:1937:BRB

[Gil37] Joseph M. Gillman. Book review: [untitled]. Journal of the American Statistical Association, 32(200):817–819, December 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279414.

Gini:1930:PCF

[Gin30] Corrado Gini. Present condition and future progress of statistics. Journal of the American Statistical Association, 25(171):295–304, September 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278192.

Girshick:1936:PC

[Gir36] M. A. Girshick. Principal components. Journal of the American Statistical Association, 31(195):519–528, September 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278378.

Girshick:1939:SBS

[Gir39] M. A. Girshick. A scientific basis for standard sizes for Children’s garments and patterns. Journal of the American Statistical Association, 34(206):362–364, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278858.

Givens:1931:BRB

[Giv31a] Meredith B. Givens. Book review: [untitled]. Journal of the American Statistical Association, 26(175):369–370, September 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277928.

Givens:1931:SMS

[Giv31b] Meredith B. Givens. Statistical measures of social aspects of unemployment. Journal of the American Statistical Association, 26 (175):303–318, September 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277908.

Givens:1934:CGSa

[Giv34a] Meredith B. Givens. Committee on government statistics and information services. Journal of the American Statistical Association, 29(186):194–196, June 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278294.

Givens:1934:CGSb

[Giv34b] Meredith B. Givens. Committee on government statistics and information services. Journal of the American Statistical Association, 29(187):319–320, September 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278077.

Givens:1934:EAS

[Giv34c] Meredith B. Givens. An experiment in advisory service: The committee on government statistics and information services. Journal of the American Statistical Association, 29(188):394–404, December 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278610.

Givens:1935:SIS

[Giv35] Meredith B. Givens. Statistical implications of the social security program. Journal of the American Statistical Association, 30 (192):651–661, December 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277718.

Glasser:1930:BRB

[Gla30] Harold Glasser. Book review: [untitled]. Journal of the American Statistical Association, 25(172):498–499, December 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277991.

Glover:1931:BRB

[Glo31] James W. Glover. Book review: [untitled]. Journal of the American Statistical Association, 26(175):353–356, September 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277918.

Gok:1936:BRB

[Gok36] Barbara Gok. Book review: [untitled]. Journal of the American Statistical Association, 31(194):444–446, June 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278596.

Gold:1934:NEU

[Gol34] Norman Leon Gold. Net emigration from the United States to Canada: 1909–1931. Journal of the American Statistical Association, 29(187):282–287, September 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278064.

Goldenweiser:1936:SLF

[Gol36] E. A. Goldenweiser. Significance of the lending function of the Federal Reserve Banks. Journal of the American Statistical Association, 31(193):95–102, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278330.

Goldenweiser:1938:FDB

[Gol38] E. A. Goldenweiser. The function of deposit banking. Journal of the American Statistical Association, 33(202):303–310, June 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278904.

Goldenweiser:1939:JH

[Gol39] E. A. Goldenweiser. Joseph A. Hill. Journal of the American Statistical Association, 34(206):407–411, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278863.

Goodrich:1934:BRB

[Goo34] Carter Goodrich. Book review: [untitled]. Journal of the American Statistical Association, 29(185):116–117, March 1934. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278482.

Gosnell:1937:HCB

[Gos37] F. A. Gosnell. Highlights of the 1935 census of business. Journal of the American Statistical Association, 32(198):263–270, June 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279568.

Gosnell:1939:CBG

[Gos39] Fred A. Gosnell. The 1939 census of business gets under way. Journal of the American Statistical Association, 34(208):667–673, December 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279811.

Gourrich:1936:SRS

[Gou36] Paul P. Gourrich. Statistics relating to security markets and corporations. Journal of the American Statistical Association, 31(193):88–94, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278329.

Graham:1938:BRB

[Gra38] Frank D. Graham. Book review: [untitled]. Journal of the American Statistical Association, 33(204):771–773, December 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279089.

Greenwood:1932:OL

[Gre32] M. Greenwood. An open letter. Journal of the American Statistical Association, 27(179):310, September 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278166.

Green:1934:RPI

[Gre34] Howard Whipple Green. The real property inventory. Journal of the American Statistical Association, 29(185):105–110, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277813.

Green:1936:BRB

[Gre36a] Howard Whipple Green. Book review: [untitled]. Journal of the American Statistical Association, 31(193):232–234, March 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278359.

Greene:1936:BRB

[Gre36b] James H. Greene. Book review: [untitled]. Journal of the American Statistical Association, 31(195):643–646, September 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278412.

Green:1937:BRB

[Gre37] Howard Whipple Green. Book review: [untitled]. Journal of the American Statistical Association, 32(199):620–621, September 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278802.

Greville:1938:BRB

[Gre38] T. N. E. Greville. Book review: [untitled]. Journal of the American Statistical Association, 33(204):752–753, December 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279077.

Greville:1939:BRB

[Gre39] Thomas N. E. Greville. Book review: [untitled]. Journal of the American Statistical Association, 34(206):447, June 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278886.

Griffin:1930:NCS

[Gri30] Harold D. Griffin. Nomograms for correcting simple and multiple correlation coefficients. Journal of the American Statistical Association, 25(171):316–319, September 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278195.

Griffin:1932:CPC

[Gri32] Harold D. Griffin. On the coefficient of part correlation. Journal of the American Statistical Association, 27(179):298–301, September 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278163.

Grinstead:1935:BRB

[Gri35] L. H. Grinstead. Book review: [untitled]. Journal of the American Statistical Association, 30(189):140–141, March 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278554.

Griffin:1936:BRB

[Gri36] C. E. Griffin. Book review: [untitled]. Journal of the American Statistical Association, 31(193):214–217, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278348.

Gosnell:1936:FCA

[GS36] Harold F. Gosnell and Margaret J. Schmidt. Factorial and correlational analysis of the 1934 vote in Chicago. Journal of the American Statistical Association, 31(195):507–518, September 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278377.

H:1930:PWCa

[H.30a] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 25(169):90–93, March 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278505.

H:1930:PWCb

[H.30b] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 25(171):344–347, September 1930.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278202.

H:1930:PWCc

[H.30c] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 25(172):468–471, December 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277976.

H:1931:PWCa

[H.31a] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 26(173):70–73, March 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278264.

H:1931:PWCb

[H.31b] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 26(174):201–204, June 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277772.

H:1931:PWCc

[H.31c] J. A. H. Progress of work in the Census Bureau: Stage of progress on decennial census. Journal of the American Statistical Association, 26(176):466–467, December 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277950.

H:1932:PWCa

[H.32a] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 27(177):90–92, March 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277887.

H:1932:PWCb

[H.32b] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 27(178):201–204, June 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277672.

H:1932:PWCc

[H.32c] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 27(180):436–438, December 1932.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278645.

H:1933:PWCa

[H.33a] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 28(182):221–223, June 1933. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278041.

H:1933:PWCb

[H.33b] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 28(183):345–346, September 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278434.

H:1933:PWCc

[H.33c] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 28(184):444–448, December 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278137.

H:1934:MN

[H.34a] J. A. H. Miscellaneous notes. Journal of the American Statistical Association, 29(186):210–218, June 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278298.

H:1934:PWCb

[H.34b] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 29(186):207–210, June 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278297.

H:1934:PWCc

[H.34c] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 29(187):325–327, September 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278080.

H:1934:PWCa

[H.34d] J. A. H. Progress of work in the Census Bureau: a census of American business. Journal of the American Statistical Association, 29(185):86– 89, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278468.

H:1934:PWCd

[H.34e] J. A. H. Progress of work in the Census Bureau: Progress census of agriculture. Journal of the American Statistical Association, 29(188):425–428, December 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278621.

H:1935:PWCa

[H.35a] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 30(189):103–106, March 1935. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278536.

H:1935:PWCc

[H.35b] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 30(191):594–597, September 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278106.

H:1935:PWCd

[H.35c] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 30(192):711–714, December 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277729.

H:1935:PWCb

[H.35d] J. A. H. Progress of work in the Census Bureau: Census of agriculture. Journal of the American Statistical Association, 30(190):451– 453, June 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277697.

H:1936:PWCa

[H.36a] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 31(193):141–143, March 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278341.

H:1936:PWCb DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278568.

H:1936:PWCc

[H.36c] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 31(195):558–560, September 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278386.

H:1937:PWCa

[H.37a] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 32(197):135–137, March 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279874.

H:1937:PWCb

[H.37b] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 32(198):377–379, June 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279587.

H:1937:PWCc

[H.37c] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 32(199):543–545, September 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278775.

H:1937:PWCd

[H.37d] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 32(200):719–720, December 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279377.

H:1938:PWCa

[H.38a] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 33(201):230–232, March 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279129.

H:1938:PWCb DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278921.

H:1938:PWCc

[H.38c] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 33(204):722–724, December 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279068.

Haas:1936:SSS

[Haa36] George C. Haas. Special studies on statistics of income. Journal of the American Statistical Association, 31(193):132–134, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278339.

Haber:1932:FED

[Hab32] William Haber. Fluctuations in employment in Detroit factories, 1921–1931. Journal of the American Statistical Association, 27(178):141–152, June 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277655.

Hadley:1938:BRB

[Had38] Clausin D. Hadley. Book review: [untitled]. Journal of the American Statistical Association, 33(204):757–759, December 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279081.

Hager:1937:BRB

[Hag37] Beatrice Hager. Book review: [untitled]. Journal of the American Statistical Association, 32(199):622–624, September 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278804.

Hagood:1939:BRB

[Hag39] Margaret Jarman Hagood. Book review: [untitled]. Journal of the American Statistical Association, 34(205):213–215, March 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279197.

Haig:1932:BRB

[Hai32] Robert Murray Haig. Book review: [untitled]. Journal of the American Statistical Association, 27(177):103–106, March 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277892.

Haig:1933:BRB

[Hai33] Robert Murray Haig. Book review: [untitled]. Journal of the American Statistical Association, 28(184):467–470, December 1933. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278146.

Hall:1933:BRB

[Hal33] Ray Ovid Hall. Book review: [untitled]. Journal of the American Statistical Association, 28(181):111–112, March 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278018.

Halligan:1936:PST

[Hal36] C. W. Halligan. The practical side of trade association statistics. Journal of the American Statistical Association, 31(195):491–495, September 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278373.

Hall:1937:BRB

[Hal37] Ray Ovid Hall. Book review: [untitled]. Journal of the American Statistical Association, 32(200):757–759, December 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279383.

Hall:1938:BRB

[Hal38] Ray Ovid Hall. Book review: [untitled]. Journal of the American Statistical Association, 33(204):769–771, December 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279088.

Ham:1933:BRB

[Ham33a] William T. Ham. Book review: [untitled]. Journal of the American Statistical Association, 28(184):475–477, December 1933. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278150.

Hamilton:1933:ADN

[Ham33b] C. Horace Hamilton. Algebraic derivation of the normal equations involved in multiple and partial correlation. Journal of the American Statistical Association, 28(182):204–205, June 1933. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278035.

Hamilton:1935:BRB

[Ham35] C. Horace Hamilton. Book review: [untitled]. Journal of the American Statistical Association, 30(192):765–766, December 1935. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277745.

Hamilton:1936:BRB

[Ham36] Earl J. Hamilton. Book review: [untitled]. Journal of the American Statistical Association, 31(196):798–800, December 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278705.

Hansen:1930:BRB

[Han30] Alvin H. Hansen. Book review: [untitled]. Journal of the American Statistical Association, 25(170):233–234, June 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277644.

Hansen:1933:BRB

[Han33] Alvin H. Hansen. Book review: [untitled]. Journal of the American Statistical Association, 28(182):253–254, June 1933. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278055.

Haney:1934:PBR

[Han34a] Lewis H. Haney. The processes of business revival in the light of 1933 conditions. Journal of the American Statistical Association, 29(185): 75–80, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277807.

Hanna:1934:ASB

[Han34b] Hugh S. Hanna. Adequacy of the sample in budgetary studies. Journal of the American Statistical Association, 29(185):131–134, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277818.

Hansen:1935:BRB

[Han35] Alvin H. Hansen. Book review: [untitled]. Journal of the American Statistical Association, 30(189):120–122, March 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278541.

Hansen:1937:BRB

[Han37] Alvin H. Hansen. Book review: [untitled]. Journal of the American Statistical Association, 32(199):601–602, September 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278791.

Hanna:1939:SNN

[Han39] Frank A. Hanna. Statistical news and notes. Journal of the American Statistical Association, 34(207):549–576, September 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279489.

Harriman:1930:BRB

[Har30] John W. Harriman. Book review: [untitled]. Journal of the American Statistical Association, 25(170):235–236, June 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277645.

Harris:1932:BRB

[Har32a] S. E. Harris. Book review: [untitled]. Journal of the American Statistical Association, 27(179):335–338, September 1932. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278173.

Hartkemeier:1932:BRB

[Har32b] Harry Pelle Hartkemeier. Book review: [untitled]. Journal of the American Statistical Association, 27(180):449–451, December 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278652.

Harris:1934:BRB

[Har34a] Abram L. Harris. Book review: [untitled]. Journal of the American Statistical Association, 29(187):351–353, September 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278092.

Harwood:1934:OLC

[Har34b] E. C. Harwood. An open letter on “cause and control of the business cycle”. Journal of the American Statistical Association, 29(185):83– 85, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278465.

Harbeson:1936:INP

[Har36a] Robert W. Harbeson. Index numbers and public utility valuation. Journal of the American Statistical Association, 31(194):245–257, June 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278556.

Hartkemeier:1936:BRB

[Har36b] Harry Pelle Hartkemeier. Book review: [untitled]. Journal of the American Statistical Association, 31(194):454–456, June 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278604.

Hardy:1937:BRBa

[Har37a] Charles O. Hardy. Book review: [untitled]. Journal of the American Statistical Association, 32(197):225–227, March 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279885.

Hardy:1937:BRBb

[Har37b] Charles O. Hardy. Book review: [untitled]. Journal of the American Statistical Association, 32(199):611–613, September 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278797.

Hartsough:1937:BRB

[Har37c] Mildred Hartsough. Book review: [untitled]. Journal of the American Statistical Association, 32(200):773, December 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279388.

Harris:1939:BRB

[Har39a] S. E. Harris. Book review: [untitled]. Journal of the American Statistical Association, 34(207):615–616, September 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279517.

Hartkemeier:1939:BRB

[Har39b] Harry Pelle Hartkemeier. Book review: [untitled]. Journal of the American Statistical Association, 34(206):424–425, June 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278869.

Hartkemeier:1939:SPT

[Har39c] Harry Pelle Hartkemeier. Some problems in teaching elementary statistics: Discussion. Journal of the American Statistical Association, 34(206):306–308, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278849.

Hartshorne:1939:BRB

[Har39d] E. Y. Hartshorne. Book review: [untitled]. Journal of the American Statistical Association, 34(208):781–782, December 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279846.

Hauser:1937:BRB

[Hau37] Ernest O. Hauser. Book review: [untitled]. Journal of the American Statistical Association, 32(199):617–619, September 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278800.

Hayes:1936:BRB

[Hay36a] H. Gordon Hayes. Book review: [untitled]. Journal of the American Statistical Association, 31(195):642–643, September 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278411.

Hayford:1936:SUS

[Hay36b] F. Leslie Hayford. Some uses of statistics in executive control. Journal of the American Statistical Association, 31(193):31–36, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278319.

Holsen:1934:ICA

[HCW^{+}34] James N. Holsen, Alonzo B. Cox, Ralph J. Watkins, Waldo E. Grimes, Paul T. Cherington, Casimir A. Sienkiewicz, and F. L. Carmichael. The industrial and commercial area vs. the political area as the unit for the collection of business statistics. Journal of the American Statistical Association, 29(185):14–23, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277796.

Hanson:1930:WDL

[HD30] Alice C. Hanson and Paul H. Douglas. The wages of domestic labor in Chicago, 1890–1929. Journal of the American Statistical Association, 25(169):47–50, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278493.

Hedlund:1936:BRB

[Hed36] G. W. Hedlund. Book review: [untitled]. Journal of the American Statistical Association, 31(195):631–633, September 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278405.

Heiberg:1931:ELL

[Hei31] Povl Heiberg. Has Engel’s law its limitations? Journal of the American Statistical Association, 26(174):175–177, June 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277766.

Heidingsfield:1939:BRB

[Hei39] Myron S. Heidingsfield. Book review: [untitled]. Journal of the American Statistical Association, 34(208):780–781, December 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279845.

Henderson:1932:BRB

[Hen32] Robert Henderson. Book review: [untitled]. Journal of the American Statistical Association, 27(177):116, March 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277899.

Henig:1939:BRB

[Hen39] Harry Henig. Book review: [untitled]. Journal of the American Statistical Association, 34(205):196–197, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279182.

Herwitz:1930:BRB

[Her30] H. K. Herwitz. Book review: [untitled]. Journal of the American Statistical Association, 25(169):128, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278522.

Hewes:1930:BRB

[Hew30] Amy Hewes. Book review: [untitled]. Journal of the American Statistical Association, 25(169):114–117, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278512.

Harris:1930:IAC

[HG30] J. Arthur Harris and Borghild Gunstad. Illustration of the application of a coefficient measuring the correlation between a variable and the deviation of an associated but not dependent variable from its probable value. Journal of the American Statistical Association, 25(171):305–307, September 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278193.

Hotelling:1931:CBR

[HH31] Harold Hotelling and Floy Hotelling. Causes of birth rate fluctuations. Journal of the American Statistical Association, 26(174):135– 149, June 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277762.

Hickernell:1939:ICM

[Hic39] Warren F. Hickernell. An index of 10 commodities, monthly, 1815–60 in the New York City wholesale markets. Journal of the American Statistical Association, 34(205):71–78, March 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279164.

Hill:1931:ESM

[Hil31] A. C. Hill, Jr. Employment statistics as measures of unemployment. Journal of the American Statistical Association, 26(174):184–187, June 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277768.

Hillhouse:1938:BRB

[Hil38] A. M. Hillhouse. Book review: [untitled]. Journal of the American Statistical Association, 33(201):299–300, March 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279152.

Hinrichs:1932:BRB

[Hin32] A. F. Hinrichs. Book review: [untitled]. Journal of the American Statistical Association, 27(178):229–230, June 1932. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277682.

Hinrichs:1938:SBP

[Hin38] A. F. Hinrichs. Statistical bias in primary data and public policy. Journal of the American Statistical Association, 33(201):143–152, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279121.

Hincks:1939:BRB

[Hin39] Edward B. Hincks. Book review: [untitled]. Journal of the American Statistical Association, 34(208):758–759, December 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279826.

Hiscock:1932:BRB

[His32] Ira V. Hiscock. Book review: [untitled]. Journal of the American Statistical Association, 27(179):351, September 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278184.

Hobson:1939:BRB

[Hob39] Asher Hobson. Book review: [untitled]. Journal of the American Statistical Association, 34(207):597–599, September 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279500.

Hoffer:1934:BRB

[Hof34a] Irwin S. Hoffer. Book review: [untitled]. Journal of the American Statistical Association, 29(188):451–453, December 1934. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278633.

Hoffman:1934:BRB

[Hof34b] Frederick L. Hoffman. Book review: [untitled]. Journal of the American Statistical Association, 29(186):222–224, June 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278302.

Hogg:1930:SIE

[Hog30] Margaret H. Hogg. Sources of incomparability and error in employment-unemployment surveys. Journal of the American Statistical Association, 25(171):284–294, September 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278191.

Hogg:1931:BRBa

[Hog31a] Margaret H. Hogg. Book review: [untitled]. Journal of the American Statistical Association, 26(174):230–231, June 1931. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277782.

Hogg:1931:BRBb

[Hog31b] Margaret H. Hogg. Book review: [untitled]. Journal of the American Statistical Association, 26(176):494, December 1931. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277958.

Hogg:1931:DCL

[Hog31c] Margaret H. Hogg. A distortion in the cost of living index. Journal of the American Statistical Association, 26(173):52–57, March 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278259.

Hogg:1934:BRB

[Hog34a] Margaret H. Hogg. Book review: [untitled]. Journal of the American Statistical Association, 29(185):110–112, March 1934. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278478.

Hogg:1934:RWE

[Hog34b] Margaret H. Hogg. Revising the wage earners’ cost-of-living index. Journal of the American Statistical Association, 29(185):120–124, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277816.

Hollander:1932:BRB

[Hol32] Jacob H. Hollander. Book review: [untitled]. Journal of the American Statistical Association, 27(177):109, March 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277894.

Holzinger:1936:BRB

[Hol36] Karl J. Holzinger. Book review: [untitled]. Journal of the American Statistical Association, 31(195):620–622, September 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278399.

Holzinger:1937:NPK

[Hol37] Karl J. Holzinger. Note on Professor Kelley’s method. Journal of the American Statistical Association, 32(198):360–362, June 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279578.

Homan:1933:ACD

[Hom33] Jeanette Homan. Adjusting for the changing data of Easter in economic series. Journal of the American Statistical Association, 28 (183):328–332, September 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278428.

Homan:1934:BRB

[Hom34] Paul T. Homan. Book review: [untitled]. Journal of the American Statistical Association, 29(188):439–440, December 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278625.

Hoover:1933:BRB

[Hoo33] Calvin B. Hoover. Book review: [untitled]. Journal of the American Statistical Association, 28(181):109–110, March 1933. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278016.

Hoover:1937:BRB

[Hoo37] Edgar M. Hoover, Jr. Book review: [untitled]. Journal of the American Statistical Association, 32(198):429–430, June 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279599.

Hoover:1938:BRBa

[Hoo38a] Calvin B. Hoover. Book review: [untitled]. Journal of the American Statistical Association, 33(201):282–284, March 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279142.

Hoover:1938:BRBb

[Hoo38b] Edgar M. Hoover, Jr. Book review: [untitled]. Journal of the American Statistical Association, 33(204):749–751, December 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279075.

Hoover:1939:BRB

[Hoo39] Edgar M. Hoover, Jr. Book review: [untitled]. Journal of the American Statistical Association, 34(208):767–768, December 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279833.

Horst:1931:OCS

[Hor31] Paul Horst. Obtaining comparable scores from distributions of dissimilar shape. Journal of the American Statistical Association, 26 (176):455–460, December 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277947.

Horst:1932:GME

[Hor32] Paul Horst. A general method for evaluating multiple regression constants. Journal of the American Statistical Association, 27 (179):270–278, September 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278160.

Horst:1935:BRB

[Hor35a] Paul Horst. Book review: [untitled]. Journal of the American Statistical Association, 30(192):781–782, December 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277756.

Horst:1935:CGM

[Hor35b] Paul Horst. A correction: a general method for evaluating multiple regression constants. Journal of the American Statistical Association, 30(189):102, March 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278534.

Hotelling:1930:BRB

[Hot30a] Harold Hotelling. Book review: [untitled]. Journal of the American Statistical Association, 25(171):381–382, September 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278215.

Hotelling:1930:BSS

[Hot30b] Harold Hotelling. British statistics and statisticians today. Journal of the American Statistical Association, 25(170):186–190, June 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277631.

Hotelling:1932:BRBa

[Hot32a] Harold Hotelling. Book review: [untitled]. Journal of the American Statistical Association, 27(178):215–217, June 1932. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277676.

Hotelling:1932:BRBb

[Hot32b] Harold Hotelling. Book review: [untitled]. Journal of the American Statistical Association, 27(180):451–452, December 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278653.

Hotelling:1933:BRBb

[Hot33a] Harold Hotelling. Book review: The Triumph of Mediocrity in Business, by Horace Secrist. Journal of the American Statistical Association, 28(184):463–465, December 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278144.

Hotelling:1933:BRBa

[Hot33b] Harold Hotelling. Book review: [untitled]. Journal of the American Statistical Association, 28(183):374–375, September 1933. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278451.

Hotelling:1935:BRBa

[Hot35a] Harold Hotelling. Book review: [untitled]. Journal of the American Statistical Association, 30(189):117–118, March 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278538.

Hotelling:1935:BRBb

[Hot35b] Harold Hotelling. Book review: [untitled]. Journal of the American Statistical Association, 30(189):118, March 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278539.

Hotelling:1935:BRBc

[Hot35c] Harold Hotelling. Book review: [untitled]. Journal of the American Statistical Association, 30(190):480–483, June 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277704.

Hotelling:1935:BRBd

[Hot35d] Harold Hotelling. Book review: [untitled]. Journal of the American Statistical Association, 30(192):771–772, December 1935. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277749.

Hotelling:1937:BRBa

[Hot37a] Harold Hotelling. Book review: [untitled]. Journal of the American Statistical Association, 32(197):218–219, March 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279881.

Hotelling:1937:BRBb

[Hot37b] Harold Hotelling. Book review: [untitled]. Journal of the American Statistical Association, 32(199):580–582, September 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278780.

Hotelling:1938:BRB

[Hot38] Harold Hotelling. Book review: [untitled]. Journal of the American Statistical Association, 33(204):744–747, December 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279073.

Hotelling:1939:BRB

[Hot39] Harold Hotelling. Book review: [untitled]. Journal of the American Statistical Association, 34(206):423–424, June 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278868.

Houghton:1939:BRB

[Hou39] Harrison F. Houghton. Book review: [untitled]. Journal of the American Statistical Association, 34(206):428–429, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278873.

Howell:1936:BRBa

[How36a] J. W. Howell. Book review: [untitled]. Journal of the American Statistical Association, 31(193):227–228, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278355.

Howell:1936:BRBb

[How36b] Paul L. Howell. Book review: [untitled]. Journal of the American Statistical Association, 31(196):809–811, December 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278712.

Howell:1937:BRB

[How37] James W. Howell. Book review: [untitled]. Journal of the American Statistical Association, 32(197):236–237, March 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279891.

Hoyt:1939:BRB

[Hoy39] Homer Hoyt. Book review: [untitled]. Journal of the American Statistical Association, 34(205):215–216, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279198.

Holmes:1931:SNI

[HP31] S. J. Holmes and S. L. Parker. The stabilized natural increase of the Negro. Journal of the American Statistical Association, 26(174):159– 171, June 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277764.

Hubbard:1937:BRB

[Hub37] Joseph B. Hubbard. Book review: [untitled]. Journal of the American Statistical Association, 32(200):824–827, December 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279418.

Hudson:1939:TPL

[Hud39] Philip G. Hudson. The technical problems and limitations to the construction of indexes of physical production. Journal of the American Statistical Association, 34(206):239–251, June 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278841.

Humphrey:1937:SAC

[Hum37] Don D. Humphrey. Some adjustments in census data on unemployment. Journal of the American Statistical Association, 32 (200):702–712, December 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279373.

Humphrey:1939:PRS

[Hum39] Don D. Humphrey. Price reduction as a stimulus to sales of durable consumers’ goods. Journal of the American Statistical Association, 34(206):252–260, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278842.

Huntington:1932:IEF

[Hun32] Edward V. Huntington. An improved equal-frequency map of the normal correlation surface, using circles instead of ellipses. Journal of the American Statistical Association, 27(179):251–255, September 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2278158.

Hurlin:1934:BRB

[Hur34a] Ralph G. Hurlin. Book review: [untitled]. Journal of the American Statistical Association, 29(188):450–451, December 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278632.

Hurlin:1934:SFC

[Hur34b] Ralph G. Hurlin. State and Federal cooperation in the collection of social welfare statistics. Journal of the American Statistical Association, 29(185):94–98, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277810.

Hartkemeier:1935:NRC

[HW35] Harry Pelle Hartkemeier and R. Clyde White. Note on R. Clyde White’s “Social Statistics”. Journal of the American Statistical Association, 30(191):586–589, September 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278104.

Immer:1937:CBM

[Imm37] F. R. Immer. Correlation between means and standard deviations in field experiments. Journal of the American Statistical Association, 32(199):525–531, September 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278772.

Inghram:1939:BRB

[Ing39] Howell A. Inghram. Book review: [untitled]. Journal of the American Statistical Association, 34(205):210, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279193.

Inman:1939:BRB

[Inm39] M. K. Inman. Book review: [untitled]. Journal of the American Statistical Association, 34(206):450–451, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278889.

Isaacs:1934:BRB

[Isa34] Asher Isaacs. Book review: [untitled]. Journal of the American Statistical Association, 29(188):442–444, December 1934. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278627.

Isaacs:1936:BRB

[Isa36] Nathan Isaacs. Book review: [untitled]. Journal of the American Statistical Association, 31(196):774–775, December 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278694.

Isaacs:1938:BRB

[Isa38] Asher Isaacs. Book review: [untitled]. Journal of the American Statistical Association, 33(204):773–774, December 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279090.

Ise:1936:BRB

[Ise36] John Ise. Book review: [untitled]. Journal of the American Statistical Association, 31(196):812–813, December 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278714.

Jaffe:1939:RDN

[Jaf39] A. J. Jaffe. Religious differentials in the net reproduction rate. Journal of the American Statistical Association, 34(206):335–342, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2278855.

James:1935:BRB

[Jam35] Clifford L. James. Book review: [untitled]. Journal of the American Statistical Association, 30(189):137–138, March 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278552.

Janssen:1938:CMS

[Jan38] Theodore A. Janssen. Comparability of mortality statistics. Journal of the American Statistical Association, 33(202):399–403, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278917.

Jarchow:1937:EPS

[Jar37] Christian E. Jarchow. The Employer’s problem in social security reporting. Journal of the American Statistical Association, 32(197): 28–31, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279856.

Jenks:1934:BRB

[Jen34] Leland H. Jenks. Book review: [untitled]. Journal of the American Statistical Association, 29(185):112–113, March 1934. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278479.

Jerome:1931:BRB

[Jer31] Harry Jerome. Book review: [untitled]. Journal of the American Statistical Association, 26(174):241, June 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277792.

Jerome:1932:BRB

[Jer32] Harry Jerome. Book review: [untitled]. Journal of the American Statistical Association, 27(177):120, March 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277902.

Jerome:1936:BRB

[Jer36] Harry Jerome. Book review: [untitled]. Journal of the American Statistical Association, 31(193):217–218, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278349.

Johnson:1935:TLG

[Joh35] Norris O. Johnson. A trend line for growth series. Journal of the American Statistical Association, 30(192):717, December 1935. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277731.

Johnson:1936:TLG

[Joh36] Norris O. Johnson. A trend line for growth series, further remarks. Journal of the American Statistical Association, 31(196):731, December 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278673.

Johnson:1938:BRB

[Joh38a] E. A. J. Johnson. Book review: [untitled]. Journal of the American Statistical Association, 33(203):608–609, September 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279711.

Johnson:1938:NIP

[Joh38b] Norris O. Johnson. New indexes of production and trade. Journal of the American Statistical Association, 33(202):341–348, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278909.

Johnson:1939:BRB

[Joh39] Ernest A. Johnson. Book review: [untitled]. Journal of the American Statistical Association, 34(206):448, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278887.

Jones:1937:BRB

[Jon37] Joseph M. Jones. Book review: [untitled]. Journal of the American Statistical Association, 32(198):419–421, June 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279594.

Jones:1938:SPB

[Jon38] Homer Jones. Some problems of bank supervision. Journal of the American Statistical Association, 33(202):334–340, June 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278908.

Joy:1930:BRB

[Joy30a] Aryness Joy. Book review: [untitled]. Journal of the American Statistical Association, 25(172):499–500, December 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277992.

Joy:1930:IPM

[Joy30b] Aryness Joy. Index of production of manufactures derived from census data, 1927. Journal of the American Statistical Association, 25(172):453–459, December 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277971.

Joy:1933:BRB

[Joy33] Aryness Joy. Book review: [untitled]. Journal of the American Statistical Association, 28(181):115–116, March 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278021.

Joy:1934:IPM

[Joy34a] Aryness Joy. Index of production of manufactures derived from census data, 1929, 1931. Journal of the American Statistical Association, 29(187):305–307, September 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278068.

Joy:1934:RPE

[Joy34b] Aryness Joy. Recent progress in employment statistics. Journal of the American Statistical Association, 29(188):355–371, December 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278606.

Jasny-Philippi:1936:BRB

[JP36] Marie Jasny-Philippi. Book review: [untitled]. Journal of the American Statistical Association, 31(194):448–449, June 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278599.

Joy:1939:CUU

[JW39] Aryness Joy and Loring Wood. Concepts used in unemployment surveys: Discussion. Journal of the American Statistical Association, 34(205):59–61, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279162.

Kahler:1939:BRB

[Kah39] Alfred Kahler. Book review: [untitled]. Journal of the American Statistical Association, 34(208):760–762, December 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279828.

Kaplan:1938:EPU

[Kap38] A. D. H. Kaplan. Expenditure patterns of urban families. Journal of the American Statistical Association, 33(201):81–100, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279116.

Kellogg:1933:BRB

[Kel33] Lester S. Kellogg. Book review: [untitled]. Journal of the American Statistical Association, 28(184):481–483, December 1933. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278153.

Kellogg:1934:BRB

[Kel34] Lester S. Kellogg. Book review: [untitled]. Journal of the American Statistical Association, 29(187):345–346, September 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278088.

Kellogg:1936:BRB

[Kel36] Lester S. Kellogg. Book review: [untitled]. Journal of the American Statistical Association, 31(194):418–420, June 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278577.

Kelley:1938:BRB

[Kel38] Truman L. Kelley. Book review: [untitled]. Journal of the American Statistical Association, 33(204):753–755, December 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279078.

Kellogg:1939:SPT

[Kel39] Lester S. Kellogg. Some problems in teaching elementary statistics. Journal of the American Statistical Association, 34(206):299– 306, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278848.

Kemp:1934:SMS

[Kem34] W. B. Kemp. Some methods for statistical analysis. Journal of the American Statistical Association, 29(186):147–158, June 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278284.

Kennedy:1936:ISM

[Ken36] S. J. Kennedy. Industry statistics in marketing management. Journal of the American Statistical Association, 31(196):702–707, December 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2278667.

Kester:1936:BRB

[Kes36] Roy B. Kester. Book review: [untitled]. Journal of the American Statistical Association, 31(196):772–774, December 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278693.

Killough:1938:BRB

[Kil38] Hugh B. Killough. Book review: [untitled]. Journal of the American Statistical Association, 33(204):776–777, December 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279092.

King:1930:MAS

[Kin30a] Robert B. King. A method of appraising short-term forecasts. Journal of the American Statistical Association, 25(171):333–334, September 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2278198.

King:1930:BM

[Kin30b] Willford I. King. Bank mergers. Journal of the American Statistical Association, 25(169):86–89, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278504.

King:1930:TBC

[Kin30c] Willford I. King. Is there a business cycle? Journal of the American Statistical Association, 25(170):207–209, June 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277636.

King:1930:MBD

[Kin30d] Willford I. King. Meeting on better data on merchandise distribution. Journal of the American Statistical Association, 25(171):336–339, September 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278200.

King:1930:MCV

[Kin30e] Willford I. King. Meeting on cyclical variations in employment. Journal of the American Statistical Association, 25(171):339–344, September 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278201.

King:1930:RAP

[Kin30f] Willford I. King. Railroads and American prosperity. Journal of the American Statistical Association, 25(170):210–213, June 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277637.

King:1930:WP

[Kin30g] Willford I. King. Whither prices? Journal of the American Statistical Association, 25(172):465–468, December 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277974.

King:1931:BRB

[Kin31a] Willford I. King. Book review: [untitled]. Journal of the American Statistical Association, 26(176):499–500, December 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277963.

King:1931:MAC

[Kin31b] Willford I. King. Methods of analyzing consumer attitudes. Journal of the American Statistical Association, 26(176):463–465, December 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277949.

King:1931:MDS

[Kin31c] Willford I. King. Modern developments in statistical theory. Journal of the American Statistical Association, 26(175):331–333, September 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277911.

King:1931:PCP

[Kin31d] Willford I. King. Prospective changes in the population of the United States. Journal of the American Statistical Association, 26(174):192– 197, June 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277770.

King:1931:PUP

[Kin31e] Willford I. King. Public utility problems. Journal of the American Statistical Association, 26(174):188–192, June 1931. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277769.

King:1931:SAD

[Kin31f] Willford I. King. The statistical aspects of disease. Journal of the American Statistical Association, 26(174):197–201, June 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277771.

King:1931:WSG

[Kin31g] Willford I. King. What stock groups will lead the next bull market? Journal of the American Statistical Association, 26(173):66–70, March 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278263.

King:1932:BRB

[Kin32a] Willford I. King. Book review: [untitled]. Journal of the American Statistical Association, 27(179):347–348, September 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278181.

King:1932:FLR

[Kin32b] Willford I. King. Financial liquidation and recovery. Journal of the American Statistical Association, 27(177):85–90, March 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277886.

King:1932:FMS

[Kin32c] Willford I. King. Forecasting methods successfully used since 1928. Journal of the American Statistical Association, 27(179):315–319, September 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278168.

King:1932:OPL

[Kin32d] Willford I. King. The outlook for the price level. Journal of the American Statistical Association, 27(180):431–436, December 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278644.

King:1932:PLS

[Kin32e] Willford I. King. The price level slumps of 1920 and 1931 — what caused them? Journal of the American Statistical Association, 27(177):78–84, March 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277885.

King:1932:SMA

[Kin32f] Willford I. King. Statistical methods in advertising research. Journal of the American Statistical Association, 27(179):319–321, September 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2278169.

King:1932:UOE

[Kin32g] Willford I. King. Unemployment — its origin and its elimination. Journal of the American Statistical Association, 27(179):310–315, September 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278167.

King:1933:AVM

[Kin33a] Willford I. King. The adequacy of various methods of relieving the unemployed. Journal of the American Statistical Association, 28(182):210–215, June 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278038.

King:1933:WMM

[Kin33b] Willford I. King. Are we menaced by machines? Journal of the American Statistical Association, 28(182):219–221, June 1933. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278040.

King:1933:BRB

[Kin33c] Willford I. King. Book review: [untitled]. Journal of the American Statistical Association, 28(182):259–260, June 1933. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278059.

King:1933:CGU

[Kin33d] Willford I. King. The cost of government in the United States and its measurement. Journal of the American Statistical Association, 28(181):87–91, March 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278007.

King:1933:ECF

[Kin33e] Willford I. King. Economic change and financial readjustment. Journal of the American Statistical Association, 28(181):82–87, March 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2278006.

King:1933:I

[Kin33f] Willford I. King. Inflation. Journal of the American Statistical Association, 28(183):338–342, September 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278432.

King:1933:NRA

[Kin33g] Willford I. King. The national recovery act and retail prices. Journal of the American Statistical Association, 28(184):440–444, December 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2278136.

King:1933:NDI

[Kin33h] Willford I. King. New developments in insurance revealed by statistics. Journal of the American Statistical Association, 28(182):215– 219, June 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278039.

King:1934:CBB

[Kin34a] Willford I. King. Are consumers being burdened by monopolies created by the NRA? Journal of the American Statistical Association, 29(188):420–421, December 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278615.

King:1934:DMH

[Kin34b] Willford I. King. Dinner meeting in honor of visiting members of the International Statistical Institute. Journal of the American Statistical Association, 29(185):85, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278466.

King:1934:MAB

[Kin34c] Willford I. King. Minutes of the annual business meeting. Journal of the American Statistical Association, 29(185):193–197, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277833.

King:1934:OSE

[Kin34d] Willford I. King. Ought the security exchanges to be regulated by the Federal Government? Journal of the American Statistical Association, 29(186):205–207, June 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278296.

King:1934:OUS

[Kin34e] Willford I. King. Ought the United States to resume lending abroad? Journal of the American Statistical Association, 29(187):320–322, September 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278078.

King:1934:TMF

[Kin34f] Willford I. King. Technical methods of forecasting stock prices. Journal of the American Statistical Association, 29(187):323–325, September 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278079.

King:1935:RME

[Kin35a] Willford I. King. Recent monetary experiments and their effect upon the theory of money and prices. Journal of the American Statistical Association, 30(190):387–400, June 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277690.

King:1935:SFM

[Kin35b] Willford I. King. Should a Federal monetary authority be established to exercise complete control over the currency? Journal of the American Statistical Association, 30(189):99–100, March 1935. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278532.

King:1936:COG

[Kin36a] Willford I. King. Consolidating our gains. Journal of the American Statistical Association, 31(193):1–9, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278317.

King:1936:ES

[Kin36b] Willford I. King. Edgar Sydenstricker. Journal of the American Statistical Association, 31(194):411–414, June 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278574.

Kindleberger:1939:BRB

[Kin39a] Charles P. Kindleberger. Book review: [untitled]. Journal of the American Statistical Association, 34(206):457, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278895.

King:1939:BRB

[Kin39b] Willford I. King. Book review: [untitled]. Journal of the American Statistical Association, 34(206):427–428, June 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278872.

King:1939:CPA

[Kin39c] Willford I. King. Can production of automobiles be stabilized by making their prices flexible? Journal of the American Statistical Association, 34(208):641–651, December 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279807.

King:1939:CDP

[Kin39d] Willford I. King. The concept of demand and price elasticity — the dynamics of automobile demand: a rejoinder. Journal of the American Statistical Association, 34(208):664–665, December 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279809.

Kiser:1932:FSC

[Kis32] Clyde V. Kiser. Fertility of social classes in various types of communities of the east north central states in 1900. Journal of the American Statistical Association, 27(180):371–382, December 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278637.

Kiser:1934:BRB

[Kis34a] Clyde V. Kiser. Book review: [untitled]. Journal of the American Statistical Association, 29(186):226–227, June 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278304.

Kiser:1934:PSP

[Kis34b] Clyde V. Kiser. Pitfalls in sampling for population study. Journal of the American Statistical Association, 29(187):250–256, September 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2278062.

Kiser:1936:BRB

[Kis36] Clyde V. Kiser. Book review: [untitled]. Journal of the American Statistical Association, 31(196):762–765, December 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278686.

Knight:1937:BRB

[Kni37] Frank H. Knight. Book review: [untitled]. Journal of the American Statistical Association, 32(199):600–601, September 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278790.

Knight:1938:BRB

[Kni38] Frank H. Knight. Book review: [untitled]. Journal of the American Statistical Association, 33(204):759–760, December 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279082.

Knight:1939:BRB

[Kni39] M. M. Knight. Book review: [untitled]. Journal of the American Statistical Association, 34(207):609–610, September 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279510.

Koenig:1936:BRB

[Koe36] William H. Koenig. Book review: [untitled]. Journal of the American Statistical Association, 31(194):440–441, June 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278592.

Kolesnikoff:1935:IPM

[Kol35] Vladimir S. Kolesnikoff. Index of production of manufactures derived from census data, 1933. Journal of the American Statistical Association, 30(191):581–583, September 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278102.

Kolesnikoff:1937:IMP

[Kol37] Vladimir S. Kolesnikoff. Index of manufacturing production derived from census data, 1935. Journal of the American Statistical Association, 32(200):713–714, December 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279374.

Kopf:1930:BRB

[Kop30] E. W. Kopf. Book review: [untitled]. Journal of the American Statistical Association, 25(171):379–380, September 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278213.

Kopf:1931:BRB

[Kop31] Edwin W. Kopf. Book review: [untitled]. Journal of the American Statistical Association, 26(174):223–226, June 1931. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277779.

Kopf:1932:BRB

[Kop32a] E. W. Kopf. Book review: [untitled]. Journal of the American Statistical Association, 27(177):115, March 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277898.

Kopf:1932:BDM

[Kop32b] Edwin W. Kopf. Birth, death and marriage rates of large German cities in 1931. Journal of the American Statistical Association, 27(178):192, June 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277669.

Kossoris:1939:SAA

[Kos39] Max D. Kossoris. A statistical approach to accident prevention. Journal of the American Statistical Association, 34(207):524–532, September 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279485.

Kreps:1931:BRB

[Kre31] Theodore J. Kreps. Book review: [untitled]. Journal of the American Statistical Association, 26(176):491–494, December 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277957.

Kreps:1932:BRB

[Kre32] Theodore J. Kreps. Book review: [untitled]. Journal of the American Statistical Association, 27(179):343–344, September 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278177.

Kreps:1934:EUD

[Kre34] Theodore J. Kreps. Estimates of unemployment during the last four years. Journal of the American Statistical Association, 29(185):81–85, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277808.

Krout:1934:BRB

[Kro34] John A. Krout. Book review: [untitled]. Journal of the American Statistical Association, 29(186):240–241, June 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278314.

Kneeland:1936:PSC

[KSF36] Hildegarde Kneeland, Erika H. Schoenberg, and Milton Friedman. Plans for a study of the consumption of goods and services by American families. Journal of the American Statistical Association, 31(193):135–140, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278340.

Kuczynski:1930:RDL

[Kuc30] Robert R. Kuczynski. A reply to Dr. Lotka’s review of “The Balance of Births and Deaths”. Journal of the American Statistical Association, 25(169):84–85, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278502.

Kuczynski:1931:BRB

[Kuc31a] R. R. Kuczynski. Book review: [untitled]. Journal of the American Statistical Association, 26(173):104–105, March 1931. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278273.

Kuczynski:1931:CMS

[Kuc31b] R. R. Kuczynski. Charts and maps of the Social and Economic Museum of Vienna. Journal of the American Statistical Association, 26(173):66, March 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278262.

Kulp:1934:BRB

[Kul34] C. A. Kulp. Book review: [untitled]. Journal of the American Statistical Association, 29(185):104–105, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278474.

Kulp:1938:BRB

[Kul38] C. A. Kulp. Book review: [untitled]. Journal of the American Statistical Association, 33(203):642–644, September 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279727.

Kuvin:1930:SPI

[Kuv30] Leonard Kuvin. Stock price indexes of the New York Curb Exchange. Journal of the American Statistical Association, 25(169): 51–62, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278494.

Kuvin:1936:ERP

[Kuv36] Leonard Kuvin. Effect of N.R.A. on the physical volume of production. Journal of the American Statistical Association, 31(193):58–60, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278324.

Kuznets:1930:BRBa

[Kuz30a] Simon Kuznets. Book review: [untitled]. Journal of the American Statistical Association, 25(171):372–373, September 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278209.

Kuznets:1930:BRBb

[Kuz30b] Simon Kuznets. Book review: [untitled]. Journal of the American Statistical Association, 25(172):486, December 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277981.

Kuznets:1931:BRB

[Kuz31] Simon Kuznets. Book review: [untitled]. Journal of the American Statistical Association, 26(174):231–233, June 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277783.

Kuznets:1932:BRB

[Kuz32a] Simon Kuznets. Book review: [untitled]. Journal of the American Statistical Association, 27(177):116–118, March 1932. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277900.

Kuznets:1932:SPS

[Kuz32b] Simon Kuznets. Seasonal pattern and seasonal amplitude: Measurement of their short-time variations. Journal of the American Statistical Association, 27(177):9–20, March 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277876.

Kuznets:1933:BRB

[Kuz33] Simon Kuznets. Book review: [untitled]. Journal of the American Statistical Association, 28(183):363–364, September 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278442.

Kuznets:1935:AMA

[Kuz35a] Simon Kuznets. The attitudes of members of the American Statistical Association toward the question of calendar reform. Journal of the American Statistical Association, 30(190):437–445, June 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277694.

Kuznets:1935:BRB

[Kuz35b] Simon Kuznets. Book review: [untitled]. Journal of the American Statistical Association, 30(191):642–643, September 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278122.

Kuznets:1938:BRB

[Kuz38a] Simon Kuznets. Book review: [untitled]. Journal of the American Statistical Association, 33(202):456–458, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278926.

Kuznets:1938:DWR

[Kuz38b] Simon Kuznets. Dr. Warburton’s review of national income and capital formation — a comment. Journal of the American Statistical Association, 33(204):714–719, December 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279065.

Kyrk:1930:BRB

[Kyr30] Hazel Kyrk. Book review: [untitled]. Journal of the American Statistical Association, 25(170):239–240, June 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277648.

Kyrk:1933:BRB

[Kyr33] Hazel Kyrk. Book review: [untitled]. Journal of the American Statistical Association, 28(183):369–370, September 1933. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278446.

Kyrk:1937:MMV

[Kyr37] Hazel Kyrk. Methods of measuring variations in family expenditures: Discussion. Journal of the American Statistical Association, 32(197): 47–49, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279860.

Landau:1931:OSP

[Lan31] Ludwik Landau. Organization of statistics in Poland. Journal of the American Statistical Association, 26(173):41–51, March 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278258.

Lang:1933:BRB

[Lan33] Richard O. Lang. Book review: [untitled]. Journal of the American Statistical Association, 28(184):472–474, December 1933. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278148.

Lange:1935:BRB

[Lan35] Oscar Lange. Book review: [untitled]. Journal of the American Statistical Association, 30(191):635–637, September 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278119.

Latimer:1930:BRB

[Lat30] Murray W. Latimer. Book review: [untitled]. Journal of the American Statistical Association, 25(172):479–485, December 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277980.

Latimer:1934:BRB

[Lat34] Murray W. Latimer. Book review: [untitled]. Journal of the American Statistical Association, 29(186):234–235, June 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278310.

Leavens:1931:FDC

[Lea31] Dickson H. Leavens. Frequency distributions corresponding to time series. Journal of the American Statistical Association, 26 (176):407–415, December 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277941.

Leavens:1938:BRB

[Lea38] Dickson H. Leavens. Book review: [untitled]. Journal of the American Statistical Association, 33(204):756–757, December 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279080.

Leavens:1939:BRB

[Lea39] Dickson H. Leavens. Book review: [untitled]. Journal of the American Statistical Association, 34(208):769–770, December 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279835.

Lederer:1936:BRB

[Led36] Emil Lederer. Book review: [untitled]. Journal of the American Statistical Association, 31(195):629–631, September 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278404.

Leong:1932:CSI

[Leo32a] Y. S. Leong. A comparative study of the indexes of production. Journal of the American Statistical Association, 27(179):256–269, September 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278159.

Leong:1932:IPV

[Leo32b] Y. S. Leong. Indexes of the physical volume production of producers’ and consumers’ goods. Journal of the American Statistical Association, 27(177):21–36, March 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277877.

Leong:1935:IPV

[Leo35] Y. S. Leong. Indexes of the physical volume production of producers’ goods, consumers’ goods, durable goods and transient goods. Journal of the American Statistical Association, 30(190):361–376, June 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277688.

Leong:1936:SCU

[Leo36a] Y. S. Leong. Statistics on copper in the United States. Journal of the American Statistical Association, 31(196):665–676, December 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278664.

Leontief:1936:BRB

[Leo36b] Wassily Leontief. Book review: [untitled]. Journal of the American Statistical Association, 31(194):422–423, June 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278580.

Leontief:1937:BRB

[Leo37] Wassily Leontief. Book review: [untitled]. Journal of the American Statistical Association, 32(197):219–220, March 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279882.

Lerner:1938:BRB

[Ler38] A. P. Lerner. Book review: [untitled]. Journal of the American Statistical Association, 33(203):612–614, September 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279713.

Leven:1930:BRB

[Lev30] Maurice Leven. Book review: [untitled]. Journal of the American Statistical Association, 25(169):122–123, March 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278517.

Leven:1937:BRB

[Lev37] Maurice Leven. Book review: [untitled]. Journal of the American Statistical Association, 32(197):242–243, March 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279894.

Lewis:1932:EFN

[Lew32] Edward E. Lewis. Economic factors in Negro migration. Journal of the American Statistical Association, 27(177):45–53, March 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277879.

Lewis:1937:MPP

[Lew37] Edward E. Lewis. The measurement of purchasing power. Journal of the American Statistical Association, 32(198):337–343, June 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279574.

Lewis:1938:BRBb

[Lew38a] Ben W. Lewis. Book review: [untitled]. Journal of the American Statistical Association, 33(203):606–607, September 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279709.

Lewis:1938:BRBa

[Lew38b] Cleona Lewis. Book review: [untitled]. Journal of the American Statistical Association, 33(202):455–456, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278925.

Lewis:1939:BRB

[Lew39] Ben W. Lewis. Book review: [untitled]. Journal of the American Statistical Association, 34(207):625–627, September 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279526.

Linders:1935:PN

[LH35] F. J. Linders and Ralph G. Hurlin. Personal notes. Journal of the American Statistical Association, 30(192):751–754, December 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277738.

Li:1934:SMF

[Li34] Chen-Nan Li. Summation method of fitting parabolic curves and calculating linear and curvilinear correlation coefficients on a scatter-diagram. Journal of the American Statistical Association, 29(188):405–409, December 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278611.

Lindsay:1932:BRB

[Lin32] Samuel McCune Lindsay. Book review: [untitled]. Journal of the American Statistical Association, 27(178):234–236, June 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277686.

Linder:1937:EAN

[Lin37a] Forrest E. Linder. The effect of allocation of nonresident deaths on death rates. Journal of the American Statistical Association, 32(197):105–112, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279868.

Linder:1937:OMN

[Lin37b] Forrest E. Linder. Outline of 1936 mortality and natality tabulations made by the Bureau of the Census. Journal of the American Statistical Association, 32(199):480–488, September 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278768.

Linnenberg:1939:BRB

[Lin39] Clem C. Linnenberg, Jr. Book review: [untitled]. Journal of the American Statistical Association, 34(207):604–605, September 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279506.

Livermore:1938:BRB

[Liv38] Shaw Livermore. Book review: [untitled]. Journal of the American Statistical Association, 33(204):781–782, December 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279097.

Lively:1939:BRB

[Liv39] C. E. Lively. Book review: [untitled]. Journal of the American Statistical Association, 34(206):463–464, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278900.

Lloyd:1938:CDT

[Llo38] Edward L. Lloyd. Current domestic trade statistics. Journal of the American Statistical Association, 33(203):579–582, September 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279699.

Lockley:1936:BRB

[Loc36] Lawrence C. Lockley. Book review: [untitled]. Journal of the American Statistical Association, 31(194):449–450, June 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278600.

Loman:1938:CSD

[Lom38] H. J. Loman. Changes in secondary distribution of equity securities. Journal of the American Statistical Association, 33(201):21–30, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279110.

Lorimer:1935:AIA

[Lor35] Frank Lorimer. Adjustment for the influence of age composition on estimates by Burks’s method of the distribution of sizes of completed fraternities. Journal of the American Statistical Association, 30(192):688–690, December 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277722.

Lorimer:1939:BRB

[Lor39] Frank Lorimer. Book review: [untitled]. Journal of the American Statistical Association, 34(206):441–442, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278882.

Lotka:1930:RDL

[Lot30] Alfred J. Lotka. A reply to Dr. Lotka’s review of “The Balance of Births and Deaths”: Rejoinder. Journal of the American Statistical Association, 25(169):85–86, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278503.

Lotka:1933:O

[Lot33] Alfred J. Lotka. Obituary. Journal of the American Statistical Association, 28(184):456–457, December 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278140.

Lotka:1936:GDI

[Lot36] Alfred J. Lotka. The geographic distribution of intrinsic natural increase in the United States, and an examination of the relation between several measures of net reproductivity. Journal of the American Statistical Association, 31(194):273–294, June 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278558.

Lotka:1938:SRR

[Lot38] Alfred J. Lotka. Some recent results in population analysis. Journal of the American Statistical Association, 33(201):164–178, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279123.

Louden:1939:BRB

[Lou39] Mary V. Louden. Book review: [untitled]. Journal of the American Statistical Association, 34(206):426–427, June 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278871.

Lovitt:1936:BRB

[Lov36] W. V. Lovitt. Book review: [untitled]. Journal of the American Statistical Association, 31(194):425–426, June 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278583.

Lubin:1930:BRB

[Lub30] Isador Lubin. Book review: [untitled]. Journal of the American Statistical Association, 25(172):496–497, December 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277989.

Lubin:1934:WSC

[Lub34] Isador Lubin. Who should collect economic statistics for recovery and planning? Journal of the American Statistical Association, 29(185):180–183, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277830.

Lundberg:1934:BRB

[Lun34] George A. Lundberg. Book review: [untitled]. Journal of the American Statistical Association, 29(187):343, September 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278086.

Lundberg:1935:BRB

[Lun35] George A. Lundberg. Book review: [untitled]. Journal of the American Statistical Association, 30(189):118–120, March 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278540.

Lundberg:1936:BRB

[Lun36a] George A. Lundberg. Book review: [untitled]. Journal of the American Statistical Association, 31(194):423–424, June 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278581.

Lundenberg:1936:BRB

[Lun36b] George A. Lundenberg. Book review: [untitled]. Journal of the American Statistical Association, 31(196):819, December 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278717.

Lundberg:1937:BRB

[Lun37] George A. Lundberg. Book review: [untitled]. Journal of the American Statistical Association, 32(200):813–814, December 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279412.

Lundberg:1938:BRB

[Lun38] George A. Lundberg. Book review: [untitled]. Journal of the American Statistical Association, 33(204):787–788, December 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279101.

Lundberg:1939:BRB

[Lun39] George A. Lundberg. Book review: [untitled]. Journal of the American Statistical Association, 34(208):752–753, December 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279821.

Lutz:1933:BRB

[Lut33] H. L. Lutz. Book review: [untitled]. Journal of the American Statistical Association, 28(181):110–111, March 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278017.

Lutz:1937:BRB

[Lut37] H. L. Lutz. Book review: [untitled]. Journal of the American Statistical Association, 32(197):244–245, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279896.

Lutz:1939:BRB

[Lut39] F. A. Lutz. Book review: [untitled]. Journal of the American Statistical Association, 34(207):607, September 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279508.

Lyon:1931:BRB

[Lyo31] Leverett S. Lyon. Book review: [untitled]. Journal of the American Statistical Association, 26(173):111–112, March 1931. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278278.

Lyon:1936:BRB

[Lyo36] Leverett S. Lyon. Book review: [untitled]. Journal of the American Statistical Association, 31(196):780–783, December 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278697.

MacIver:1933:BRB

[Mac33] R. M. MacIver. Book review: [untitled]. Journal of the American Statistical Association, 28(181):104–106, March 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278014.

MacNeil:1938:BRB

[Mac38] Douglas H. MacNeil. Book review: [untitled]. Journal of the American Statistical Association, 33(203):637–639, September 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279725.

Machlup:1939:BRB

[Mac39a] Fritz Machlup. Book review: [untitled]. Journal of the American Statistical Association, 34(205):186–188, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279175.

Maclaurin:1939:BRB

[Mac39b] W. Rupert Maclaurin. Book review: [untitled]. Journal of the American Statistical Association, 34(205):222–223, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279202.

Madow:1938:BRB

[Mad38] William G. Madow. Book review: [untitled]. Journal of the American Statistical Association, 33(203):622–623, September 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279717.

Magnusson:1933:IIC

[Mag33] Leifur Magnusson. An international inquiry into costs of living: a comparative study of workers’ living costs in Detroit and fourteen European cities. Journal of the American Statistical Association, 28(182):123–138, June 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278027.

Magdoff:1939:PMM

[Mag39] Harry Magdoff. The purpose and method of measuring productivity. Journal of the American Statistical Association, 34(206):309–318, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278850.

Maher:1930:ECI

[Mah30] Amy G. Maher. Effect of concentration of industry on seasonal variation of employment. Journal of the American Statistical Association, 25(169):41–46, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278492.

Malott:1935:BRB

[Mal35] Deane W. Malott. Book review: [untitled]. Journal of the American Statistical Association, 30(190):498, June 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277715.

Malzberg:1936:RMD

[Mal36] Benjamin Malzberg. Rates of mental disease among certain population groups in New York State. Journal of the American Statistical Association, 31(195):545–548, September 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278381.

Mangus:1934:SFR

[Man34] A. Raymond Mangus. Sampling in the field of rural relief. Journal of the American Statistical Association, 29(188):410–415, December 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2278612.

Mann:1945:NPC

[Man45] H. B. Mann. Note on a paper by C. W. Cotterman and L. H. Snyder. Annals of Mathematical Statistics, 16(3):311–312, September 1945. CODEN AASTAD. ISSN 0003-4851. URL http://projecteuclid.

org/euclid.aoms/1177731097. See [CS39].

Marget:1930:BRB

[Mar30] Arthur W. Marget. Book review: [untitled]. Journal of the American Statistical Association, 25(169):104–107, March 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278508.

Marget:1933:BRB

[Mar33] Arthur W. Marget. Book review: [untitled]. Journal of the American Statistical Association, 28(182):243–244, June 1933. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278050.

Marget:1934:BRB

[Mar34a] Arthur W. Marget. Book review: [untitled]. Journal of the American Statistical Association, 29(187):349–350, September 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278091.

Marshall:1934:JCS

[Mar34b] L. C. Marshall. Judicial criminal statistics (abstract). Journal of the American Statistical Association, 29(185):99, March 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277811.

Marget:1935:BRBa

[Mar35a] Arthur W. Marget. Book review: [untitled]. Journal of the American Statistical Association, 30(189):122–123, March 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278542.

Marget:1935:BRBb

[Mar35b] Arthur W. Marget. Book review: [untitled]. Journal of the American Statistical Association, 30(192):758–760, December 1935. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277741.

Marget:1939:BRB

[Mar39a] Arthur W. Marget. Book review: [untitled]. Journal of the American Statistical Association, 34(205):189–190, March 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279177.

Martin:1939:BRBa

[Mar39b] James W. Martin. Book review: [untitled]. Journal of the American Statistical Association, 34(205):200–201, March 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279185.

Martin:1939:BRBb

[Mar39c] Robert F. Martin. Book review: [untitled]. Journal of the American Statistical Association, 34(206):437–439, June 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278880.

Mason:1933:BRB

[Mas33] Edward S. Mason. Book review: [untitled]. Journal of the American Statistical Association, 28(184):462–463, December 1933. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278143.

Mason:1935:BRB

[Mas35] Edward S. Mason. Book review: [untitled]. Journal of the American Statistical Association, 30(190):487–489, June 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277708.

Mason:1936:BRB

[Mas36] Edward S. Mason. Book review: [untitled]. Journal of the American Statistical Association, 31(195):641–642, September 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278410.

Mason:1937:BRB

[Mas37] Edward S. Mason. Book review: [untitled]. Journal of the American Statistical Association, 32(199):602–604, September 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278792.

Mason:1938:BRB

[Mas38] Edward S. Mason. Book review: [untitled]. Journal of the American Statistical Association, 33(203):607–608, September 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279710.

Maverick:1932:GPS

[Mav32] Lewis A. Maverick. Graphic presentation of standard deviation. Journal of the American Statistical Association, 27(179):287–297, September 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278162.

McCormick:1934:CID

[McC34] Thomas C. McCormick. A coefficient of independent determination. Journal of the American Statistical Association, 29(185):76–78, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278463.

McCabe:1936:BRB

[McC36a] George K. McCabe. Book review: [untitled]. Journal of the American Statistical Association, 31(196):805–806, December 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278709.

McCahan:1936:BRB

[McC36b] David McCahan. Book review: [untitled]. Journal of the American Statistical Association, 31(193):237–238, March 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278362.

McDiarmid:1938:BRB

[McD38] O. J. McDiarmid. Book review: [untitled]. Journal of the American Statistical Association, 33(202):485–487, June 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278943.

McDiarmid:1939:BRB

[McD39a] John O. McDiarmid. Book review: [untitled]. Journal of the American Statistical Association, 34(205):209–210, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279192.

McDonough:1939:BRB

[McD39b] J. E. McDonough. Book review: [untitled]. Journal of the American Statistical Association, 34(207):619–620, September 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279520.

McFall:1932:CD

[McF32] Robert J. McFall. Census of distribution. Journal of the American Statistical Association, 27(179):302–306, September 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278164.

McIntyre:1936:BRB

[McI36] Francis McIntyre. Book review: [untitled]. Journal of the American Statistical Association, 31(195):619–620, September 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278398.

McIntyre:1937:ACC

[McI37] Francis McIntyre. Automatic checks in correlation analysis. Journal of the American Statistical Association, 32(197):119–123, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279870.

McIntyre:1938:BRB

[McI38a] Francis McIntyre. Book review: [untitled]. Journal of the American Statistical Association, 33(204):751–752, December 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279076.

McIntyre:1938:PSP

[McI38b] Francis McIntyre. The problem of the stock price index number. Journal of the American Statistical Association, 33(203):557–563, September 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279694.

McKay:1936:BRB

[McK36] Evelyn C. McKay. Book review: [untitled]. Journal of the American Statistical Association, 31(193):230–231, March 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278357.

McLaughlin:1932:ICD

[McL32] Glenn E. McLaughlin. The inadequacy of census data for individual industries. Journal of the American Statistical Association, 27(177): 37–44, March 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277878.

McLaughlin:1936:BRB

[McL36] Glenn E. McLaughlin. Book review: [untitled]. Journal of the American Statistical Association, 31(193):220–223, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278351.

McLaughlin:1938:BRB

[McL38] Glenn E. McLaughlin. Book review: [untitled]. Journal of the American Statistical Association, 33(201):287–288, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279144.

McLaughlin:1939:BRB

[McL39] Glenn E. McLaughlin. Book review: [untitled]. Journal of the American Statistical Association, 34(205):231–232, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279208.

Mead:1931:BRB

[Mea31] E. S. Mead. Book review: [untitled]. Journal of the American Statistical Association, 26(173):98–101, March 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278270.

Means:1933:BRB

[Mea33] Gardiner C. Means. Book review: [untitled]. Journal of the American Statistical Association, 28(182):251–253, June 1933. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278054.

Means:1935:PIR

[Mea35a] Gardiner C. Means. Price inflexibility and the requirements of a stabilizing monetary policy. Journal of the American Statistical Association, 30(190):401–413, June 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277691.

Mears:1935:FTS

[Mea35b] Eliot G. Mears. The foreign trade statistics of the United States. Journal of the American Statistical Association, 30(191):501–516, September 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278096.

Means:1937:BRB

[Mea37] Gardiner C. Means. Book review: [untitled]. Journal of the American Statistical Association, 32(200):777–781, December 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279392.

Menge:1930:BRB

[Men30] Walter O. Menge. Book review: [untitled]. Journal of the American Statistical Association, 25(171):382–383, September 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278216.

Menge:1932:BRB

[Men32] W. O. Menge. Book review: [untitled]. Journal of the American Statistical Association, 27(179):344–345, September 1932. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278178.

Menge:1935:BRB

[Men35] W. O. Menge. Book review: [untitled]. Journal of the American Statistical Association, 30(192):777, December 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277752.

Mendershausen:1939:BRB

[Men39a] Horst Mendershausen. Book review: [untitled]. Journal of the American Statistical Association, 34(206):433–434, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278877.

Mendershausen:1939:CVC

[Men39b] Horst Mendershausen. Clearing variates in confluence analysis. Journal of the American Statistical Association, 34(205):93–105, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279167.

Meriam:1933:BRB

[Mer33] R. S. Meriam. Book review: [untitled]. Journal of the American Statistical Association, 28(183):362–363, September 1933. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278441.

Methorst:1936:NSP

[Met36] H. W. Methorst. The new system of population accounting in The Netherlands. Journal of the American Statistical Association, 31 (196):719–722, December 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278669.

Methorst:1938:NSP

[Met38] H. W. Methorst. The new system of population accounting in The Netherlands. Journal of the American Statistical Association, 33 (204):713–714, December 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279064.

Meyer:1938:BRB

[Mey38] Gerhard Meyer. Book review: [untitled]. Journal of the American Statistical Association, 33(204):762–764, December 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279085.

McClure:1935:ADC

[MF35] Howard H. McClure and Richard L. Funkhouser. The activities of the division of current business statistics of the central statistical board. Journal of the American Statistical Association, 30(192):701–706, December 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277725.

Mudgett:1934:RFS

[MG34] Bruce D. Mudgett and S. R. Gevorkiantz. Reliability of forest surveys. Journal of the American Statistical Association, 29(187):257–281, September 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278063.

Myers:1939:SCS

[MH39] R. J. Myers and A. Herschfeld. The summation check in statistical calculations. Journal of the American Statistical Association, 34(207):545–548, September 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279488.

Mills:1930:BRB

[Mil30] Frederick C. Mills. Book review: [untitled]. Journal of the American Statistical Association, 25(170):231, June 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277642.

Mills:1935:SL

[Mil35] Frederick C. Mills. Statistics and Leviathan. Journal of the American Statistical Association, 30(189):1–11, March 1935. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278525.

Mills:1937:IPP

[Mil37a] Frederick C. Mills. Industrial productivity and prices. Journal of the American Statistical Association, 32(198):247–262, June 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279567.

Mills:1937:MMO

[Mil37b] Frederick C. Mills. Movements of mail-order prices. Journal of the American Statistical Association, 32(197):131–132, March 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279872.

Millison:1938:BRB

[Mil38] Earl G. Millison. Book review: [untitled]. Journal of the American Statistical Association, 33(202):489–490, June 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278945.

Miner:1930:BRB

[Min30] John R. Miner. Book review: [untitled]. Journal of the American Statistical Association, 25(169):113–114, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278511.

Miner:1933:BRB

[Min33] John R. Miner. Book review: [untitled]. Journal of the American Statistical Association, 28(182):238–239, June 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278047.

Miner:1935:BRB

[Min35] John R. Miner. Book review: [untitled]. Journal of the American Statistical Association, 30(192):777–778, December 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277753.

Mitchell:1930:BRB

[Mit30] Broadus Mitchell. Book review: [untitled]. Journal of the American Statistical Association, 25(172):500–501, December 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277993.

Mitchell:1932:BRB

[Mit32] George Sinclair Mitchell. Book review: [untitled]. Journal of the American Statistical Association, 27(179):346–347, September 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278180.

Mitchell:1936:BRB

[Mit36] Broadus Mitchell. Book review: [untitled]. Journal of the American Statistical Association, 31(196):811–812, December 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278713.

Mitchell:1939:CII

[Mit39a] Walter Mitchell, Jr. Collection and interpretation of inventory data. Journal of the American Statistical Association, 34(206):283–290, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278846.

Mitchell:1939:FAR

[Mit39b] Walter Mitchell, Jr. Factors affecting the rate of return on mailed questionnaires. Journal of the American Statistical Association, 34 (208):683–692, December 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279813.

Mitchell:1939:IWH

[Mit39c] Walter Mitchell, Jr. Industrial wage and hour surveys for management guidance. Journal of the American Statistical Association, 34(207):482–491, September 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279482.

Monroe:1937:AFC

[MK37] Day Monroe and Hazel Kyrk. Analyzing families by composition type with respect to consumption. Journal of the American Statistical Association, 32(197):35–39, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279858.

Martin:1931:NGF

[ML31] F. C. Martin and Dickson H. Leavens. A new grid for fitting a normal probability curve to a given frequency distribution. Journal of the American Statistical Association, 26(174):178–183, June 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277767.

Montgomery:1938:BRB

[Mon38] Royal E. Montgomery. Book review: [untitled]. Journal of the American Statistical Association, 33(201):288–289, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279145.

Moore:1937:BRB

[Moo37] O. Ernest Moore. Book review: [untitled]. Journal of the American Statistical Association, 32(199):613–615, September 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278798.

Morrison:1936:BRB

[Mor36a] L. A. Morrison. Book review: [untitled]. Journal of the American Statistical Association, 31(196):776–778, December 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278695.

Morton:1936:BRB

[Mor36b] W. A. Morton. Book review: [untitled]. Journal of the American Statistical Association, 31(196):802–803, December 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278707.

Morse:1937:GPC

[Mor37] John W. Morse. Geometric presentation of correlation. Journal of the American Statistical Association, 32(198):364–365, June 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279580.

Morley:1939:BRB

[Mor39] Burton R. Morley. Book review: [untitled]. Journal of the American Statistical Association, 34(205):218–221, March 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279200.

Mosak:1937:SDE

[Mos37] Jacob L. Mosak. On the simultaneous determination of the elementary regressions and their standard errors in subsets of variables. Journal of the American Statistical Association, 32(199):517–524, September 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278771.

Mosak:1938:CSD

[Mos38] Jacob L. Mosak. A correction: On the simultaneous determination of the elementary regressions and their standard errors in subsets of variables. Journal of the American Statistical Association, 33(201): 232, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279130.

Mosak:1939:LSS

[Mos39] Jacob L. Mosak. The least-squares standard error of the coefficient of elasticity of demand. Journal of the American Statistical Association, 34(206):353–361, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278857.

Meyer:1935:ICD

[MTDD35] H. Arthur Meyer, Sc. Techn, D. F. Deming, and W. Edwards Deming. On the influence of classification on the determination of a measurable characteristic. Journal of the American Statistical Association, 30(192):671–677, December 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277720.

Mudgett:1937:BRB

[Mud37] Bruce D. Mudgett. Book review: [untitled]. Journal of the American Statistical Association, 32(199):582–586, September 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278781.

Mudgett:1939:BRB

[Mud39] Bruce D. Mudgett. Book review: [untitled]. Journal of the American Statistical Association, 34(206):416–419, June 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278865.

Muench:1934:DRS

[Mue34] Hugo Muench. Derivation of rates from summation data by the catalytic curve. Journal of the American Statistical Association, 29(185): 25–38, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278457.

Muench:1936:PDP

[Mue36] Hugo Muench. The probability distribution of protection test results. Journal of the American Statistical Association, 31(196):677–689, December 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278665.

Muench:1938:DFD

[Mue38] H. Muench. Discrete frequency distributions arising from mixtures of several single probability values. Journal of the American Statistical Association, 33(202):390–398, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278916.

Musselman:1933:BRB

[Mus33] J. R. Musselman. Book review: [untitled]. Journal of the American Statistical Association, 28(183):371, September 1933. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278448.

Myers:1934:USB

[Mye34] Howard B. Myers. Using state bureaus for the joint collection of labor statistics. Journal of the American Statistical Association, 29(185):140–145, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277820.

Myers:1939:BRB

[Mye39] Margaret G. Myers. Book review: [untitled]. Journal of the American Statistical Association, 34(206):449, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278888.

Nadler:1939:BRB

[Nad39] Marcus Nadler. Book review: [untitled]. Journal of the American Statistical Association, 34(205):199, March 1939. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279184.

Nagelberg:1939:BRB

[Nag39] Morton S. Nagelberg. Book review: [untitled]. Journal of the American Statistical Association, 34(205):207–208, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279191.

Nathan:1935:BRB

[Nat35] Otto Nathan. Book review: [untitled]. Journal of the American Statistical Association, 30(192):760–762, December 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277742.

Nathan:1937:BRB

[Nat37] Robert R. Nathan. Book review: [untitled]. Journal of the American Statistical Association, 32(200):765–766, December 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279385.

Nathan:1939:BRB

[Nat39] Robert R. Nathan. Book review: [untitled]. Journal of the American Statistical Association, 34(206):435–437, June 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278879.

Nelson:1939:BRB

[Nel39] ElRoy Nelson. Book review: [untitled]. Journal of the American Statistical Association, 34(207):620–621, September 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279521.

Neprash:1934:SPC

[Nep34] Jerry A. Neprash. Some problems in the correlation of spatially distributed variables. Journal of the American Statistical Association, 29(185):167–168, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277826.

Newcomer:1938:BRB

[New38] Mabel Newcomer. Book review: [untitled]. Journal of the American Statistical Association, 33(201):298–299, March 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279151.

Neyman:1935:DIA

[Ney35] J. Neyman. The development of industrial applications of statistical theory in Poland. Journal of the American Statistical Association, 30(192):707–708, December 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277726.

Neyman:1938:CTS

[Ney38a] J. Neyman. Contribution to the theory of sampling human populations. Journal of the American Statistical Association, 33(201):101– 116, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279117.

Neyman:1938:MWG

[Ney38b] J. Neyman. Mr. W. S. Gosset. Journal of the American Statistical Association, 33(201):226–228, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279125.

Neyman:1939:BRB

[Ney39] J. Neyman. Book review: [untitled]. Journal of the American Statistical Association, 34(208):755, December 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279823.

Nichols:1939:BRB

[Nic39] Bervard Nichols. Book review: [untitled]. Journal of the American Statistical Association, 34(207):603–604, September 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279505.

Nixon:1938:ICS

[Nix38] J. W. Nixon. An international convention on statistics of wages and hours of work. Journal of the American Statistical Association, 33(204):721–722, December 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279067.

Nagel:1936:MP

[NMDW36] Ernest Nagel, Henry Margenau, C. J. Ducasse, and S. S. Wilks. The meaning of probability. Journal of the American Statistical Association, 31(193):10–30, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278318.

Neyman:1936:ESK

[NT36] J. Neyman and B. Tokarska. Errors of the second kind in testing “Student’s” hypothesis. Journal of the American Statistical Association, 31(194):318–326, June 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278560.

Nugent:1938:TCF

[Nug38] Rolf Nugent. Tendencies in consumer financing. Journal of the American Statistical Association, 33(201):42–50, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279112.

Nybolle:1937:ARI

[Nyb37] Hans Cl. Nybolle. On an attempt to reduce infant mortality in Denmark. Journal of the American Statistical Association, 32 (199):489–508, September 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278769.

Nystrom:1934:BRB

[Nys34] Paul H. Nystrom. Book review: [untitled]. Journal of the American Statistical Association, 29(185):122–123, March 1934. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278486.

Ogburn:1932:SA

[Ogb32] William F. Ogburn. Statistics and art. Journal of the American Statistical Association, 27(177):1–8, March 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277875.

Ogburn:1935:FVC

[Ogb35] William F. Ogburn. Factors in the variation of crime among cities. Journal of the American Statistical Association, 30(189):12–34, March 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278526.

Orchard:1930:BRB

[Orc30] John E. Orchard. Book review: [untitled]. Journal of the American Statistical Association, 25(172):488–489, December 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277983.

Orchard:1931:BRBa

[Orc31a] Dorothy J. Orchard. Book review: [untitled]. Journal of the American Statistical Association, 26(175):356–357, September 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277919.

Orchard:1931:BRBb

[Orc31b] John E. Orchard. Book review: [untitled]. Journal of the American Statistical Association, 26(175):361–363, September 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277923.

Orchard:1933:BRBa

[Orc33a] John E. Orchard. Book review: [untitled]. Journal of the American Statistical Association, 28(184):484–486, December 1933. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278155.

Orchard:1933:BRBb

[Orc33b] John E. Orchard. Book review: [untitled]. Journal of the American Statistical Association, 28(184):486, December 1933. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278156.

Orchard:1934:BRB

[Orc34] John E. Orchard. Book review: [untitled]. Journal of the American Statistical Association, 29(187):340–343, September 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278085.

Orchard:1935:BRB

[Orc35] John E. Orchard. Book review: [untitled]. Journal of the American Statistical Association, 30(189):137, March 1935. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278551.

Orchard:1936:BRB

[Orc36] John E. Orchard. Book review: [untitled]. Journal of the American Statistical Association, 31(194):443–444, June 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278595.

Orchard:1937:BRB

[Orc37] John E. Orchard. Book review: [untitled]. Journal of the American Statistical Association, 32(199):615–617, September 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278799.

Ostrolenk:1931:BRB

[Ost31] Bernhard Ostrolenk. Book review: [untitled]. Journal of the American Statistical Association, 26(176):497–498, December 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277961.

Palyi:1936:BRB

[Pal36] Melchior Palyi. Book review: [untitled]. Journal of the American Statistical Association, 31(196):783–784, December 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278698.

Palmer:1938:BRB

[Pal38a] Edgar Z. Palmer. Book review: [untitled]. Journal of the American Statistical Association, 33(203):628–630, September 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279722.

Palyi:1938:BRB

[Pal38b] Melchior Palyi. Book review: [untitled]. Journal of the American Statistical Association, 33(202):465–467, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278931.

Palmer:1939:BRB

[Pal39a] Edgar Z. Palmer. Book review: [untitled]. Journal of the American Statistical Association, 34(206):434–435, June 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278878.

Palmer:1939:CLO

[Pal39b] Gladys L. Palmer. The convertibility list of occupations and the problems of developing it. Journal of the American Statistical Association, 34(208):693–708, December 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279814.

Palyi:1939:BRBa

[Pal39c] Melchior Palyi. Book review: [untitled]. Journal of the American Statistical Association, 34(205):192–193, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279179.

Palyi:1939:BRBb

[Pal39d] Melchior Palyi. Book review: [untitled]. Journal of the American Statistical Association, 34(206):455–456, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278893.

Paris:1937:BRB

[Par37a] James D. Paris. Book review: [untitled]. Journal of the American Statistical Association, 32(200):769–772, December 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279387.

Parry:1937:BRB

[Par37b] Corliss L. Parry. Book review: [untitled]. Journal of the American Statistical Association, 32(200):766–769, December 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279386.

Parr:1938:BRB

[Par38] Thaddeus Parr. Book review: [untitled]. Journal of the American Statistical Association, 33(202):490–492, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278946.

Paris:1939:BRB

[Par39] James D. Paris. Book review: [untitled]. Journal of the American Statistical Association, 34(207):622–623, September 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279523.

Paton:1930:BRB

[Pat30] W. A. Paton. Book review: [untitled]. Journal of the American Statistical Association, 25(171):363–366, September 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278206.

Patterson:1934:BRBa

[Pat34a] Ernest Minor Patterson. Book review: [untitled]. Journal of the American Statistical Association, 29(185):119–121, March 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278484.

Patterson:1934:BRBb

[Pat34b] Ernest Minor Patterson. Book review: [untitled]. Journal of the American Statistical Association, 29(188):453–455, December 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278634.

Patterson:1937:BRB

[Pat37] Ernest Minor Patterson. Book review: [untitled]. Journal of the American Statistical Association, 32(200):755–757, December 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279382.

Patterson:1939:SBS

[Pat39] Donald Patterson. A statistical basis for setting wage rates. Journal of the American Statistical Association, 34(206):365–368, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278859.

Pearson:1931:TSC

[Pea31] Egon S. Pearson. The test of significance for the correlation coefficient. Journal of the American Statistical Association, 26(174):128– 134, June 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277761.

Pearson:1932:TSC

[Pea32] Egon S. Pearson. The test of significance for the correlation coefficient: Some further results. Journal of the American Statistical Association, 27(180):424–426, December 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278641.

Pearl:1936:BRB

[Pea36a] Raymond Pearl. Book review: [untitled]. Journal of the American Statistical Association, 31(195):608–609, September 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278394.

Pearl:1936:KP

[Pea36b] Raymond Pearl. Karl Pearson, 1857–1936. Journal of the American Statistical Association, 31(196):653–664, December 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278663.

Pearson:1936:SMI

[Pea36c] Egon S. Pearson. Statistical method and industry in Great Britain. Journal of the American Statistical Association, 31(194):361–366, June 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278563.

Pearson:1936:BRB

[Pea36d] F. A. Pearson. Book review: [untitled]. Journal of the American Statistical Association, 31(196):800–802, December 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278706.

Peake:1938:BRB

[Pea38] Cyrus H. Peake. Book review: [untitled]. Journal of the American Statistical Association, 33(202):470–471, June 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278934.

Peck:1931:BRB

[Pec31] Gustav Peck. Book review: [untitled]. Journal of the American Statistical Association, 26(175):367–368, September 1931. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277926.

Peck:1933:BRB

[Pec33] Gustav Peck. Book review: [untitled]. Journal of the American Statistical Association, 28(183):370, September 1933. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278447.

Peek:1937:TOD

[Pee37] R. L. Peek, Jr. Test of an observed difference in the frequency of two results. Journal of the American Statistical Association, 32 (199):532–536, September 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278773.

Penrose:1931:BRB

[Pen31] E. F. Penrose. Book review: [untitled]. Journal of the American Statistical Association, 26(175):364–367, September 1931. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277925.

Penrose:1932:BRB

[Pen32a] E. F. Penrose. Book review: [untitled]. Journal of the American Statistical Association, 27(180):453–455, December 1932. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278655.

Penrose:1932:OL

[Pen32b] E. F. Penrose. An open letter. Journal of the American Statistical Association, 27(178):186–187, June 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277664.

Person:1931:BRB

[Per31] H. S. Person. Book review: [untitled]. Journal of the American Statistical Association, 26(174):240–241, June 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277791.

Perlman:1934:SPC

[Per34] Jacob Perlman. Some problems in the construction of an index number of rents. Journal of the American Statistical Association, 29(185): 24–29, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277797.

Perlman:1936:TFA

[Per36] Jacob Perlman. Training field agents in the Bureau of Labor Statistics. Journal of the American Statistical Association, 31 (196):735–736, December 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278675.

Perlman:1937:BRB

[Per37] S. Perlman. Book review: [untitled]. Journal of the American Statistical Association, 32(197):232–233, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279888.

Peterson:1933:BRB

[Pet33] Shorey Peterson. Book review: [untitled]. Journal of the American Statistical Association, 28(184):477–479, December 1933. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278151.

Peterson:1937:MUS

[Pet37] Florence Peterson. Methods used in strike statistics. Journal of the American Statistical Association, 32(197):90–96, March 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279866.

Peterson:1939:BRB

[Pet39] Florence Peterson. Book review: [untitled]. Journal of the American Statistical Association, 34(207):633–635, September 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279530.

Phelps:1931:BRB

[Phe31] Roswell F. Phelps. Book review: [untitled]. Journal of the American Statistical Association, 26(176):494–496, December 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277959.

Phelps:1932:MNb

[Phe32] Roswell F. Phelps. Miscellaneous notes. Journal of the American Statistical Association, 27(179):321–333, September 1932. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278170.

Phelps:1934:MN

[Phe34a] Roswell F. Phelps. Miscellaneous notes. Journal of the American Statistical Association, 29(185):89–99, March 1934. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278469.

Phelps:1934:SFC

[Phe34b] Roswell F. Phelps. State and Federal cooperation in the collection of industrial statistics. Journal of the American Statistical Association, 29(185):135–139, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277819.

Phelps:1937:BRB

[Phe37a] D. M. Phelps. Book review: [untitled]. Journal of the American Statistical Association, 32(200):795–796, December 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279400.

Phelps:1937:LDP

[Phe37b] Roswell F. Phelps. Leroy D. Peavey. Journal of the American Statistical Association, 32(200):753, December 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279380.

Phelps:1939:CFG

[Phe39] Roswell F. Phelps. Charles Ferris Gettemy. Journal of the American Statistical Association, 34(208):750–751, December 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279820.

Pearson:1930:TC

[PHT^{+}30] Karl Pearson, J. Arthur Harris, Alan E. Treloar, Marian Wilder, and Karl Pearson. On the theory of contingency. Journal of the American Statistical Association, 25(171):320–327, September 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278196.

Piser:1932:MCW

[Pis32] Leroy M. Piser. A method of calculating weekly seasonal indexes. Journal of the American Statistical Association, 27(179):307–309, September 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278165.

Piser:1934:ATD

[Pis34] Leroy M. Piser. The adjustment of time data for the influence of Easter. Journal of the American Statistical Association, 29(186):190– 191, June 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278292.

Phelps:1932:MNa

[PLE^{+}32] Roswell F. Phelps, Bernice Lamb, D. C. Elliott, A. W. Lohr, and H. S. Davis. Miscellaneous notes. Journal of the American Statistical Association, 27(177):92–99, March 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277888.

Pribram:1935:BRB

[Pri35] Karl Pribram. Book review: [untitled]. Journal of the American Statistical Association, 30(192):784–785, December 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277758.

Proudfoot:1939:BRBa

[Pro39a] M. J. Proudfoot. Book review: [untitled]. Journal of the American Statistical Association, 34(208):776–777, December 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279840.

Proudfoot:1939:BRBb

[Pro39b] M. J. Proudfoot. Book review: [untitled]. Journal of the American Statistical Association, 34(208):783, December 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279847.

Perry:1930:IFG

[PS30] F. Gardiner Perry and A. G. Silverman. The “ideal” formula for “general purpose” index numbers. Journal of the American Statistical Association, 25(169):63–68, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278495.

Puffer:1939:BRB

[Puf39] Claude E. Puffer. Book review: [untitled]. Journal of the American Statistical Association, 34(205):210–211, March 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279194.

Raymond:1936:BCI

[Ray36] Fairfield E. Raymond. Bases of control for industrial operation. Journal of the American Statistical Association, 31(193):37–42, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2278320.

Reagh:1934:SCM

[Rea34] Russell R. Reagh. A simplified code for multiple card punching. Journal of the American Statistical Association, 29(186):182–183, June 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2278289.

Reed:1933:BRB

[Ree33] Louis S. Reed. Book review: [untitled]. Journal of the American Statistical Association, 28(183):372–374, September 1933. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278450.

Reed:1935:BRB

[Ree35] Harold L. Reed. Book review: [untitled]. Journal of the American Statistical Association, 30(192):762–763, December 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277743.

Reed:1937:BRB

[Ree37a] Lowell J. Reed. Book review: [untitled]. Journal of the American Statistical Association, 32(200):820–821, December 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279416.

Reed:1937:CVC

[Ree37b] Vergil D. Reed. Comparability versus change. Journal of the American Statistical Association, 32(200):625–635, December 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279367.

Reed:1939:BRB

[Ree39] Louis S. Reed. Book review: [untitled]. Journal of the American Statistical Association, 34(207):631–632, September 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279528.

Reynolds:1939:PGP

[Rey39] Lloyd G. Reynolds. Producers’ goods prices in expansion and decline. Journal of the American Statistical Association, 34(205):32–40, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279159.

Rice:1930:CAS

[RG30] Stuart A. Rice and Morris Green. Composition of the American statistical association. Journal of the American Statistical Association, 25(170):198–202, June 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277633.

Richey:1930:ESA

[Ric30a] F. D. Richey. Errata: Some applications of statistical methods to agronomic experiments. Journal of the American Statistical Association, 25(172):468, December 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277975.

Richey:1930:SAS

[Ric30b] Frederick D. Richey. Some applications of statistical methods to agronomic experiments. Journal of the American Statistical Association, 25(171):269–283, September 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278190.

Richards:1931:ASE

[Ric31] Henry I. Richards. Analysis of the spurious effect of high intercorrelation of independent variables on regression and correlation coefficients. Journal of the American Statistical Association, 26(173): 21–29, March 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278255.

Rice:1933:CGS

[Ric33a] Stuart A. Rice. Committee on government statistics and information services. Journal of the American Statistical Association, 28 (183):333–334, September 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278429.

Rice:1933:PL

[Ric33b] Stuart A. Rice. A presidential letter. Journal of the American Statistical Association, 28(182):206–208, June 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278036.

Rice:1934:SOR

[Ric34] Stuart A. Rice. Statistical opportunities and responsibilities. Journal of the American Statistical Association, 29(185):1–10, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278455.

Ricker:1937:CCF

[Ric37] William E. Ricker. The concept of confidence or fiducial limits applied to the Poisson frequency distribution. Journal of the American Statistical Association, 32(198):349–356, June 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279576.

Rice:1938:QMP

[Ric38a] Stuart A. Rice. Quantitative methods in politics. Journal of the American Statistical Association, 33(201):126–130, March 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279119.

Richardson:1938:BRB

[Ric38b] A. H. Richardson. Book review: [untitled]. Journal of the American Statistical Association, 33(203):623–624, September 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279718.

Rider:1931:NSS

[Rid31] Paul R. Rider. A note on small sample theory. Journal of the American Statistical Association, 26(174):172–174, June 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277765.

Rider:1935:RPS

[Rid35] Paul R. Rider. Recent progress in statistical method. Journal of the American Statistical Association, 30(189):58–88, March 1935. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278529.

Rider:1937:BRB

[Rid37] Paul R. Rider. Book review: [untitled]. Journal of the American Statistical Association, 32(200):807–808, December 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279407.

Rietz:1930:BRB

[Rie30] H. L. Rietz. Book review: [untitled]. Journal of the American Statistical Association, 25(169):119–120, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278515.

Rietz:1931:BRBa

[Rie31a] H. L. Rietz. Book review: [untitled]. Journal of the American Statistical Association, 26(173):110–111, March 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278277.

Rietz:1931:BRBb

[Rie31b] H. L. Rietz. Book review: [untitled]. Journal of the American Statistical Association, 26(175):360, September 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277922.

Rietz:1931:CAR

[Rie31c] H. L. Rietz. Comments on applications of recently developed theory of small samples. Journal of the American Statistical Association, 26(174):150–158, June 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277763.

Riefler:1932:BRB

[Rie32] Winfield W. Riefler. Book review: [untitled]. Journal of the American Statistical Association, 27(180):456–457, December 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278658.

Rietz:1937:BRB

[Rie37] H. L. Rietz. Book review: [untitled]. Journal of the American Statistical Association, 32(200):803–805, December 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279405.

Riggleman:1933:BRB

[Rig33a] John R. Riggleman. Book review: [untitled]. Journal of the American Statistical Association, 28(182):258–259, June 1933. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278058.

Riggleman:1933:BCU

[Rig33b] John R. Riggleman. Building cycles in the United States, 1875– 1932. Journal of the American Statistical Association, 28(182):174– 183, June 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278031.

Riggleman:1939:BRB

[Rig39] John R. Riggleman. Book review: [untitled]. Journal of the American Statistical Association, 34(207):602–603, September 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279504.

Riley:1937:IGP

[Ril37] Donald C. Riley. Improvement of government personnel. Journal of the American Statistical Association, 32(197):133–134, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279873.

Riley:1938:BRB

[Ril38] C. W. Riley. Book review: [untitled]. Journal of the American Statistical Association, 33(204):785–786, December 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279099.

Robinson:1933:BRB

[Rob33] Louis N. Robinson. Book review: [untitled]. Journal of the American Statistical Association, 28(182):241–242, June 1933. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278049.

Robinson:1934:CES

[Rob34] Leland Rex Robinson. Corporate earnings on share and borrowed capital in ratios of gross income. Journal of the American Statistical Association, 29(185):39–50, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278458.

Robinson:1936:CES

[Rob36] Leland Rex Robinson. Corporate earnings on share and borrowed capital in ratios of gross income (1918–1935). Journal of the American Statistical Association, 31(195):481–490, September 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278372.

Robinson:1938:BRB

[Rob38] Henry A. Robinson. Book review: [untitled]. Journal of the American Statistical Association, 33(201):280, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279139.

Robinson:1939:BRB

[Rob39] Leland Rex Robinson. Book review: [untitled]. Journal of the American Statistical Association, 34(207):605–606, September 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279507.

Rodkey:1932:BRB

[Rod32] R. G. Rodkey. Book review: [untitled]. Journal of the American Statistical Association, 27(178):233–234, June 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277685.

Roelse:1930:BRB

[Roe30] H. V. Roelse. Book review: [untitled]. Journal of the American Statistical Association, 25(169):126–127, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278521.

Roelse:1938:BRB

[Roe38] Harold V. Roelse. Book review: [untitled]. Journal of the American Statistical Association, 33(203):609–612, September 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279712.

Rogers:1932:GIC

[Rog32] James Harvey Rogers. Gold, international credits and depression. Journal of the American Statistical Association, 27(179):237–250, September 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278157.

Rogers:1934:BRB

[Rog34] H. O. Rogers. Book review: [untitled]. Journal of the American Statistical Association, 29(185):107–108, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278476.

Rolph:1930:BRBa

[Rol30a] John G. Rolph. Book review: [untitled]. Journal of the American Statistical Association, 25(169):120–121, March 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278516.

Rolph:1930:BRBb

[Rol30b] John G. Rolph. Book review: [untitled]. Journal of the American Statistical Association, 25(171):381, September 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278214.

Rolph:1931:BRB

[Rol31] J. G. Rolph. Book review: [untitled]. Journal of the American Statistical Association, 26(173):113–114, March 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278280.

Romanovsky:1935:AMS

[Rom35] V. Romanovsky. On the application of mathematical statistics and the theory of probability in the industries of the Soviet union. Journal of the American Statistical Association, 30(192):709–710, December 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277727.

Roos:1936:BRBa

[Roo36a] Charles F. Roos. Book review: [untitled]. Journal of the American Statistical Association, 31(194):420–421, June 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278578.

Roos:1936:BRBb

[Roo36b] Charles F. Roos. Book review: [untitled]. Journal of the American Statistical Association, 31(195):627–629, September 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278403.

Rorty:1930:PAM

[Ror30] M. C. Rorty. The program for the 1930 annual meeting. Journal of the American Statistical Association, 25(171):328–332, September 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278197.

Rorty:1931:SSM

[Ror31] Malcolm C. Rorty. Statistics and the scientific method. Journal of the American Statistical Association, 26(173):1–10, March 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278253.

Rorty:1935:NCC

[Ror35] M. C. Rorty. Notes on coal consumption. Journal of the American Statistical Association, 30(192):718, December 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277732.

Rorty:1936:UIU

[Ror36] M. C. Rorty. Unemployment insurance and unemployment. Journal of the American Statistical Association, 31(193):61–66, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278325.

Ross:1934:RSP

[Ros34] Emerson Ross. Research and statistical program of the Federal Emergency Relief Administration. Journal of the American Statistical Association, 29(187):288–294, September 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278065.

Ross:1938:CLF

[Ros38] Frank A. Ross. C. Luther Fry. Journal of the American Statistical Association, 33(203):605, September 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279707.

Royer:1933:SGC

[RT33] Elmer B. Royer and Herbert A. Toops. The statistics of geometrically coded scores. Journal of the American Statistical Association, 28(182):192–198, June 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278033.

Rubinow:1930:BRB

[Rub30] I. M. Rubinow. Book review: [untitled]. Journal of the American Statistical Association, 25(169):123–124, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278518.

Rubinow:1933:BRB

[Rub33] I. M. Rubinow. Book review: [untitled]. Journal of the American Statistical Association, 28(182):245–247, June 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278051.

Roos:1939:CDPa

[RvS39a] Charles F. Roos and Victor von Szeliski. The concept of demand and price elasticity — the dynamics of automobile demand. Journal of the American Statistical Association, 34(208):652–664, December 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279808.

Roos:1939:CDPb

[RvS39b] Charles F. Roos and Victor von Szeliski. The concept of demand and price elasticity — the dynamics of automobile demand: a further reply. Journal of the American Statistical Association, 34(208):665–666, December 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279810.

Sanger:1931:BRB

[San31] Homer F. Sanger. Book review: [untitled]. Journal of the American Statistical Association, 26(175):373, September 1931. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277931.

Sanders:1934:ICM

[San34] Barkev Sanders. The inadequacy of correlation method in genetic research. Journal of the American Statistical Association, 29(185): 59–65, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277804.

Sanders:1936:PPA

[San36a] Barkev S. Sanders. The purpose and progress of the attorney General’s survey of release procedures. Journal of the American Statistical Association, 31(196):732–734, December 1936. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278674.

Sanders:1936:BRB

[San36b] T. H. Sanders. Book review: [untitled]. Journal of the American Statistical Association, 31(196):771–772, December 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278692.

Sanders:1937:CRS

[San37] Barkev S. Sanders. Consistency of recording statistical data from prison files. Journal of the American Statistical Association, 32 (198):323–336, June 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279573.

Sanderson:1939:BRB

[San39] Fred H. Sanderson. Book review: [untitled]. Journal of the American Statistical Association, 34(207):627–631, September 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279527.

Sarle:1934:SPM

[Sar34] Charles F. Sarle. Statistical problems of the meat packing industry under an agricultural adjustment administration trade agreement. Journal of the American Statistical Association, 29(185):146–151, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277821.

Sasuly:1930:SSE

[Sas30] Max Sasuly. The simplest symmetrical expression for rate of change of time series data. Journal of the American Statistical Association, 25(169):72–75, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278497.

Schultz:1930:SEF

[Sch30] Henry Schultz. The standard error of a forecast from a curve. Journal of the American Statistical Association, 25(170):139–185, June 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277630.

Schultz:1931:BRB

[Sch31] Henry Schultz. Book review: [untitled]. Journal of the American Statistical Association, 26(176):484–491, December 1931. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277956.

Schultz:1933:AHC

[Sch33a] Arch D. Schultz. The allocation of highway costs among classes of motor vehicles for purposes of calculating graduated tax schedules. Journal of the American Statistical Association, 28(183):286–292, September 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278424.

Schultz:1933:SEC

[Sch33b] Henry Schultz. The standard error of the coefficient of elasticity of demand. Journal of the American Statistical Association, 28(181): 64–69, March 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278002.

Schneider:1935:CFC

[Sch35] David M. Schneider. Checking the 1933 Federal Census of dependent and neglected children in New York State. Journal of the American Statistical Association, 30(190):427–436, June 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277693.

Schultz:1936:BRB

[Sch36a] Henry Schultz. Book review: [untitled]. Journal of the American Statistical Association, 31(195):613–617, September 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278396.

Schumpeter:1936:BRB

[Sch36b] Joseph A. Schumpeter. Book review: The General Theory of Employment, Interest and Money, by John Maynard Keynes. Journal of the American Statistical Association, 31(196):791–795, December 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278703.

Schwenger:1936:BRB

[Sch36c] Robert B. Schwenger. Book review: [untitled]. Journal of the American Statistical Association, 31(195):637–638, September 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278408.

Schmalz:1938:BRB

[Sch38a] Carl N. Schmalz. Book review: [untitled]. Journal of the American Statistical Association, 33(203):639–642, September 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279726.

Schmid:1938:BRB

[Sch38b] Calvin F. Schmid. Book review: [untitled]. Journal of the American Statistical Association, 33(202):494–498, June 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278948.

Schenker:1939:BRBa

[Sch39a] David Schenker. Book review: [untitled]. Journal of the American Statistical Association, 34(208):778, December 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279842.

Schenker:1939:BRBb

[Sch39b] David Schenker. Book review: [untitled]. Journal of the American Statistical Association, 34(208):779, December 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279843.

Schmidt:1939:BRB

[Sch39c] Carl T. Schmidt. Book review: [untitled]. Journal of the American Statistical Association, 34(205):191–192, March 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279178.

Schneider:1939:PRL

[Sch39d] David M. Schneider. The place of research in the long-time planning of a state program for public welfare. Journal of the American Statistical Association, 34(208):674–682, December 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279812.

Schumacher:1939:BRB

[Sch39e] F. X. Schumacher. Book review: [untitled]. Journal of the American Statistical Association, 34(208):753–754, December 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279822.

Schwenger:1939:BRB

[Sch39f] Robert B. Schwenger. Book review: [untitled]. Journal of the American Statistical Association, 34(205):188–189, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279176.

Scott:1937:BRB

[Sco37] Frances V. Scott. Book review: [untitled]. Journal of the American Statistical Association, 32(200):811, December 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279410.

Scoville:1938:BRB

[Sco38] Warren C. Scoville. Book review: [untitled]. Journal of the American Statistical Association, 33(202):472–475, June 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278936.

Scott:1939:BRB

[Sco39] Frances V. Scott. Book review: [untitled]. Journal of the American Statistical Association, 34(208):756–757, December 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279824.

Sealy:1936:BRB

[Sea36] Marie P. Sealy. Book review: [untitled]. Journal of the American Statistical Association, 31(193):240–241, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278366.

Seltzer:1932:BRB

[Sel32] Lawrence H. Seltzer. Book review: [untitled]. Journal of the American Statistical Association, 27(180):452–453, December 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278654.

Seltzer:1933:BRB

[Sel33] Lawrence H. Seltzer. Book review: [untitled]. Journal of the American Statistical Association, 28(181):119–121, March 1933. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278025.

Seltzer:1934:SPA

[Sel34] Lawrence H. Seltzer. Statistical problems arising out of the NRA codes: The automobile industry. Journal of the American Statistical Association, 29(185):152–156, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277822.

Sellin:1938:BRB

[Sel38] Thorsten Sellin. Book review: [untitled]. Journal of the American Statistical Association, 33(204):755–756, December 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279079.

Shewhart:1931:SME

[She31] W. A. Shewhart. Statistical method from an engineering viewpoint. Journal of the American Statistical Association, 26(175):262–269, September 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277904.

Shewhart:1939:BRB

[She39] Walter A. Shewhart. Book review: [untitled]. Journal of the American Statistical Association, 34(207):589–591, September 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279494.

Shields:1937:BRBa

[Shi37a] Murray Shields. Book review: [untitled]. Journal of the American Statistical Association, 32(198):421–422, June 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279595.

Shields:1937:BRBb

[Shi37b] Murray Shields. Book review: [untitled]. Journal of the American Statistical Association, 32(200):773–774, December 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279389.

Shields:1938:BRB

[Shi38] Murray Shields. Book review: [untitled]. Journal of the American Statistical Association, 33(203):637, September 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279724.

Shlakman:1939:BRB

[Shl39] Vera Shlakman. Book review: [untitled]. Journal of the American Statistical Association, 34(205):232–233, March 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279209.

Shoup:1933:BRB

[Sho33] Carl Shoup. Book review: [untitled]. Journal of the American Statistical Association, 28(184):474–475, December 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278149.

Shoup:1935:BRB

[Sho35] Carl Shoup. Book review: [untitled]. Journal of the American Statistical Association, 30(189):130–131, March 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278546.

Shoup:1936:BRB

[Sho36] Carl Shoup. Book review: [untitled]. Journal of the American Statistical Association, 31(194):453–454, June 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278603.

Short:1937:CDC

[Sho37] Oliver C. Short. Consolidation of the Division of Current Business Statistics and the Division of Manufactures, Bureau of the Census. Journal of the American Statistical Association, 32(200):717–718, December 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279376.

Secrist:1934:OL

[SHR^{+}34] Horace Secrist, Harold Hotelling, M. C. Rorty, Corrado Gini, and Willford I. King. Open letters. Journal of the American Statistical Association, 29(186):196–205, June 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278295.

Shryock:1939:BRB

[Shr39] Henry S. Shryock, Jr. Book review: [untitled]. Journal of the American Statistical Association, 34(208):779–780, December 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279844.

Sibley:1930:AGP

[Sib30] Elbridge Sibley. An apparatus for graphic portrayal of series of products or quotients without computation. Journal of the American Statistical Association, 25(172):460–463, December 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277972.

Sienkiewicz:1939:BRB

[Sie39] C. A. Sienkiewicz. Book review: [untitled]. Journal of the American Statistical Association, 34(208):765, December 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279831.

Simmons:1937:BRB

[Sim37] Carroll D. Simmons. Book review: [untitled]. Journal of the American Statistical Association, 32(199):610–611, September 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278796.

Simmons:1938:BRB

[Sim38] Carroll D. Simmons. Book review: [untitled]. Journal of the American Statistical Association, 33(204):781, December 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279096.

Sissman:1936:DPS

[Sis36] Louise Sissman. Development of the postal savings system. Journal of the American Statistical Association, 31(196):708–718, December 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2278668.

Sloan:1930:BRB

[Slo30] Laurence H. Sloan. Book review: [untitled]. Journal of the American Statistical Association, 25(172):487–488, December 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277982.

Sloan:1931:BRB

[Slo31] Laurence H. Sloan. Book review: [untitled]. Journal of the American Statistical Association, 26(174):239–240, June 1931. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277790.

Sloan:1937:CSB

[Slo37] Laurence H. Sloan. Corporation statements as the basis of investment policy. Journal of the American Statistical Association, 32(197):8–14, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279853.

Smart:1933:BCB

[Sma33] O. M. Smart. Business curve based on interest rates and prices: An adaptation of the quantity theory of money. Journal of the American Statistical Association, 28(181):76–81, March 1933. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278004.

Smart:1934:BRB

[Sma34] L. Edwin Smart. Book review: [untitled]. Journal of the American Statistical Association, 29(188):455–456, December 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278635.

Smith:1931:FIB

[Smi31] Bradford B. Smith. A forecasting index for business. Journal of the American Statistical Association, 26(174):115–127, June 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277760.

Smith:1932:BRB

[Smi32] James G. Smith. Book review: [untitled]. Journal of the American Statistical Association, 27(178):223–225, June 1932. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277680.

Smith:1933:BRB

[Smi33] James Gerald Smith. Book review: [untitled]. Journal of the American Statistical Association, 28(182):248–249, June 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278052.

Smith:1934:EEF

[Smi34] E. Dillon Smith. The exponential equation fitted by the mean value method. Journal of the American Statistical Association, 29(186):184–187, June 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278290.

Smith:1935:BRBa

[Smi35a] James G. Smith. Book review: [untitled]. Journal of the American Statistical Association, 30(189):135–136, March 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278550.

Smith:1935:BRBb

[Smi35b] James G. Smith. Book review: [untitled]. Journal of the American Statistical Association, 30(191):639–642, September 1935. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278121.

Smith:1935:TAG

[Smi35c] Mapheus Smith. Trends in the ages of gainful workers, by occupation, 1910–1930. Journal of the American Statistical Association, 30(192):678–687, December 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277721.

Smith:1936:BRB

[Smi36] James G. Smith. Book review: [untitled]. Journal of the American Statistical Association, 31(194):427–429, June 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278584.

Smith:1937:BRBb

[Smi37a] T. Lynn Smith. Book review: [untitled]. Journal of the American Statistical Association, 32(199):594–596, September 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278788.

Smith:1937:BRBa

[Smi37b] William E. Smith. Book review: [untitled]. Journal of the American Statistical Association, 32(198):434–435, June 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279603.

Smith:1939:BRB

[Smi39] James G. Smith. Book review: [untitled]. Journal of the American Statistical Association, 34(208):763–764, December 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279830.

Sydenstricker:1930:DFA

[SN30] Edgar Sydenstricker and Frank W. Notestein. Differential fertility according to social class: a study of 69,620 native white married women under 45 years of age based upon the United States Census returns of 1910. Journal of the American Statistical Association, 25(169):9–32, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278490.

Snedecor:1934:MEN

[Sne34] George W. Snedecor. The method of expected numbers for tables of multiple classification with disproportionate subclass numbers. Journal of the American Statistical Association, 29(188):389–393, December 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278609.

Snedecor:1936:IST

[Sne36] George W. Snedecor. The improvement of statistical techniques in biology. Journal of the American Statistical Association, 31 (196):690–701, December 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278666.

Snedecor:1939:BRB

[Sne39] George W. Snedecor. Book review: [untitled]. Journal of the American Statistical Association, 34(205):206–207, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279190.

Snyder:1934:ILT

[Sny34] Carl Snyder. The increase of long-term debt in the United States (from 1880). Journal of the American Statistical Association, 29 (186):166–174, June 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278286.

Sogge:1933:ICU

[Sog33] Tillman M. Sogge. Industrial classes in the United States in 1930. Journal of the American Statistical Association, 28(182):199–203, June 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278034.

Soule:1930:BRB

[Sou30] George Soule. Book review: [untitled]. Journal of the American Statistical Association, 25(171):375–377, September 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278211.

Soule:1934:BRB

[Sou34] George Soule. Book review: [untitled]. Journal of the American Statistical Association, 29(187):346–347, September 1934. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278089.

Soule:1936:BRB

[Sou36] George Soule. Book review: [untitled]. Journal of the American Statistical Association, 31(196):787–788, December 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278700.

Snyder:1931:IVT

[SP31] Carl Snyder and Leroy M. Piser. The index of the volume of trade: Third revision. Journal of the American Statistical Association, 26 (176):436–442, December 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277944.

Sydenstricker:1934:CFI

[SP34] Edgar Sydenstricker and G. St. J. Perrott. Changes in family income and rental during the economic depression. Journal of the American Statistical Association, 29(185):43–46, March 1934. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277801.

Schoenberg:1937:MPS

[SP37] Erika H. Schoenberg and Mildred Parten. Methods and problems of sampling presented by the urban study of consumer purchases. Journal of the American Statistical Association, 32(198):311–322, June 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279572.

Spahr:1935:BRB

[Spa35] Walter E. Spahr. Book review: [untitled]. Journal of the American Statistical Association, 30(189):131–132, March 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278547.

Spearman:1930:BRB

[Spe30] C. Spearman. Book review: [untitled]. Journal of the American Statistical Association, 25(169):107–110, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278509.

Spengler:1931:BRB

[Spe31] Joseph J. Spengler. Book review: [untitled]. Journal of the American Statistical Association, 26(173):107–108, March 1931. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278275.

Spengler:1939:BRB

[Spe39] J. J. Spengler. Book review: [untitled]. Journal of the American Statistical Association, 34(205):211–212, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279195.

Spier:1934:AVS

[Spi34] J. Spier. The average value of the sum of absolute differences of paired ordinals of two ordinal series. Journal of the American Statistical Association, 29(186):180–181, June 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278288.

Spiegelman:1939:BRB

[Spi39] Mortimer Spiegelman. Book review: [untitled]. Journal of the American Statistical Association, 34(206):442–445, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278883.

Sprague:1934:SE

[Spr34] O. M. W. Sprague. Statistics and economics. Journal of the American Statistical Association, 29(185):162–164, March 1934. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277824.

Spurr:1937:GMM

[Spu37] William A. Spurr. A graphic method of measuring seasonal variation. Journal of the American Statistical Association, 32(198):281– 289, June 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279570.

Schultz:1933:AVE

[SS33] Theodore W. Schultz and George W. Snedecor. Analysis of variance as an effective method of handling the time element in certain economic statistics. Journal of the American Statistical Association, 28(181):14–30, March 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277999.

Stephan:1931:ASP

[Ste31] Frederick F. Stephan. Alternative statements of percentage data in the fitting of logarithmic curves. Journal of the American Statistical Association, 26(173):58–61, March 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278260.

Stephan:1932:SMF

[Ste32] Frederick F. Stephan. Summation methods in fitting parabolic curves. Journal of the American Statistical Association, 27(180):413–423, December 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278640.

Stecker:1934:BRB

[Ste34a] Margaret L. Stecker. Book review: [untitled]. Journal of the American Statistical Association, 29(185):114–116, March 1934. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278481.

Stephan:1934:BRB

[Ste34b] Frederick F. Stephan. Book review: [untitled]. Journal of the American Statistical Association, 29(188):449–450, December 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278631.

Stephan:1934:SEI

[Ste34c] Frederick F. Stephan. Sampling errors and interpretations of social data ordered in time and space. Journal of the American Statistical Association, 29(185):165–166, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277825.

Stevens:1934:OGP

[Ste34d] W. S. Stevens. Obituary: George Pendleton Watkins. Journal of the American Statistical Association, 29(185):99–100, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278471.

Stephan:1935:DAA

[Ste35a] Frederick F. Stephan. Development of the activities of the American statistical association. Journal of the American Statistical Association, 30(189):102–103, March 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278535.

Stephan:1935:EH

[Ste35b] Frederick F. Stephan. Employment and health. Journal of the American Statistical Association, 30(189):100–102, March 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278533.

Stephan:1935:NC

[Ste35c] Frederick F. Stephan. Nominating committee. Journal of the American Statistical Association, 30(191):600, September 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278109.

Stewart:1935:BRB

[Ste35d] Bryce M. Stewart. Book review: [untitled]. Journal of the American Statistical Association, 30(190):494–495, June 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277712.

Stephan:1936:NC

[Ste36a] Frederick F. Stephan. Nominating committee. Journal of the American Statistical Association, 31(194):390, June 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278570.

Stewart:1936:BRBa

[Ste36b] Bryce M. Stewart. Book review: [untitled]. Journal of the American Statistical Association, 31(195):646–648, September 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278413.

Stewart:1936:BRBb

[Ste36c] Paul W. Stewart. Book review: [untitled]. Journal of the American Statistical Association, 31(196):788–789, December 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278701.

Stephan:1937:NC

[Ste37a] Frederick F. Stephan. Nominating committee. Journal of the American Statistical Association, 32(199):579, September 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278779.

Stewart:1937:BRB

[Ste37b] Bryce M. Stewart. Book review: [untitled]. Journal of the American Statistical Association, 32(197):229–232, March 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279887.

Stephan:1938:NC

[Ste38a] Frederick F. Stephan. Nominating committee. Journal of the American Statistical Association, 33(203):605, September 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279708.

Stephan:1938:SMA

[Ste38b] Frederick F. Stephan. Special meeting of the American statistical association. Journal of the American Statistical Association, 33(203):571–572, September 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279696.

Stephan:1939:NC

[Ste39a] Frederick F. Stephan. Nominating committee. Journal of the American Statistical Association, 34(207):587, September 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279492.

Stephan:1939:RSL

[Ste39b] Frederick F. Stephan. Representative sampling in large-scale surveys. Journal of the American Statistical Association, 34(206):343– 352, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278856.

Stewart:1939:BRB

[Ste39c] Blair Stewart. Book review: [untitled]. Journal of the American Statistical Association, 34(206):459–460, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278897.

Stiebeling:1934:SAF

[Sti34] Hazel K. Stiebeling. A statistical analysis of the food requirements for the population of the United States. Journal of the American Statistical Association, 29(185):186–192, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277832.

Stigler:1939:LSD

[Sti39] George J. Stigler. The limitations of statistical demand curves. Journal of the American Statistical Association, 34(207):469–481, September 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279481.

Stocking:1931:BRB

[Sto31] George Ward Stocking. Book review: [untitled]. Journal of the American Statistical Association, 26(174):236–237, June 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277787.

Stockder:1932:BRB

[Sto32] A. H. Stockder. Book review: [untitled]. Journal of the American Statistical Association, 27(178):225–229, June 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277681.

Stouffer:1933:TSA

[Sto33] Samuel A. Stouffer. Tests of significance in applying Westergaard’s method of expected cases to sociological data. Journal of the American Statistical Association, 28(183):293–302, September 1933. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278425.

Stouffer:1934:CCP

[Sto34a] Samuel A. Stouffer. A coefficient of “combined partial correlation” with an example from sociological data. Journal of the American Statistical Association, 29(185):70–71, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278461.

Stouffer:1934:FFR

[Sto34b] Samuel A. Stouffer. Fertility of families on relief. Journal of the American Statistical Association, 29(187):295–300, September 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278066.

Stouffer:1934:PAC

[Sto34c] Samuel A. Stouffer. Problems in the application of correlation to sociology. Journal of the American Statistical Association, 29(185): 52–58, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277803.

Stouffer:1936:EEI

[Sto36] Samuel A. Stouffer. Evaluating the effect of inadequately measured variables in partial correlation analysis. Journal of the American Statistical Association, 31(194):348–360, June 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278562.

Stocking:1937:BRB

[Sto37] Collis Stocking. Book review: [untitled]. Journal of the American Statistical Association, 32(200):784–785, December 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279394.

Stocking:1938:BRB

[Sto38] S. B. Stocking. Book review: [untitled]. Journal of the American Statistical Association, 33(202):488–489, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278944.

Stock:1939:MGI

[Sto39a] J. Stevens Stock. A method of graphic interpolation. Journal of the American Statistical Association, 34(208):709–713, December 1939.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279815.

Stocking:1939:BRB

[Sto39b] George Ward Stocking. Book review: [untitled]. Journal of the American Statistical Association, 34(207):611–612, September 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279512.

Stratton:1937:BRB

[Str37] Samuel S. Stratton. Book review: [untitled]. Journal of the American Statistical Association, 32(200):781–783, December 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279393.

Strauss:1939:BRB

[Str39] Frederick Strauss. Book review: [untitled]. Journal of the American Statistical Association, 34(206):425–426, June 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278870.

Sturges:1934:NCR

[Stu34] Alexander Sturges. Note on correlation in random series. Journal of the American Statistical Association, 29(188):421–422, December 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2278616.

Studenski:1937:BRB

[Stu37a] Paul Studenski. Book review: [untitled]. Journal of the American Statistical Association, 32(198):436–438, June 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279605.

Sturm:1937:HSE

[Stu37b] Wm. A. Sturm. Handbook on sources of economic data pertaining to California. Journal of the American Statistical Association, 32(198):375–376, June 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279585.

Subramanian:1934:CCF

[Sub34] S. Subramanian. On a certain conclusion of Frisch’s. Journal of the American Statistical Association, 29(187):316–317, September 1934.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278072.

Swart:1936:BRB

[Swa36] Jacob Swart. Book review: [untitled]. Journal of the American Statistical Association, 31(193):239–240, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278365.

Swayzee:1939:BRB

[Swa39] Cleon O. Swayzee. Book review: [untitled]. Journal of the American Statistical Association, 34(205):216–217, March 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279199.

Sweeney:1935:BRB

[Swe35] Henry W. Sweeney. Book review: [untitled]. Journal of the American Statistical Association, 30(190):489–491, June 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277709.

Swineford:1938:BRB

[Swi38] Frances Swineford. Book review: [untitled]. Journal of the American Statistical Association, 33(203):621–622, September 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279716.

T:1930:PN

[T.30] A. E. T. Personal notes. Journal of the American Statistical Association, 25(171):356–360, September 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278204.

T:1934:IPD

[T.34] L. E. T. Index of population data. Journal of the American Statistical Association, 29(185):86, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278467.

Taeuber:1938:BRB

[Tae38] Conrad Taeuber. Book review: [untitled]. Journal of the American Statistical Association, 33(201):293–294, March 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279148.

Taylor:1930:BRB

[Tay30a] Maurice Taylor. Book review: [untitled]. Journal of the American Statistical Association, 25(172):489–491, December 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277984.

Taylor:1930:EMC

[Tay30b] Paul S. Taylor. Employment of Mexicans in Chicago and the Calumet region. Journal of the American Statistical Association, 25(170):206– 207, June 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277635.

Taylor:1931:BRB

[Tay31] Maurice Taylor. Book review: [untitled]. Journal of the American Statistical Association, 26(174):237–238, June 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277788.

Taylor:1932:NAT

[Tay32] Paul S. Taylor. Note on age at time of marriage in two Mexican localities of divergent types. Journal of the American Statistical Association, 27(180):427–430, December 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278642.

Taylor:1934:CCS

[Tay34] James S. Taylor. Construction cost statistics. Journal of the American Statistical Association, 29(185):38–42, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277800.

Taylor:1935:BRB

[Tay35] Maurice Taylor. Book review: [untitled]. Journal of the American Statistical Association, 30(189):134–135, March 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278549.

Taylor:1937:SMB

[Tay37] Amos E. Taylor. Statistical methods in balance-of-payments estimates. Journal of the American Statistical Association, 32(197):65– 74, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279863.

Taylor:1938:BRB

[Tay38] K. W. Taylor. Book review: [untitled]. Journal of the American Statistical Association, 33(201):300–302, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279153.

Tucker:1938:BBG

[TBSM38] Rufus S. Tucker, Alfred L. Bernheim, Margaret Grant Schneider, and Gardiner C. Means. Big business, its growth and its place. Journal of the American Statistical Association, 33(202):406–411, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278919.

Teagarden:1938:BRBa

[Tea38a] Florence M. Teagarden. Book review: [untitled]. Journal of the American Statistical Association, 33(204):789, December 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279102.

Teagarden:1938:BRBb

[Tea38b] Florence M. Teagarden. Book review: [untitled]. Journal of the American Statistical Association, 33(204):789–791, December 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279103.

Teagarden:1938:BRBc

[Tea38c] Florence M. Teagarden. Book review: [untitled]. Journal of the American Statistical Association, 33(204):791, December 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279104.

Teele:1937:BRBa

[Tee37a] Stanley F. Teele. Book review: [untitled]. Journal of the American Statistical Association, 32(200):776–777, December 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279391.

Teele:1937:BRBb

[Tee37b] Stanley F. Teele. Book review: [untitled]. Journal of the American Statistical Association, 32(200):785–787, December 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279395.

Thorp:1930:BRB

[Tho30] Willard L. Thorp. Book review: [untitled]. Journal of the American Statistical Association, 25(172):502–503, December 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277994.

Thomas:1932:BRBa

[Tho32a] Woodlief Thomas. Book review: [untitled]. Journal of the American Statistical Association, 27(177):112–115, March 1932. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277897.

Thomas:1932:BRBb

[Tho32b] Woodlief Thomas. Book review: [untitled]. Journal of the American Statistical Association, 27(178):230–232, June 1932. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277683.

Thompson:1932:BRB

[Tho32c] Warren S. Thompson. Book review: [untitled]. Journal of the American Statistical Association, 27(177):102–103, March 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277891.

Thorne:1932:BRB

[Tho32d] G. B. Thorne. Book review: [untitled]. Journal of the American Statistical Association, 27(180):459–461, December 1932. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278660.

Thorp:1932:BRB

[Tho32e] Willard L. Thorp. Book review: [untitled]. Journal of the American Statistical Association, 27(177):111–112, March 1932. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277896.

Thomas:1933:EDU

[Tho33] Carl Thomas. The effect of the depression upon bond yields. Journal of the American Statistical Association, 28(183):261–271, September 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2278422.

Thorp:1934:BRB

[Tho34a] Willard L. Thorp. Book review: [untitled]. Journal of the American Statistical Association, 29(185):121–122, March 1934. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278485.

Thorp:1934:SPW

[Tho34b] Willard L. Thorp. Statistics as a public work. Journal of the American Statistical Association, 29(185):157–161, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277823.

Thomas:1935:BRB

[Tho35a] Woodlief Thomas. Book review: [untitled]. Journal of the American Statistical Association, 30(189):124–129, March 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278544.

Thompson:1935:BRB

[Tho35b] Warren S. Thompson. Book review: [untitled]. Journal of the American Statistical Association, 30(190):491–493, June 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277710.

Thomas:1936:SC

[Tho36a] Dorothy Swaine Thomas. The Swedish census of 1935–6. Journal of the American Statistical Association, 31(195):541–544, September 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2278380.

Thorp:1936:WIS

[Tho36b] Willard L. Thorp. Wanted — industrial statistics. Journal of the American Statistical Association, 31(193):47–52, March 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278322.

Thomas:1937:AIM

[Tho37a] Dorothy Swaine Thomas. The analysis of internal migration in the Swedish Census of 1930. Journal of the American Statistical Association, 32(197):124–130, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279871.

Thorndike:1937:VAC

[Tho37b] Edward L. Thorndike. Variations among cities in per capita income. Journal of the American Statistical Association, 32(199):471–479, September 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278767.

Thompson:1938:TBE

[Tho38a] Donald S. Thompson. Trends of bank earnings and expenses. Journal of the American Statistical Association, 33(202):327–333, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278907.

Thorp:1938:BRB

[Tho38b] Willard L. Thorp. Book review: [untitled]. Journal of the American Statistical Association, 33(201):269–271, March 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279134.

Thornthwaite:1939:BRB

[Tho39a] C. W. Thornthwaite. Book review: [untitled]. Journal of the American Statistical Association, 34(206):446–447, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278885.

Thorp:1939:BRBa

[Tho39b] Willard L. Thorp. Book review: [untitled]. Journal of the American Statistical Association, 34(206):451–452, June 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278890.

Thorp:1939:BRBb

[Tho39c] Willard L. Thorp. Book review: [untitled]. Journal of the American Statistical Association, 34(208):766–767, December 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279832.

Thresher:1939:BRB

[Thr39] B. A. Thresher. Book review: [untitled]. Journal of the American Statistical Association, 34(205):229–230, March 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279206.

Timoshenko:1930:BRB

[Tim30] V. P. Timoshenko. Book review: [untitled]. Journal of the American Statistical Association, 25(171):373–375, September 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278210.

Timoshenko:1931:BRB

[Tim31] V. P. Timoshenko. Book review: [untitled]. Journal of the American Statistical Association, 26(175):372–373, September 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277930.

Timoshenko:1932:BRBa

[Tim32a] V. P. Timoshenko. Book review: [untitled]. Journal of the American Statistical Association, 27(179):352–353, September 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278185.

Timoshenko:1932:BRBb

[Tim32b] V. P. Timoshenko. Book review: [untitled]. Journal of the American Statistical Association, 27(180):461–462, December 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278661.

Timoshenko:1933:BRB

[Tim33] V. P. Timoshenko. Book review: [untitled]. Journal of the American Statistical Association, 28(181):114–115, March 1933. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278020.

Timoshenko:1935:BRB

[Tim35] V. P. Timoshenko. Book review: [untitled]. Journal of the American Statistical Association, 30(191):644–646, September 1935. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278124.

Timoshenko:1936:BRB

[Tim36] V. P. Timoshenko. Book review: [untitled]. Journal of the American Statistical Association, 31(194):442–443, June 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278594.

Timoshenko:1937:BRB

[Tim37] V. P. Timoshenko. Book review: [untitled]. Journal of the American Statistical Association, 32(198):425–428, June 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279597.

Tippetts:1936:BRB

[Tip36] Charles S. Tippetts. Book review: [untitled]. Journal of the American Statistical Association, 31(195):635–637, September 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278407.

Tippetts:1937:BRB

[Tip37] Charles S. Tippetts. Book review: [untitled]. Journal of the American Statistical Association, 32(199):605–609, September 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278794.

Tippetts:1939:BRB

[Tip39a] Charles S. Tippetts. Book review: [untitled]. Journal of the American Statistical Association, 34(206):454–455, June 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278892.

Tippetts:1939:PPD

[Tip39b] Charles S. Tippetts. Postponable purchasing of durable consumers’ goods and the business cycle. Journal of the American Statistical Association, 34(205):27–31, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279158.

Titus:1936:KIE

[Tit36] Walter F. Titus. The kind of information an executive needs to operate a factory. Journal of the American Statistical Association, 31(193):43–46, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278321.

Tough:1933:SMS

[TK33] Evelyn G. Tough and E. L. Kirkpatrick. Scales for measuring the standard of living. Journal of the American Statistical Association, 28(181):55–63, March 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278001.

Tolles:1934:URI

[Tol34] Arnold Tolles. Unemployment rates, insurance benefit, and contribution scales in Germany and Great Britain. Journal of the American Statistical Association, 29(186):159–165, June 1934. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278285.

Toops:1935:BRB

[Too35] Herbert A. Toops. Book review: [untitled]. Journal of the American Statistical Association, 30(190):483–484, June 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277705.

Toops:1936:BRB

[Too36] Herbert A. Toops. Book review: [untitled]. Journal of the American Statistical Association, 31(195):624–625, September 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278401.

Toops:1937:BRB

[Too37] Herbert A. Toops. Book review: [untitled]. Journal of the American Statistical Association, 32(200):812–813, December 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279411.

Treloar:1938:BRB

[Tre38] A. E. Treloar. Book review: [untitled]. Journal of the American Statistical Association, 33(201):271–273, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279135.

Troxel:1939:BRB

[Tro39] C. Emery Troxel. Book review: [untitled]. Journal of the American Statistical Association, 34(207):607–608, September 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279509.

Truesdell:1931:TFD

[Tru31] Leon E. Truesdell. Tabulation of family data from the 1930 census. Journal of the American Statistical Association, 26(175):325–330, September 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277910.

Truesdell:1933:FCR

[Tru33] Leon E. Truesdell. Fifteenth census reports on population. Journal of the American Statistical Association, 28(183):334–336, September 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2278430.

Truesdell:1935:MTP

[Tru35] Leon E. Truesdell. The mechanics of the tabulation of the population census. Journal of the American Statistical Association, 30(189):89– 94, March 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278530.

Truesdell:1938:RRV

[Tru38] Leon E. Truesdell. Residual relationships and velocity of change as pitfalls in the field of statistical forecasting. Journal of the American Statistical Association, 33(202):373–379, June 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278913.

Truesdell:1939:BRB

[Tru39] Leon E. Truesdell. Book review: [untitled]. Journal of the American Statistical Association, 34(206):445–446, June 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278884.

Taeuber:1938:MCF

[TT38] Conrad Taeuber and Irene B. Taeuber. Measures of changes in fertility in Germany. Journal of the American Statistical Association, 33(204):709–712, December 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279062.

Tucker:1936:BRB

[Tuc36a] R. S. Tucker. Book review: [untitled]. Journal of the American Statistical Association, 31(194):430–432, June 1936. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278586.

Tucker:1936:RWA

[Tuc36b] Rufus S. Tucker. Real wages of artisans in London, 1729–1935. Journal of the American Statistical Association, 31(193):73–84, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2278327.

Tucker:1937:BRB

[Tuc37] Rufus S. Tucker. Book review: [untitled]. Journal of the American Statistical Association, 32(197):223–225, March 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279884.

Tufts:1937:BRB

[Tuf37] Edith Miller Tufts. Book review: [untitled]. Journal of the American Statistical Association, 32(197):244, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279895.

Tulin:1933:BRB

[Tul33] Justine Wise Tulin. Book review: [untitled]. Journal of the American Statistical Association, 28(182):249–251, June 1933. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278053.

Turner:1936:NIB

[Tur36] A. Willard Turner. Notes on an index of bond prices. Journal of the American Statistical Association, 31(195):555–558, September 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278385.

Turner:1939:TCB

[Tur39] A. Willard Turner. Trends in Canadian bond, stock and commodity markets 1928 to 1938. Journal of the American Statistical Association, 34(206):291–298, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278847.

Usher:1933:BRB

[Ush33] Abbott Payson Usher. Book review: [untitled]. Journal of the American Statistical Association, 28(184):480–481, December 1933. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278152.

Vaile:1936:BRB

[Vai36] Roland S. Vaile. Book review: [untitled]. Journal of the American Statistical Association, 31(196):815–819, December 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278716.

Valk:1930:IIS

[Val30] Willem L. Valk. The importance of an index of specific relative growth of industries. Journal of the American Statistical Association, 25(169):82–83, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278500.

VanBuren:1935:BRU

[Van35] George H. Van Buren. Book review: [untitled]. Journal of the American Statistical Association, 30(192):778–780, December 1935. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277754.

Vanderblue:1939:BRB

[Van39] Homer B. Vanderblue. Book review: [untitled]. Journal of the American Statistical Association, 34(207):614–615, September 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279516.

vonBeckerath:1939:BRB

[vB39] Herbert von Beckerath. Book review: [untitled]. Journal of the American Statistical Association, 34(206):456, June 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278894.

Veenstra:1930:MAT

[Vee30] Theodore A. Veenstra. A method of adjusting turnover rates. Journal of the American Statistical Association, 25(172):407–412, December 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277966.

Veenstra:1936:BRB

[Vee36] Theodore A. Veenstra. Book review: [untitled]. Journal of the American Statistical Association, 31(194):438–439, June 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278591.

Veenstra:1939:BRB

[Vee39] Theodore A. Veenstra. Book review: [untitled]. Journal of the American Statistical Association, 34(205):221–222, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279201.

vonHuhn:1930:BRBa

[vH30a] R. von Huhn. Book review: [untitled]. Journal of the American Statistical Association, 25(169):125, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278519.

vonHuhn:1930:BRBb

[vH30b] R. von Huhn. Book review: [untitled]. Journal of the American Statistical Association, 25(170):241–242, June 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277651.

vonHuhn:1930:RBA

[vH30c] R. von Huhn. The relation between the arithmetic and geometric average of two index numbers. Journal of the American Statistical Association, 25(169):76–79, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278498.

vonHuhn:1931:TMC

[vH31] R. von Huhn. A trigonometrical method for computing the scales of statistical charts to improve visualization. Journal of the American Statistical Association, 26(175):319–324, September 1931. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277909.

vonHuhn:1932:BRB

[vH32] R. von Huhn. Book review: [untitled]. Journal of the American Statistical Association, 27(179):348–350, September 1932. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278182.

vonHuhn:1933:RRT

[vH33a] R. von Huhn. The Richards–Roope tangent meter. Journal of the American Statistical Association, 28(181):82, March 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278005.

vonHuhn:1933:SCM

[vH33b] R. von Huhn. Secondary curves as a measure of the lag or phase difference between two primary curves. Journal of the American Statistical Association, 28(183):312–327, September 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278427.

vonHuhn:1933:SDM

[vH33c] R. von Huhn. The standard deviation as a measure of the intensity of seasonal adjustment. Journal of the American Statistical Association, 28(181):70–75, March 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278003.

vonHuhn:1934:BRBa

[vH34a] R. von Huhn. Book review: [untitled]. Journal of the American Statistical Association, 29(186):227–228, June 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278305.

vonHuhn:1934:BRBb

[vH34b] R. von Huhn. Book review: [untitled]. Journal of the American Statistical Association, 29(187):343–345, September 1934. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278087.

vonHuhn:1938:BRB

[vH38a] R. von Huhn. Book review: [untitled]. Journal of the American Statistical Association, 33(203):619–621, September 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279715.

vonHuhn:1938:MIS

[vH38b] R. von Huhn. A mechanical intensity shading map. Journal of the American Statistical Association, 33(204):713, December 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279063.

Vial:1937:BRB

[Via37] E. E. Vial. Book review: [untitled]. Journal of the American Statistical Association, 32(200):802–803, December 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279404.

Vickery:1938:PCT

[Vic38] C. W. Vickery. Punched card technique for the correction of bias in sampling. Journal of the American Statistical Association, 33 (203):552–556, September 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279693.

Vickery:1939:NPC

[Vic39] C. W. Vickery. Note on punched card technique. Journal of the American Statistical Association, 34(205):110–111, March 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279170.

Viner:1930:BRB

[Vin30] Jacob Viner. Book review: [untitled]. Journal of the American Statistical Association, 25(169):111–113, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278510.

Vold:1939:SPE

[Vol39] George B. Vold. The statistical presentation of the effectiveness of institutional treatment. Journal of the American Statistical Association, 34(205):79–86, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279165.

vonSzeliski:1930:CVD

[vS30] Victor S. von Szeliski. Comment on “the variate difference method of seasonal variation”. Journal of the American Statistical Association, 25(169):83–84, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278501.

vonSzeliski:1931:BRB

[vS31] Victor von Szeliski. Book review: [untitled]. Journal of the American Statistical Association, 26(174):235–236, June 1931. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277786.

Wadleigh:1934:BRB

[Wad34] H. J. Wadleigh. Book review: [untitled]. Journal of the American Statistical Association, 29(185):105–106, March 1934. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278475.

Waite:1932:FRG

[Wai32a] Warren C. Waite. Further remarks on the graphic method of correlation: II. rejoinder. Journal of the American Statistical Association, 27(178):185, June 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277663.

Waite:1932:SCG

[Wai32b] Warren C. Waite. Some characteristics of the graphic method of correlation. Journal of the American Statistical Association, 27(177): 68–70, March 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277882.

Waite:1936:BRB

[Wai36] Warren C. Waite. Book review: [untitled]. Journal of the American Statistical Association, 31(195):617–618, September 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278397.

Waite:1938:AAP

[Wai38] Warren C. Waite. The association of agricultural price fluctuations. Journal of the American Statistical Association, 33(202):404–406, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278918.

Waines:1939:BRB

[Wai39] W. J. Waines. Book review: [untitled]. Journal of the American Statistical Association, 34(206):461–462, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278899.

Walker:1934:BCN

[Wal34] Helen M. Walker. Bi-centenary of the normal curve. Journal of the American Statistical Association, 29(185):72–75, March 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278462.

Wallis:1936:PDS

[Wal36] W. Allen Wallis. The Poisson distribution and the supreme court. Journal of the American Statistical Association, 31(194):376–380, June 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278565.

Walradt:1937:BRBa

[Wal37a] Henry F. Walradt. Book review: [untitled]. Journal of the American Statistical Association, 32(200):827–828, December 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279419.

Walradt:1937:BRBb

[Wal37b] Henry F. Walradt. Book review: [untitled]. Journal of the American Statistical Association, 32(200):828, December 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279420.

Wallace:1938:BRB

[Wal38] Donald H. Wallace. Book review: [untitled]. Journal of the American Statistical Association, 33(204):766–768, December 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279087.

Wald:1939:BRB

[Wal39a] Abraham Wald. Book review: [untitled]. Journal of the American Statistical Association, 34(207):591–592, September 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279496.

Walker:1939:BRB

[Wal39b] Helen M. Walker. Book review: [untitled]. Journal of the American Statistical Association, 34(207):593–595, September 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279497.

Walker:1939:NIE

[Wal39c] Q. Forrest Walker. Notes on the interpretation of external statistics. Journal of the American Statistical Association, 34(206):277– 282, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278845.

Wallis:1939:CRR

[Wal39d] W. Allen Wallis. The correlation ratio for ranked data. Journal of the American Statistical Association, 34(207):533–538, September 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279486.

Warren:1930:BRB

[War30] Robert B. Warren. Book review: [untitled]. Journal of the American Statistical Association, 25(169):118–119, March 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278514.

Warburton:1934:VGN

[War34a] Clark Warburton. Value of the gross national product and its components, 1919–1929. Journal of the American Statistical Association, 29(188):383–388, December 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278608.

Warne:1934:BRB

[War34b] Colston E. Warne. Book review: [untitled]. Journal of the American Statistical Association, 29(188):440–441, December 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278626.

Warren:1936:BRB

[War36] G. F. Warren. Book review: [untitled]. Journal of the American Statistical Association, 31(196):803–805, December 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278708.

Warburton:1937:BRB

[War37] Clark Warburton. Book review: [untitled]. Journal of the American Statistical Association, 32(200):759–764, December 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279384.

Warburton:1938:BRB

[War38a] Clark Warburton. Book review: [untitled]. Journal of the American Statistical Association, 33(203):630–637, September 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279723.

Warburton:1938:DWR

[War38b] Clark Warburton. Dr. Warburton’s review of national income and capital formation — a rejoinder. Journal of the American Statistical Association, 33(204):719–721, December 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279066.

Watkins:1930:BRB

[Wat30a] Ralph J. Watkins. Book review: [untitled]. Journal of the American Statistical Association, 25(170):229–230, June 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277641.

Watkins:1930:UCN

[Wat30b] Ralph J. Watkins. The use of coefficients of net determination in testing the economic validity of correlation results. Journal of the American Statistical Association, 25(170):191–197, June 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277632.

Watts:1931:IPD

[Wat31] Reginald E. Watts. The influence of population density on crime. Journal of the American Statistical Association, 26(173):11–20, March 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278254.

Watkins:1933:OIC

[Wat33] G. P. Watkins. An ordinal index of correlation. Journal of the American Statistical Association, 28(182):139–151, June 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278028.

Watson:1936:BRB

[Wat36] J. P. Watson. Book review: [untitled]. Journal of the American Statistical Association, 31(193):235–237, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278361.

Watkins:1938:BRB

[Wat38a] Leonard L. Watkins. Book review: [untitled]. Journal of the American Statistical Association, 33(204):779–780, December 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279094.

Watson:1938:BRB

[Wat38b] J. P. Watson. Book review: [untitled]. Journal of the American Statistical Association, 33(202):467–468, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278932.

Watson:1939:BRBa

[Wat39a] J. P. Watson. Book review: [untitled]. Journal of the American Statistical Association, 34(207):613–614, September 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279514.

Watson:1939:BRBb

[Wat39b] J. P. Watson. Book review: [untitled]. Journal of the American Statistical Association, 34(208):777–778, December 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279841.

Waugh:1935:SMD

[Wau35] Frederick V. Waugh. A simplified method of determining multiple regression constants. Journal of the American Statistical Association, 30(192):694–700, December 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277724.

Waugh:1936:ARS

[Wau36] Frederick V. Waugh. The analysis of regression in subsets of variables. Journal of the American Statistical Association, 31(196):729–730, December 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278672.

Walker:1932:MAT

[WD32] Helen M. Walker and Walter N. Durost. A model to aid in teaching partial correlation. Journal of the American Statistical Association, 27(177):71–73, March 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277883.

Webb:1938:BRB

[Web38] John N. Webb. Book review: [untitled]. Journal of the American Statistical Association, 33(201):291–293, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279147.

Webb:1939:CUU

[Web39] John N. Webb. Concepts used in unemployment surveys. Journal of the American Statistical Association, 34(205):49–59, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279161.

Weintraub:1932:BRBa

[Wei32a] David Weintraub. Book review: [untitled]. Journal of the American Statistical Association, 27(177):119–120, March 1932. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277901.

Weintraub:1932:BRBb

[Wei32b] David Weintraub. Book review: [untitled]. Journal of the American Statistical Association, 27(179):354–355, September 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278186.

Weintraub:1932:DWT

[Wei32c] David Weintraub. The displacement of workers through increases in efficiency and their absorption by industry, 1920–1931. Journal of the American Statistical Association, 27(180):383–400, December 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278638.

Weintraub:1938:SMC

[Wei38] David Weintraub. Some measures of changing labor productivity and their uses in economic analysis. Journal of the American Statistical Association, 33(201):153–163, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279122.

Weintraub:1939:DPI

[Wei39a] David Weintraub. Did productivity increase in the twenties?: A rejoinder. Journal of the American Statistical Association, 34(206):332– 334, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278853.

Weiss:1939:BRB

[Wei39b] Harry Weiss. Book review: [untitled]. Journal of the American Statistical Association, 34(208):775–776, December 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279839.

Welch:1930:AIS

[Wel30] Emmett H. Welch. Adjusting indexes of seasonal variation for secular trend. Journal of the American Statistical Association, 25 (172):440–446, December 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277969.

Welk:1935:BRB

[Wel35] William G. Welk. Book review: [untitled]. Journal of the American Statistical Association, 30(192):785–786, December 1935. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277759.

Welch:1936:BRB

[Wel36a] Emmett H. Welch. Book review: [untitled]. Journal of the American Statistical Association, 31(193):241–243, March 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278367.

Welk:1936:BRB

[Wel36b] William G. Welk. Book review: [untitled]. Journal of the American Statistical Association, 31(194):424–425, June 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278582.

Weld:1938:PMR

[Wel38a] L. D. H. Weld. The problem of measuring radio coverage. Journal of the American Statistical Association, 33(201):117–125, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279118.

Welsh:1938:BRB

[Wel38b] Edward C. Welsh. Book review: [untitled]. Journal of the American Statistical Association, 33(202):468–469, June 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278933.

Welch:1939:BRB

[Wel39a] Emmett H. Welch. Book review: [untitled]. Journal of the American Statistical Association, 34(207):632–633, September 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279529.

Welk:1939:BRB

[Wel39b] William G. Welk. Book review: [untitled]. Journal of the American Statistical Association, 34(207):614, September 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279515.

Welsh:1939:BRB

[Wel39c] Edward C. Welsh. Book review: [untitled]. Journal of the American Statistical Association, 34(205):202–203, March 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279187.

Whitright:1934:TWS

[WF34] G. M. Whitright and M. Farrington. Taming the weekly seasonal. Journal of the American Statistical Association, 29(187):313–315, September 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278071.

Wilson:1935:SCF

[WH35] E. B. Wilson and M. M. Hilferty. Size of completed families. Journal of the American Statistical Association, 30(191):577–580, September 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2278101.

Whelpton:1934:CBR

[Whe34a] P. K. Whelpton. The completeness of birth registration in the United States. Journal of the American Statistical Association, 29(186):125– 136, June 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278282.

Whelpton:1934:RDB

[Whe34b] P. K. Whelpton. On the rapidity of the decline in the birth rate. Journal of the American Statistical Association, 29(187):308–309, September 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278069.

Whelpton:1936:BRB

[Whe36a] P. K. Whelpton. Book review: [untitled]. Journal of the American Statistical Association, 31(193):228–230, March 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278356.

Whelpton:1936:EMC

[Whe36b] P. K. Whelpton. An empirical method of calculating future population. Journal of the American Statistical Association, 31(195): 457–473, September 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278370.

White:1935:BRB

[Whi35] Wilford L. White. Book review: [untitled]. Journal of the American Statistical Association, 30(190):499, June 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277716.

White:1937:BRB

[Whi37] Wilford L. White. Book review: [untitled]. Journal of the American Statistical Association, 32(200):796–798, December 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279401.

Whitsett:1939:BRB

[Whi39] J. M. Whitsett. Book review: [untitled]. Journal of the American Statistical Association, 34(205):197–199, March 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279183.

Wilson:1931:GF

[WHM31] Edwin B. Wilson, Margaret M. Hilferty, and Helen C. Maher. Goodness of fit. Journal of the American Statistical Association, 26 (176):443–448, December 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277945.

Willcox:1930:EBS

[Wil30a] Walter F. Willcox. The eighteenth biennial session of the international statistical institute. Journal of the American Statistical Association, 25(169):69–71, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278496.

Willcox:1930:WEE

[Wil30b] Walter F. Willcox. A Westerner’s effort to estimate the population of China, and its increase since 1650. Journal of the American Statistical Association, 25(171):255–268, September 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278189.

Wilson:1930:BRB

[Wil30c] E. B. Wilson. Book review: [untitled]. Journal of the American Statistical Association, 25(170):238, June 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277647.

Wilson:1930:MS

[Wil30d] Edwin B. Wilson. Mathematics and statistics. Journal of the American Statistical Association, 25(169):1–8, March 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278489.

Willcox:1931:AMP

[Wil31] Walter F. Willcox. An attempt to measure public opinion about repealing the eighteenth amendment. Journal of the American Statistical Association, 26(175):243–261, September 1931. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277903.

Wilcoxen:1932:WPP

[Wil32a] L. C. Wilcoxen. World prices and the precious metals. Journal of the American Statistical Association, 27(178):129–140, June 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277654.

Williams:1932:BRB

[Wil32b] Pierce Williams. Book review: [untitled]. Journal of the American Statistical Association, 27(178):232–233, June 1932. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277684.

Willis:1932:BRB

[Wil32c] H. Parker Willis. Book review: [untitled]. Journal of the American Statistical Association, 27(180):448–449, December 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278651.

Willcox:1933:BRB

[Wil33a] Walter F. Willcox. Book review: [untitled]. Journal of the American Statistical Association, 28(184):483–484, December 1933. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278154.

Willis:1933:BRBa

[Wil33b] H. Parker Willis. Book review: [untitled]. Journal of the American Statistical Association, 28(183):365–366, September 1933. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278444.

Willis:1933:BRBb

[Wil33c] H. Parker Willis. Book review: [untitled]. Journal of the American Statistical Association, 28(184):465–466, December 1933. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278145.

Wilcoxen:1934:SFF

[Wil34a] L. C. Wilcoxen. Some fundamental factors entering into the wealth and income of the United States. Journal of the American Statistical Association, 29(185):79–82, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278464.

Willcox:1934:LCI

[Wil34b] Walter F. Willcox. A letter concerning the improvement of American statistics. Journal of the American Statistical Association, 29 (188):423–424, December 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278618.

Willcox:1934:LMI

[Wil34c] Walter F. Willcox. London Meeting of the International Statistical Institute. Journal of the American Statistical Association, 29 (187):310–312, September 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278070.

Willcox:1934:MIS

[Wil34d] Walter F. Willcox. Meeting of the International Statistical Institute at Mexico City. Journal of the American Statistical Association, 29(185):67–69, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278460.

Willcox:1934:NCS

[Wil34e] Walter F. Willcox. Note on the chronology of statistical societies. Journal of the American Statistical Association, 29(188):418–420, December 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278614.

Williams:1934:MCC

[Wil34f] Faith M. Williams. Measuring changes in the cost of living of Federal employees living in Washington. Journal of the American Statistical Association, 29(185):125–130, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277817.

Willis:1934:BRB

[Wil34g] H. Parker Willis. Book review: [untitled]. Journal of the American Statistical Association, 29(187):339–340, September 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278084.

Wilson:1934:BCP

[Wil34h] Edwin B. Wilson. Boole’s challenge problem. Journal of the American Statistical Association, 29(187):301–304, September 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278067.

Wilks:1935:TCS

[Wil35a] S. S. Wilks. Test criteria for statistical hypotheses involving several variables. Journal of the American Statistical Association, 30 (191):549–560, September 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278099.

Wilson:1935:BRB

[Wil35b] E. B. Wilson. Book review: [untitled]. Journal of the American Statistical Association, 30(190):487, June 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277707.

Wilks:1936:BRBa

[Wil36a] S. S. Wilks. Book review: [untitled]. Journal of the American Statistical Association, 31(194):421–422, June 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278579.

Wilks:1936:BRBb

[Wil36b] S. S. Wilks. Book review: [untitled]. Journal of the American Statistical Association, 31(196):760–762, December 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278685.

Williams:1936:BRBa

[Wil36c] Benjamin H. Williams. Book review: [untitled]. Journal of the American Statistical Association, 31(193):231–232, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278358.

Williams:1936:BRBb

[Wil36d] Faith M. Williams. Book review: [untitled]. Journal of the American Statistical Association, 31(195):610–613, September 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278395.

Willis:1936:BRB

[Wil36e] H. Parker Willis. Book review: [untitled]. Journal of the American Statistical Association, 31(193):218–220, March 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278350.

Wilks:1937:BRB

[Wil37a] S. S. Wilks. Book review: [untitled]. Journal of the American Statistical Association, 32(200):808–809, December 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279408.

Willcox:1937:AMI

[Wil37b] Walter F. Willcox. The Athens meeting of the International Statistical Institute. Journal of the American Statistical Association, 32(197):113–118, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279869.

Williams:1937:BRB

[Wil37c] Aubrey Williams. Book review: [untitled]. Journal of the American Statistical Association, 32(197):235–236, March 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279890.

Williams:1937:MMV

[Wil37d] Faith M. Williams. Methods of measuring variations in family expenditures. Journal of the American Statistical Association, 32(197): 40–46, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279859.

Wilke:1938:CEP

[Wil38a] Ernestine Wilke. Conference on employment and payroll statistics. Journal of the American Statistical Association, 33(203):575–579, September 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279698.

Wilks:1938:BRBa

[Wil38b] S. S. Wilks. Book review: [untitled]. Journal of the American Statistical Association, 33(201):273–275, March 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279136.

Wilks:1938:BRBb

[Wil38c] S. S. Wilks. Book review: [untitled]. Journal of the American Statistical Association, 33(202):478–480, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278939.

Williams:1938:BRB

[Wil38d] Faith M. Williams. Book review: [untitled]. Journal of the American Statistical Association, 33(201):275–278, March 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279137.

Wilson:1938:HSS

[Wil38e] Edwin B. Wilson. How to study the social aspects of the depression. Journal of the American Statistical Association, 33(203):505–512, September 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279688.

Wilson:1938:SDS

[Wil38f] Edwin B. Wilson. The standard deviation of sampling for life expectancy. Journal of the American Statistical Association, 33 (204):705–708, December 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279061.

Willbach:1939:PPS

[Wil39a] Harry Willbach. The purpose of parole statistics. Journal of the American Statistical Association, 34(205):87–92, March 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279166.

Williamson:1939:BRB

[Wil39b] K. M. Williamson. Book review: [untitled]. Journal of the American Statistical Association, 34(207):623–624, September 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279524.

Wilson:1939:BRB

[Wil39c] Edwin B. Wilson. Book review: [untitled]. Journal of the American Statistical Association, 34(207):591, September 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279495.

Winslow:1936:BRB

[Win36] C.-E. A. Winslow. Book review: [untitled]. Journal of the American Statistical Association, 31(194):447–448, June 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278598.

Winslow:1939:BRB

[Win39] E. M. Winslow. Book review: [untitled]. Journal of the American Statistical Association, 34(205):204–205, March 1939. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279189.

Wisniewski:1930:NI

[Wis30] Jan K. Wisniewski. Note on interpolation. Journal of the American Statistical Association, 25(170):203–205, June 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277634.

Wisniewski:1931:EFF

[Wis31] J. K. Wisniewski. Extension of Fisher’s formula number 353 to three or more variables. Journal of the American Statistical Association, 26(173):62–65, March 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278261.

Wisniewski:1934:PCC

[Wis34] Jan K. Wisniewski. Pitfalls in the computation of the correlation ratio. Journal of the American Statistical Association, 29(188):416–417, December 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2278613.

Wishart:1936:SCA

[Wis36] John Wishart. Statistics in Chinese agricultural research. Journal of the American Statistical Association, 31(193):127–128, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278335.

Witte:1936:BRB

[Wit36] Edwin E. Witte. Book review: [untitled]. Journal of the American Statistical Association, 31(194):452–453, June 1936. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278602.

Wolfe:1932:PCB

[Wol32] A. B. Wolfe. Population censuses before 1790. Journal of the American Statistical Association, 27(180):357–370, December 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278636.

Wolfe:1935:BRB

[Wol35a] A. B. Wolfe. Book review: [untitled]. Journal of the American Statistical Association, 30(191):637–638, September 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278120.

Wolfe:1935:FEM

[Wol35b] F. E. Wolfe. The factor elimination method of deriving a coefficient of net determination. Journal of the American Statistical Association, 30(192):691–693, December 1935. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277723.

Wolman:1935:BRB

[Wol35c] Leo Wolman. Book review: [untitled]. Journal of the American Statistical Association, 30(191):630–631, September 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278114.

Wolfe:1938:BRB

[Wol38] A. B. Wolfe. Book review: [untitled]. Journal of the American Statistical Association, 33(202):459–461, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278928.

Wong:1937:EVM

[Won37] Y. K. Wong. On the elimination of variables in multiple correlation. Journal of the American Statistical Association, 32(198):357– 360, June 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279577.

Woofter:1931:WNR

[Woo31] T. J. Woofter, Jr. What is the Negro rate of increase? Journal of the American Statistical Association, 26(176):461–462, December 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277948.

Woodward:1932:FCP

[Woo32a] Donald B. Woodward. Forecasts of commodity prices. Journal of the American Statistical Association, 27(178):195–201, June 1932. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277671.

Woofter:1932:IPW

[Woo32b] T. J. Woofter, Jr. Interpolation for populations whose rate of increase is declining. Journal of the American Statistical Association, 27(178):180–182, June 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277661.

Wood:1934:NES

[Woo34] Loring Wood. Note on estimating significance of a number of samples. Journal of the American Statistical Association, 29(188):422–423, December 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278617.

Woodbury:1937:BRB

[Woo37] Robert M. Woodbury. Book review: [untitled]. Journal of the American Statistical Association, 32(197):227–229, March 1937. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279886.

Woodward:1938:CCF

[Woo38] Donald B. Woodward. Changes in capital financing. Journal of the American Statistical Association, 33(201):12–20, March 1938. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279109.

Working:1934:RDS

[Wor34] Holbrook Working. A random-difference series for use in the analysis of time series. Journal of the American Statistical Association, 29(185):11–24, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278456.

Working:1937:BRB

[Wor37] Holbrook Working. Book review: [untitled]. Journal of the American Statistical Association, 32(198):413–415, June 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2279591.

Worthing:1939:BRB

[Wor39] Marion W. Worthing. Book review: [untitled]. Journal of the American Statistical Association, 34(206):429–430, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278874.

Wright:1931:BRB

[Wri31] Sewall Wright. Book review: [untitled]. Journal of the American Statistical Association, 26(175):358–360, September 1931. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277921.

Wright:1934:UFE

[Wri34a] Helen R. Wright. The use of family expenditures in cost-of-living indexes. Journal of the American Statistical Association, 29(185):118– 119, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277815.

Wright:1934:BRB

[Wri34b] Philip G. Wright. Book review: [untitled]. Journal of the American Statistical Association, 29(185):109–110, March 1934. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278477.

Wright:1937:BRB

[Wri37] Verne Wright. Book review: [untitled]. Journal of the American Statistical Association, 32(199):619–620, September 1937. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278801.

Wright:1938:BRB

[Wri38] Verne C. Wright. Book review: [untitled]. Journal of the American Statistical Association, 33(202):498–499, June 1938. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278949.

Wright:1939:BRB

[Wri39] Ivan Wright. Book review: [untitled]. Journal of the American Statistical Association, 34(205):194–195, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279180.

Wolf:1939:BRB

[WS39] Irwin D. Wolf and Lewis M. Silverstein. Book review: [untitled]. Journal of the American Statistical Association, 34(205):225–228, March 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279204.

Wubnig:1939:MTF

[Wub39] Arthur Wubnig. The measurement of the technological factor in labor productivity. Journal of the American Statistical Association, 34(206):319–325, June 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278851.

Wurzburger:1936:CDR

[Wur36] Eugen Wurzburger. On the comparability of death rates. Journal of the American Statistical Association, 31(194):381–385, June 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278566.

Woofter:1938:HPR

[WW38] T. J. Woofter, Jr. and T. E. Whiting. Households and persons receiving relief or assistance. Journal of the American Statistical Association, 33(202):363–372, June 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278912.

Yang:1933:PV

[Yan33] Simon Yang. On partition values. Journal of the American Statistical Association, 28(182):184–191, June 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278032.

Yates:1934:AMC

[Yat34] F. Yates. The analysis of multiple classifications with unequal numbers in the different classes. Journal of the American Statistical Association, 29(185):51–66, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278459.

Young:1932:VDM

[YD32] Kimball Young and C. L. Dedrick. Variation in the duration of marriages which end in divorce, with special reference to the state of Wisconsin. Journal of the American Statistical Association, 27(178):160– 167, June 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277657.

Yntema:1930:BRBa

[Ynt30a] Theodore O. Yntema. Book review: [untitled]. Journal of the American Statistical Association, 25(169):126, March 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278520.

Yntema:1930:BRBb

[Ynt30b] Theodore O. Yntema. Book review: [untitled]. Journal of the American Statistical Association, 25(172):495, December 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277987.

Yntema:1932:BRB

[Ynt32] Theodore O. Yntema. Book review: [untitled]. Journal of the American Statistical Association, 27(177):106–108, March 1932. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277893.

Yntema:1933:MIP

[Ynt33] Dwight B. Yntema. Measures of the inequality in the personal distribution of wealth or income. Journal of the American Statistical Association, 28(184):423–433, December 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278132.

Yntema:1935:BRBa

[Ynt35a] Theodore O. Yntema. Book review: [untitled]. Journal of the American Statistical Association, 30(189):124, March 1935. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278543.

Yntema:1935:BRBb

[Ynt35b] Theodore O. Yntema. Book review: [untitled]. Journal of the American Statistical Association, 30(191):632–633, September 1935. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278116.

Yntema:1937:BRB

[Ynt37] Theodore O. Yntema. Book review: [untitled]. Journal of the American Statistical Association, 32(199):604–605, September 1937. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278793.

Yoder:1933:BRB

[Yod33] Dale Yoder. Book review: [untitled]. Journal of the American Statistical Association, 28(181):106–109, March 1933. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278015.

Yoder:1934:BRB

[Yod34] Dale Yoder. Book review: [untitled]. Journal of the American Statistical Association, 29(186):230, June 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2278307.

Yoder:1938:SS

[Yod38] Dale Yoder. Seasonality in strikes. Journal of the American Statistical Association, 33(204):687–693, December 1938. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279059.

Young:1932:BRB

[You32] W. H. Young. Book review: [untitled]. Journal of the American Statistical Association, 27(179):350–351, September 1932. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2278183.

Young:1936:BRBb

[You36a] Robert V. Young. Book review: [untitled]. Journal of the American Statistical Association, 31(196):766–767, December 1936. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278689.

Young:1936:BRBa

[You36b] W. H. Young. Book review: [untitled]. Journal of the American Statistical Association, 31(193):243–244, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278368.

Zabiello:1930:CSS

[Zab30] G. Zabiello. Concentration of statistical services in Italy. Journal of the American Statistical Association, 25(169):80–81, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2278499.

Zapoleon:1931:BRB

[Zap31] L. B. Zapoleon. Book review: [untitled]. Journal of the American Statistical Association, 26(174):234–235, June 1931. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277785.

Zelomek:1934:RVS

[ZDM^{+}34] A. W. Zelomek, M. M. Drake, Edmund J. Murphy, Howard G. Brunsman, Aryness Joy, and F. L. Carmichael. The reliability of various statistical series as city and regional business indexes. Journal of the American Statistical Association, 29(185):86–93, March 1934. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277809.

Zelomek:1936:BRB

[Zel36] A. W. Zelomek. Book review: [untitled]. Journal of the American Statistical Association, 31(193):239, March 1936. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2278364.

Zimmermann:1939:BRB

[Zim39] Erich W. Zimmermann. Book review: [untitled]. Journal of the American Statistical Association, 34(206):458–459, June 1939. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2278896.

Zubin:1939:NDS

[Zub39] Joseph Zubin. Nomographs for determining the significance of the differences between the frequencies of events in two contrasted series or groups. Journal of the American Statistical Association, 34 (207):539–544, September 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279487.